A Complete Bibliography of Publications in The Journal of Supercomputing: 1987–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

06 April 2023 Version 3.01

Title word cross-reference

1 [299, 421]. 2 [299, 421, 478, 537, 806]. 3 [251, 299, 314, 691, 768]. 5 [299]. 9 [154]. _{RP} [669]. d [537]. H4 [460]. K [414, 487, 498, 707, 826]. kr ⟶ r [255]. μ [447]. N [257, 312, 414, 768, 826]. p [693]. r ⟶ kr [255].

-Ary [414, 498, 768, 826]. -Body [257]. -Cubes [414, 768, 826]. -D [299, 314, 421, 478, 691]. -Dimensional [312]. -means [487, 707]. -Point [154]. -sized [693].

[?]CoarseKonstantopoulos:2009:EPT.

1.0 [203]. 14/300 [142]. 1985 [1]. 1987 [25].

2 [5, 22, 58, 67, 79, 80, 91, 104, 112, 122, 150, 186, 580]. 2004 [656]. 21st [382, 423].

3 [117]. 3-Hierarchical [839].

40 [79]. 416 [80].

5 [139, 140, 166].

802.11 [807, 808]. 860 [125].

90 [348, 471].

absolute [836]. accelerate [92].

Accelerator [566]. Accelerators [486]. Access [73, 202, 239, 749, 813, 822, 875, 878]. access-efficient [875]. access-weights [822]. Accuracy [232, 500, 505]. Accurate [346, 602]. Achieving [306]. ACO [492]. Acquisition [562]. Across [624]. action [792]. Active [368, 751, 893]. Activity [533]. Ad [579, 609, 678, 779, 829, 901]. ad-hoc [779]. adaptation [808]. Adapted [358]. Adapting [156, 227]. Adaptive [63, 291, 325, 382, 390, 399, 414, 423, 528, 576, 696, 715, 790, 795, 817, 823, 824, 833, 847, 892]. Addition [551]. Address [327]. Addressable [355]. Addressing [692]. Adjust [320]. Adjusted [601]. adjustment [735]. Admission [546, 585]. Advanced [9, 604, 876]. Advances [117, 782]. Adventures [232]. Aerodynamic [548]. aerodynamics [69]. Aerosol [212]. Affine [327]. Affinity [675]. Agent [368, 544, 573, 773, 853]. agent-based [773]. Aggregation [648, 783]. Aggregation/Dissaggregation [648]. aggressive [785]. AIAC [646]. Aid [44]. Air [531]. Alamos [381, 422]. Algebra [106, 743]. Algebraic [40, 445]. Algorithm [5, 11, 48, 58, 92, 105, 126, 129, 161, 244, 251, 260, 268, 288, 289, 318, 325, 360, 390, 412, 455, 463, 465, 482, 513, 543, 545, 552, 559, 597, 598, 606, 616, 642, 680, 681, 685, 691, 693, 695, 707, 712, 753, 773, 775, 777, 799, 804, 817, 857, 884]. Algorithm-to-Machine [244]. Algorithmic [701]. Algorithms [10, 15, 26, 52, 91, 133, 155, 164, 178, 200, 201, 212, 215, 247, 249, 257, 269, 286, 292, 298, 299, 306, 330, 370, 398, 434, 462, 488, 509, 525, 544, 546, 567, 637, 646, 647, 655, 659, 678, 694, 713, 717, 727, 731, 774, 780, 790, 806, 813, 818]. Align [464]. Alignment [231, 371, 400, 464, 570]. all-pairs [725]. Allocating [155, 624]. Allocation [149, 320, 354, 477, 675, 706, 735, 759, 801, 802, 850, 854]. Alloyed [522]. Alternative [27]. alternatives [730]. amalgamation [830]. AMIDAR [593]. among [98]. Analysis [32, 36, 65, 77, 105, 126, 163, 234, 246, 253, 254, 262, 270, 328, 333, 377, 387, 402, 428, 445, 462, 467, 483, 509, 595, 603, 647, 682, 697, 728, 730, 737, 741, 753, 788, 819, 820, 842, 897]. Analytical [252, 441, 724]. Analyzed [350]. Annealing [404, 477]. announcement [39, 46]. anomaly [808]. Ant [462, 670, 873]. Ant-based [670]. Anycast [546]. aperiodic [789]. Application [104, 247, 307, 321, 347, 356, 376, 389, 441, 451, 481, 579, 622, 859]. Application-Driven [347]. Application-Level [307]. Application-Tailored [356, 389]. Applications [9, 95, 195, 197, 273, 287, 346, 475, 479, 519, 521, 527, 528, 540, 556, 566, 568, 595, 612, 622, 643, 656, 658, 662, 681, 696, 699, 700, 715, 726, 748, 752, 795, 796, 834]. Applied [59, 687, 726]. Approach [48, 86, 207, 224, 227, 351, 496, 544, 548, 550, 595, 605, 641, 655, 666, 729, 763, 794, 801, 849]. Approaches [504, 618, 639]. Approximate [59, 469, 470, 526, 557, 764]. approximation [698, 797, 823]. Architectural [219, 684, 722, 764, 862]. Architecture [68, 102, 139, 177, 208, 245, 250, 340, 356, 389, 450, 489, 492, 495, 496, 502, 514, 567, 596, 601, 607, 631, 649, 650, 653, 667, 702, 716, 729, 798, 824, 839, 848, 851, 881, 896]. Architecture-Independent [340]. Architectures [15, 20, 27, 44, 103, 111, 167, 270, 283, 284, 352, 449, 494, 536, 542, 586, 746, 825, 831, 894]. Archived [223]. Area [365, 629, 701, 713, 761]. area-efficient [701]. Arising [154]. Arithmetic [57, 71, 314]. Arithmetics [346]. Array [84, 95, 180, 187, 203, 224, 240, 255, 259, 301, 319, 348, 371, 417–419, 434, 457, 474, 551, 586, 657, 688, 725, 729, 828, 832, 865]. Arrays [83, 188, 202, 296, 312, 352, 488, 575, 577, 632, 660, 744, 862, 886]. Articles [311]. Artificial [113]. Ary [414, 498, 768, 826]. Aspects [233, 701]. Assembly [326]. Assembly-Level [326]. Assessment [367, 597]. Assignment [395, 571, 651]. assisted [862]. associated [747]. association [719]. Asynchronous [223, 365, 442, 501, 745]. Atomic [157, 902]. Attacks [666]. Auditing [336]. Augmented [498]. Authentication [609]. Authors [3, 18, 23, 30, 37, 45, 50, 55, 60, 66, 72, 82, 88, 93, 100, 108, 116, 121, 127, 134, 143, 147, 152, 160, 165, 181, 189, 194, 199, 204, 209, 216, 221, 228, 235, 242, 435, 517, 518, 555, 563, 569, 574, 583, 588, 594, 599, 611, 615, 620, 628, 672, 683]. Auto [42, 700]. Auto-CFD-NOW [700]. Auto-Scheduling [42]. Automatic [56, 85, 149, 261, 324, 339, 340, 452, 488, 516, 587, 662, 867]. availability [868]. Available [629]. Average [162]. Aware [585, 624, 668, 767, 808, 827, 875].

b [407, 898]. Backup [501, 670]. Balance [530]. Balanced [49]. Balancing [158, 246, 291, 295, 463, 504, 529, 581, 637, 755, 795, 840, 887]. band [886]. Banded [470, 541]. Bandwidth [601, 624, 629, 901]. Bandwidth-Aware [624]. Bandwidth-based [901]. Banyan [603]. Banyan-Type [603]. Barrier [111]. base [898]. base- [898]. Based [70, 113, 159, 250, 267, 274, 309, 316, 337, 355, 363, 461, 465, 466, 471, 472, 481, 530, 547, 549, 566, 596, 610, 616, 618, 623, 632, 640, 648, 650, 662, 664, 670, 688, 694, 731, 739, 748, 754, 766, 773, 775, 791, 809, 812, 823, 834, 843– 845, 849, 851, 855, 857, 859, 866, 868, 869, 881, 889, 892, 899, 901]. Bases [743]. Basic [324]. BC [816]. BCOOPL [448]. Beating [230]. Behavior [162, 711, 874]. Behaviors [437]. Beliefs [617]. Benchmark [65, 275]. Benchmarking [79, 627]. Benchmarks [277]. Between [162, 315, 759]. Beyond [192]. Biconnected [99]. Bidirectional [597]. Biharmonic [59]. Bin [49]. Binary [578, 720]. bisecting [707]. Bistability [146]. BLAS [70, 112, 133, 145]. Block [293, 370, 444, 496, 598, 613, 815].

Block-Cyclic [293, 613, 815].

Block-Toeplitz [598]. Block-tridiagonal [444]. Block/Machine [496]. blockcipher [712]. Blocking [63]. Blood [230]. Body [257]. Boiler [271]. Book [17]. Boolean [240]. Boost [91, 723]. Boosting [359]. bottleneck [818]. Boundary [230, 470, 506]. bounded [745]. Bounding [251]. Bounds [575]. Bounteous [241]. Broadcast [297, 554, 649, 678, 902]. Broadcasting [379, 478]. Broker [267, 734, 845]. brokering [878]. BSP [329, 343, 870]. BSR [392, 694]. BSR-based [694]. Buffering [354]. build [661, 714]. Building [396, 732, 877]. Bulk [270, 393, 645, 870]. Bursty [484]. Bus [240, 296, 434, 551, 865]. Buses [458]. Business [550].

C [96, 187]. C4 [178]. C4/XA [178]. Cache [159, 168, 417, 512, 578, 642, 669, 711, 717]. Caches [241, 714]. Caching [650, 807]. Calculation [606]. Can [320]. canceller [809]. Candidate [480]. Capability [498, 649]. Carlo [69, 107, 126, 213, 335, 413, 806]. CAS [472]. CAS-System [472]. Case [244, 248, 331, 441, 561, 658, 743, 788]. Causal [336, 416]. CC [338]. CC-NUMA [338]. Cedar [75]. Cell [150, 213]. CEMA [573]. center [747]. Century [382, 423]. CertiVeR [771]. CFD [408, 472, 700]. CFT77 [71]. CGM [870]. Chain [410]. Challenges [9]. Channel [320, 407, 813, 826]. channels [809]. Character [179]. Characteristic [743]. Characterization [193, 294, 624, 692, 825]. Checkpointing [225, 308, 416, 605, 835, 892]. checkpointing-related [835]. Checkpoints [640]. Chip [560, 825]. choice [791]. Cholesky [133, 269, 441, 459]. Chromosome [220]. Chromosomes [247]. Chronos [294]. Circuit [379, 460]. Circuit-Switched [379]. Circuits [78, 280, 313, 314, 568, 641]. Circulation [214]. Class [261, 280, 593]. Classification [233]. classroom [749]. client [714]. Cluster [220, 266, 276, 316, 321, 362, 363, 365, 366, 390, 466, 472, 476, 595, 621, 627, 630, 633, 661, 665, 675, 724, 737, 784]. Cluster-based [316]. Clustered [679]. Clustering [176, 281, 364, 487, 559, 665, 707, 777, 829, 867]. Clusters [274, 334, 415, 500, 577, 614, 624, 637, 671, 709, 730, 793, 841, 859, 899, 900]. CM [67, 166]. CM-2 [67]. CM-5 [166]. CMP [825]. CNI [227]. Co [418, 624, 657, 665, 735]. Co-Allocating [624]. co-allocation [735]. Co-Array [418, 657]. co-processing [665]. Coarse [249, 495, 862, 870]. Coarse-Grained [249, 495]. Code [125, 253, 291, 471, 738, 899]. Codebook [113]. Codes [79, 150]. coding [811]. cognitive [842]. Coherence [159]. Collaboration [223]. Collaborative [408, 409, 431, 565]. Collection [61, 416]. Collective [515, 844, 855]. Collision [213, 619]. collisions [705]. colony [873]. Color [562]. Column [459]. Column-Oriented [459]. COMA [205]. Combating [818]. Combined [443]. Comment [797]. Comments [372, 436]. Commercial [479]. Commodity [276, 777]. Common [260, 586]. Communication [149, 195, 208, 282, 286, 300, 306, 309, 321, 342, 347, 356, 357, 371, 374, 389, 475, 484, 499, 567, 627, 687, 716]. Communication-Free [300, 371]. Communication/Computation [306]. Communications [98, 336, 515, 601]. COMOPS [275]. compaction [769]. Comparative [401]. Comparison [22, 67, 80, 171, 197, 418, 646, 709]. comparisons [716]. Competitive [113, 246]. Compilation [41, 68, 95, 132, 141, 351, 399, 443, 591, 659]. Compile [258, 330]. Compiler [71, 85, 138, 259, 700, 729, 862, 895]. Compilers [42, 254, 300, 382, 423, 543, 733, 819, 820]. Compiling [33, 34, 76, 140, 400, 748].

Complement [307]. Complete

[166, 375, 582]. Complex [299, 399, 507, 610]. Complexity [65, 293, 333, 473]. Component [448, 450, 889]. component-based [889]. Components [64, 99, 440]. Composite [179]. Compositing [344]. composition [721, 854]. comprehensive [730]. Compressed [541]. compressing [738]. compression [721]. Computation [230, 286, 306, 341, 343, 529, 539, 596, 693, 778, 895]. Computational [62, 86, 102, 233, 293, 385, 412, 426, 468, 566, 576, 635, 792, 845]. Computations [8, 10, 219, 223, 345, 393, 500, 592, 626, 645, 711, 863]. Computer [76, 130, 142, 215, 245, 285, 337, 381, 422, 560, 656, 743, 792]. Computers [10, 29, 79, 175, 202, 213, 290, 299, 318, 526, 627]. Computing [64, 81, 84, 96, 99, 117, 177, 183, 222, 227, 236, 266, 271, 276, 288, 305, 311, 362, 364, 365, 399, 441, 445, 468, 474, 479, 480, 502, 504, 521, 562, 573, 585, 605, 621, 630, 645, 671, 673, 697, 706, 725, 743, 758, 762, 763, 782, 784, 801, 831, 835, 865, 876, 893]. Concepts [263, 295, 760]. Concurrency [454, 533]. Concurrent [12, 636, 654, 684, 722]. conditional [816, 883, 890]. cone [864]. Conferencing [207]. Confidence [461]. Confidence-Interval [461]. Configurable [485, 491, 496, 564, 589]. Configuration [282]. Configuring [398]. Conflict [119, 723, 794]. Conflicts [417]. Conformational [126]. Congestion [440, 668]. congestion-aware [668]. Conjugate [289]. Connect [499]. Connected [64, 318, 710]. Connection [67, 150, 573]. Connections [639]. connectivity [860, 898]. Connectors [53]. Conquer [40, 248, 543, 790]. Conscious [316, 403]. Conservation [373]. Consistency [168, 692]. Consistent [442, 562]. Constant [240]. constrained [805]. Constrains [625]. Constraints [266]. Constructing [460]. Construction [535].

Constructs [43]. consumption [730, 752, 825]. Content [355, 385, 426]. Contention [340, 554, 733]. Contention-Free [340, 554, 733]. Contents [534]. Context [27, 493, 808, 812]. context-aware [808]. continuation [851]. continuation-based [851]. continued [11]. Continuous [561]. Contraction [407]. contrarian [763]. Contributing [3, 18, 23, 30, 37, 45, 50, 55, 60, 66, 72, 82, 88, 93, 100, 108, 116, 121, 127, 134, 143, 147, 152, 160, 165, 181, 189, 194, 199, 204, 209, 216, 221, 228, 235, 242, 435, 517, 555, 563, 569, 574, 583, 588, 594, 599, 611, 615, 620, 628, 672, 683]. Control [185, 298, 415, 454, 533, 546, 585, 750, 779, 789, 878, 882]. controllable [872]. Controlling [448]. Convection [526]. Conventional [20, 27, 277]. Convergence [492, 561]. Convergent [52]. Convex [178, 392]. Convolution [288]. Cooperative [650]. cooperatively [671]. Coordinated [451]. Coordination [447, 449, 453, 455]. coprocessor [718]. CORBA [295]. Core [689]. Correcting [212]. Correction [433]. Correlation [244]. Corruptions [265]. coscheduling [723]. Coset [53]. Cost [12, 287, 378, 832, 879]. Cost-Effective [287]. Cost-effectiveness [12]. Costs [293]. Coupled [584]. Cover [114]. Covering [486]. CPU [752]. crash [897]. CRAY [5, 16, 22, 47, 58, 71, 79, 80, 91, 104, 112, 118, 122, 128, 151, 193, 215, 842]. CRAY-2 [5, 22, 58, 79, 80, 91, 104, 112, 122]. Creating [762]. Creative [697]. Critical [177, 461]. Cross [499, 802]. Cross-Connect [499]. crossed [883, 888]. CT [691, 857]. cube [4, 515]. Cubes [414, 768, 826, 883]. Current [757, 775, 831]. Curve [115]. Custom [183]. Customisable [591]. Customised [368]. Cycle [602]. Cycle-Accurate [602]. cycles [883]. Cyclic [293, 613, 815]. Cydra [139, 140].

D [251, 299, 314, 421, 478, 537, 691, 806].

Daniel [1, 211]. Data [8, 90, 95, 105, 149, 155, 163, 185, 187, 288, 327, 354, 380, 391, 400, 403, 411, 464, 472, 477, 481, 489, 535, 541, 632, 639, 665, 676, 687, 706, 713, 718, 721, 727, 735, 744, 775, 776, 783, 815, 819, 820, 822, 832, 837, 846, 852]. Data-Driven [185]. Data-Parallel [187, 327]. data-path [718]. Database [501, 533, 706, 800, 897]. databases [719]. Datagram [377]. Datasets [234]. dBuses [312]. DDM [648]. DDM-Based [648]. DDoS [666]. Deadline [682, 789]. Deadlock [572]. Deadlock-Free [572]. Debug [267, 590]. Debug-By-Design [590]. Debugging [169]. Decentralized [173, 879]. Decision [720]. Decomposition [107, 292, 335]. Decoupled [284, 702]. decrements [838]. decryption [712]. Dedication [1, 25]. Deep [457]. defined [383, 424]. Defining [757]. Definition [95, 451]. Degradation [686, 808]. Degree [256]. Delay [484, 546]. Delivering [385, 426]. Delivery [513]. Delta [166]. Dense [14]. Dependable [309]. Dependence [32, 77, 90, 575, 819, 820]. Dependence-Analysis [77]. dependent [805, 879]. Deployment [402]. Depth [212, 280, 460]. Depth-Size [280, 460]. Derivative [482]. Derived [261]. Deriving [301]. description [889]. Design [20, 113, 176, 305, 315, 352, 386, 388, 402, 427, 429, 548, 587, 590, 597, 604, 653, 677, 684, 692, 702, 705, 716, 727, 729, 737, 832, 845]. Designed [356, 389]. Designing [619]. Designs [172, 638]. Desktop [623]. destination [705]. destruction [697]. detecting [666]. detection [774, 821]. Determining [575, 630, 712]. Deterministic [296, 508, 647, 678]. Deterministically [526]. developing [658]. Development [328, 334, 402, 695, 834, 850]. device [87, 751]. device-level [751]. DGMonitor [623]. DHTs [787]. diagnosability [816]. Diagnostic [77]. Diagonals [541]. Diagrammatic [654].

Diagrams [720]. Dialects [75]. Different [197, 622]. Differential [47]. Diffusion [322, 433, 526]. Digit [6, 314, 414]. Digit-Reversal [414]. Digital [562]. Dimensional [48, 115, 213, 218, 249, 288, 312, 319, 323, 379, 526, 575, 630, 727, 769, 791]. Dimensioning [760]. Direct [544, 558, 662, 843]. Directions [7]. directive [899]. directive-based [899]. directories [758]. Directory [652]. Discovering [635]. Discovery [368, 652, 758, 873]. Discrete [83, 699]. Discretization [154]. Disease [480]. Disjoint [619]. Disk [403, 660, 661, 832]. Disk-Resident [403]. Disks [675]. Dissaggregation [648]. Distance [461, 474, 537, 708]. Distance- [537]. distance-hereditary [708]. Distributed [33, 157, 167, 172, 175, 198, 210, 222, 223, 225, 258, 263, 265, 283, 285, 287, 291, 293, 299, 302, 303, 310, 328, 331, 336, 344, 360, 380, 400, 421, 442, 455, 467, 500, 504, 511, 526, 530, 540, 542, 547, 550, 556, 577, 597, 613, 617, 622, 625, 632, 644, 648, 667, 676, 679, 684, 696, 706, 721, 744, 745, 762, 789, 813, 835, 846, 854, 864, 868, 891]. Distributed-Memory [33, 157, 167, 175, 299, 303]. Distributed-Shared-Memory [198]. Distribution [257, 259, 410, 477, 632, 744, 853, 856]. Distribution-Independent [257]. Distributions [16, 301, 535]. Distributive [268, 390]. Divide [40, 248, 543, 790]. Divide-and-Conquer [40, 248, 543, 790]. divisible [760]. Do [457]. DOALL [158]. Document [559]. Domain [335, 494, 809]. Domain-Specific [494]. domains [767]. dominant [799]. domination [708]. down [387, 428]. Downlink [802]. Driven [185, 347, 522]. DSM [273, 640, 677]. DTM [524, 709]. Dual [245, 515, 610, 798]. Dual-Cube [515]. Dual-execution [798]. Dual-Tree [610]. due [808]. duplicate [821]. Duplication [466, 799]. DVS [750].

Dynamic [62, 290, 320, 322, 350, 353, 384, 393, 402, 425, 450, 512, 529, 541, 561, 571, 581, 585, 645, 669, 670, 675, 696, 705, 735, 759, 793, 795, 808, 822, 838, 850, 866, 882, 887]. dynamical [874]. Dynamically [356, 389, 490, 553]. Dynamics [150, 232, 566, 841].

Earliest [682]. Earliest-Deadline-First [682]. Early [666]. Earth [536]. economic [766]. economy [892]. Edge [465, 768, 865, 890]. edge-pancyclicity [768]. Edge-Region-Based [465]. Editor [15, 64, 70, 86, 87, 394, 468, 479]. Editorial [2, 19, 24, 38, 51, 89, 101, 148, 182, 190, 217, 222, 236, 243, 272, 362, 372, 405, 436, 468, 519, 540, 556, 600]. Editors’ [136, 405, 485, 491, 519, 770]. EDPEPPS [294]. Effect [637]. Effective [141, 287, 642, 677, 877]. Effectively [163, 700]. Effectiveness [12, 239]. Effects [35, 87]. Efficiency [440]. Efficiency/Fairness [440]. Efficient [26, 107, 254, 255, 260, 278, 288, 296, 319, 323, 327, 342, 344, 360, 370, 380, 416, 421, 434, 453, 458, 463, 495, 511, 513, 515, 538, 545, 553, 558, 592, 598, 604, 616, 632, 687, 694, 701, 708, 721, 734, 774, 783, 784, 806, 810, 854, 870, 875, 878, 885]. Effort [438]. Eigenvalues [286]. EKMR [632]. EKMR-Based [632]. El’brus [117]. El’brus-3 [117]. election [745]. Electromagnetic [213]. Electron [124]. Electron-Solid [124]. Element [131, 291, 322, 520, 843, 871]. Elements [417]. Eliminates [641]. email [867]. Embedded [304, 305, 309, 548, 602, 616, 643, 718]. Embedding [883]. embeddings [164]. Empirical [522, 597]. Enabled [642]. Enabling [385, 426, 715]. Encoding [490]. encryption [712]. encryption/decryption [712]. End [514, 789]. End-to-end [514, 789]. Energy [86, 373, 494, 495, 592, 730, 752, 767, 783, 825]. Energy-aware [767]. Energy-Efficient [495, 592, 783]. Enforced [492].

Engineering [349, 394, 412, 485, 491, 527, 564, 589]. Enhanced [404, 464, 633, 860]. Enhancement [201]. Enterprise [622]. Environment [186, 265, 451, 602, 667, 713, 762, 868]. Environments [224, 367, 441, 519, 573, 585, 633, 635, 646, 734, 735, 737, 749, 754, 772, 875]. Equation [47, 129]. Equations [14, 59, 269, 413, 526, 584]. ER-TCP [784]. Erlang [823]. Erratum [388, 430, 497, 855]. Error [433]. Errors [438]. Establishment [573]. Estimates [461]. Estimating [212]. Estimation [118, 151, 293, 494]. Estimators [254]. Ethernet [266]. Euclidean [474]. Euler [123, 167]. Evaluation [7, 142, 219, 376, 386, 427, 439, 482, 505, 576, 621, 625, 631, 639, 644, 648, 653, 702, 850, 894]. Event [267, 522]. Event-Based [267]. Event-Driven [522]. every [856]. Examine [244]. Exchange [166, 573, 858]. Exclusion [310, 877]. Execute [614]. Execution [35, 104, 393, 503, 544, 798]. executor [844, 855]. Experience [487]. Experiences [77, 105, 657]. Experiment [275]. Experimental [142, 201, 439, 467]. Experiments [373, 532]. Explanation [686]. Explicit [195, 557, 562]. Exploitation [326, 777]. Exploiting [163, 338, 466, 714]. Exploration [481, 729, 862]. Exploring [659, 718, 796]. Extended [454, 462]. extensible [824]. Extensions [85]. External [74].

Facility [384, 425]. Factor [269]. factorials [693]. Factoring [10, 11, 795]. Factorisation [369]. Factorization [63, 130, 133, 293, 421, 441, 459, 688]. Factorized [469]. Factors [443, 562]. failures [897]. fair [779]. Fairness [440]. Family [4, 580]. Farewell [217]. Fast [5, 200, 212, 302, 460, 470, 482, 545, 614, 661, 669, 693, 887, 896]. Fast-path [896]. fat [739]. fat-tree-based [739]. Fault [177, 287, 304, 305, 307, 308, 310, 366, 378, 478, 498, 619, 680, 753, 784, 827, 856, 883, 890, 900]. Fault-aware [827]. fault-free [883]. fault-tolerance [784]. Fault-Tolerant [177, 304, 305, 308, 366, 378, 478, 498, 680, 753, 890, 900]. Faults [680, 705, 883, 890, 892]. faulty [745]. FDDI [239]. FDDI-M [239]. federated [671]. Feedback [561, 750]. Fermat [179]. Fernbach [94]. FFT [91, 107, 178, 248, 299, 651]. FFTS [48, 74, 122]. Field [188]. File [265, 287, 308, 853, 856]. File-I [308]. File-I/O [308]. files [837]. Finding [459]. Fine [41, 878]. Fine-Grain [41]. fine-grained [878]. Finite [131, 291, 322, 843]. Firewall [355]. Firmware [483]. Firmware-Level [483]. First [76, 292, 510, 682]. First-Order [76]. Fixed [256]. Flames [406]. Flexible [356, 389, 495, 704, 815]. Flow [27, 105, 163, 230, 345, 415, 520, 546, 871]. Fluid [345, 407, 566, 871]. Flux [552]. Focus [311]. focused [834]. forecast [747]. Foreword [394, 456, 485, 491, 564, 589, 663, 770]. fork [56]. fork-join [56]. Form [370]. Formal [32, 897]. formalism [889]. Formally [350]. Fortran [75, 145, 348, 418, 471, 657]. Fourier [5, 482, 584]. FPGA [355, 488, 590, 643, 681, 688, 689, 725]. FPGA-Based [355, 688]. FPGAs [580, 592]. fraction [11]. Fragmentation [438]. Framework [264, 380, 396, 397, 447, 521, 585, 587, 660, 761, 776, 812, 824, 833, 848, 850, 862]. frameworks [658]. Free [300, 340, 371, 499, 554, 572, 733, 883]. Free-Space [499]. Frequent [676]. FT64 [849]. Fully [432]. Functional [234, 571, 690]. Functions [348, 445, 837]. fundamental [694]. further [794]. fusion [785]. Future [110, 356, 389]. Fuzzy [461, 750]. fuzzy-DVS [750].

Garbage [61, 416]. Gate [188]. gathering [783]. Gauss [325]. Gaussian [97, 288]. GEN BLOCK [419]. General [47, 278, 626, 698, 797, 819, 820, 895]. general-purpose [895]. Generalized [354, 433, 470, 674, 695, 898]. generated [817]. Generation [56, 327, 475, 903]. generator [899]. Generators [16, 634]. generic [729]. Genes [480]. Genetic [161, 231, 655, 685]. Geometry [62]. Gigabit [483]. Global [144, 202, 224]. Globally [52]. Glossary [31]. Gossiping [678]. GPU [665]. GPUs [895]. Gradient [289]. Grain [41]. Grained [249, 495, 862, 870, 878]. Granularity [21, 52]. Graph [247, 322, 434, 549, 616, 694]. Graph-Theoretic [434]. Graphical [340]. Graphs [102, 131, 291, 582, 708, 710, 890]. Grid [247, 552, 604, 621, 623, 631, 635, 652, 667, 673, 674, 676, 704, 734, 735, 737, 743, 748, 749, 754, 756– 758, 763, 766, 772, 775, 801, 817, 827, 833, 839, 872, 873, 875, 876, 878, 879, 881, 882, 892, 893]. Grid-based [748, 775]. Grid-Pyramid [674]. Grids [123, 292, 385, 426, 760, 771, 822, 841, 845, 877]. Gröbner [743]. Group [102, 336, 409, 453, 530]. Group-Based [530]. Group-Oriented [336]. Grouping [317, 614]. Growing [201]. GSL [861]. Guest [136, 362, 394, 405, 468, 479, 485, 491, 519, 556, 600, 770]. Guided [503, 561]. Guidelines [144].

H [343]. H-BSP [343]. hadrodynamics [128]. Hamiltonian [616]. Hamiltonicity [890]. handle [858]. Handling [472]. handoff [807]. Hardware [185, 193, 349–351, 353, 356, 389, 396, 398, 399, 481, 490, 591, 643, 701, 725, 739, 777, 778]. Hardware/Software [349, 356, 389, 701]. harmonic [699]. Hash [580, 650, 681]. heap [120]. Heart [230]. Heat [345]. hereditary [708]. Hessenberg [370]. Heterogeneity [466, 727]. Heterogeneous [238, 261, 285, 366, 441, 538, 554, 645, 668, 690, 724, 793, 799, 810, 814, 868, 902]. Heuristic [253, 554, 597, 791, 801]. Heuristics [114, 759]. Hiding [205, 315]. Hierarchical [74, 176, 257, 281, 343, 528, 578, 581, 665, 702, 758, 772, 776, 829, 839, 882]. Hierarchically [168]. Hierarchy [29, 525, 853]. High [5, 9, 48, 84, 86, 115, 117, 145, 178, 202, 215, 222, 226, 227, 236, 299, 345, 355, 362, 364–366, 381, 383, 412, 422, 424, 479, 480, 500, 527, 531, 590, 605, 629, 634, 639, 656, 681, 689, 703, 706, 718, 727, 741, 746, 764, 779, 782, 831, 835, 859, 896]. High-Dimensional [115, 727]. high-energy [86]. High-Level [383, 424, 689]. High-Performance [5, 9, 117, 145, 178, 202, 215, 227, 236, 299, 365, 366, 381, 412, 422, 479, 480, 656, 703, 706, 718, 835, 859]. High-Speed [84, 362, 629, 681, 741]. High-Visibility [590]. Higher [433, 526]. Higher-Order [433]. Highly [43, 218]. History [137, 503, 747]. Hoc [579, 609, 678, 779, 829, 901]. home [812]. homology [731]. Homotopy [115]. Hop [579]. Host [54]. Hot [508]. Hot-Spot [508]. Hough [81, 545]. HPC [900]. HPCVIEW [387, 428]. HPF [464]. human [803]. Hybrid [487, 559, 595, 881, 884]. Hyper [281, 740, 823]. hyper-Erlang [823]. Hyper-Ring [281]. hyper-threading [740]. Hypercube [26, 28, 29, 44, 49, 52, 54, 64, 81, 99, 207, 378, 463, 765, 840, 898]. Hypercube-to-Host [54]. Hypercubes [62, 131, 164, 286, 297, 508, 860, 888]. Hypermatrix [830]. Hypermesh [369]. Hyperplane [614].

I-Match [821]. I-O [818]. I/O [203, 308, 403, 664, 746, 751, 831, 835, 844, 855, 896]. I/O-Conscious [403]. IBM [145]. IceT [431]. ICGLS [542]. ICI [809]. identification [864]. Identifying [480]. Idling [373]. IEEE [807, 808]. IF [71]. II [491, 606]. III [564]. ILU [444]. Image [201, 233, 244, 252, 288, 292, 691, 721, 727].

Image-Space [292]. Imagery [212, 524]. Images [251, 562]. Immersed [230]. Impact [727, 741]. impacts [751]. Implement [568]. Implementation [13, 26, 112, 113, 123, 139, 150, 157, 167, 188, 191, 214, 230, 231, 376, 386, 427, 492, 520, 580, 587, 604, 631, 677, 681, 691, 737, 832, 844– 846, 853, 855]. implementations [657, 715]. Implemented [145]. Implementing [525, 646, 688]. Implicit [432, 723]. Improve [346, 771]. Improved [270]. improvement [741]. Improvements [643, 735]. Improving [232, 263, 326, 361, 612, 769, 785, 814, 849]. In-Place [278]. Incomplete [164, 888]. Incompressible [520, 871]. Incremental [610]. Independent [257, 340, 416]. Index [371, 420, 731]. industrial [726]. Inference [547]. InfiniBand [739]. Information [363, 513, 518, 604, 631, 669, 734, 812, 845, 881, 893]. infrastructure [802]. Initialization [290, 501, 579]. Insight [787]. inspector [844, 855]. inspector-executor [844, 855]. inspired [842, 873]. Instance [486]. Instance-Specific [486]. Institute [381, 422]. Instruction [135, 137, 142, 359, 503, 602, 738, 904]. Instruction-Level [135, 137, 142, 904]. Instruction-Set [602]. Instrumentation [339]. Integer [11, 57, 420, 575]. Integer-Valued [575]. Integrated [608, 752, 761]. Integration [577, 878]. Intel [214, 268]. intelligent [812]. intensive [796]. Inter [514, 817]. inter-Grid [817]. Inter-Tier [514]. Interaction [315]. Interactions [124, 448, 803]. Interactive [648, 747]. Interconnection [4, 29, 315, 342, 502, 549, 603, 616, 619, 724, 755, 825, 840]. interconnections [753]. Interconnects [283, 728, 741]. Interface [227, 388, 429, 716, 834]. interfaces [871]. Interleaving [303]. Interlocking [688]. Intermediate [73]. Internet [634]. Interoperable [331]. Interpolation [524, 826]. Interprocedural [36]. Interval [461]. Intrinsic [348]. introducing [856]. Introduction [4, 32, 132, 135, 136, 183, 229, 304, 311, 328, 381, 422, 446, 479]. Intrusion [607]. Intrusion-Tolerant [607]. Invalidation [347]. Inverse [470, 557, 712, 764]. Inverses [469]. Investigation [622]. Ionizing [87]. IPBGA [881]. iPSC [125]. iPSC/860 [125]. IPSec [681]. irreducibility [836]. Irregular [197, 200, 322, 475, 543, 572, 664, 733, 773]. Ising [806]. Issue [135, 170, 210, 222, 236, 362, 519, 527, 556, 756, 872, 904]. Issues [52, 132, 310, 716, 831]. Itemsets [676]. iterations [793]. Iterative [261, 443, 519, 523, 637, 785, 857]. ITRA [514]. IV [589].

J [25]. Jacobi [286]. Java [516, 786]. job [814]. Jobs [624]. Join [56, 268, 390]. Jordan [325]. Journal [872, 904].

Kascic [25]. Katsevich [691]. key [810]. knapsack [791]. Known [376].

L [1]. Label [608]. laceability [860]. LACSI [656]. Lagrange [826]. lags [805]. Lambda [760]. Language [43, 85, 95, 250, 263, 285, 383, 395, 424, 448, 654]. Language-Level [263]. Large [8, 14, 298, 345, 411, 471, 472, 500, 648, 665, 714]. Large-Scale [471, 472, 648]. LARPBS [545, 865]. Last [344]. Latency [73, 205, 315, 375, 483, 508]. Lattice [128]. Layer [373, 653, 802]. Layout [660]. LCAN [758]. LCAN-mapped [758]. Learning [113, 617, 776]. Least [520, 542, 792, 887]. Least-Squares [520]. Letter [15, 64, 70, 86, 87]. Level [70, 112, 135, 137, 142, 145, 175, 263, 300, 307, 326, 342, 383, 417, 424, 483, 689, 716, 751, 904]. Lexicon [821]. Liana [879]. libraries [709]. Library [525, 629]. lieu [71]. lifetime [868]. Lifting [341]. Limited [78]. limits [796].

Line [261, 606, 760, 848]. lineality [864]. Linear [14, 76, 84, 92, 180, 240, 289, 296, 298, 371, 413, 434, 445, 474, 510, 551, 725, 804, 843, 865]. Link [238, 438, 752, 838, 883]. link/CPU [752]. LINPACK [70]. Linux [472, 709, 899]. Linux-Cluster [472]. Lists [43]. LIW [68]. Load [158, 246, 291, 295, 463, 504, 529, 530, 581, 637, 755, 760, 795, 840, 879, 887]. Load-Balancing [291]. load-dependent [879]. loaded [887]. Loads [320]. Local [91, 437, 713, 857]. Locality [338, 391, 507, 553, 849]. localization [815]. Locating [265]. location [775, 847]. Log [447, 640]. Log-Based [640]. Logging [302, 416]. Logic [187]. Logical [71, 310]. long [698, 797, 823]. long-tailed [698, 797, 823]. Longest [260, 586]. Loop [264, 371, 488, 507, 553, 561, 570, 633, 685, 754, 785, 793, 903]. Loops [103, 158, 180, 237, 391, 457, 614, 655]. Loss [377]. Low [73, 286, 302, 378, 490, 571, 638, 832]. Low-Cost [378]. Low-Latency [73]. LU [293]. Lyapunov [269].

M [117, 239]. Machine [67, 146, 150, 244, 496]. Machine-2 [150]. Machines [41, 132, 149, 156, 183, 258, 303, 510, 613]. MAGNeT [386, 427]. makespan [759]. Mammalian [230]. manage [759]. manageable [872]. Management [238, 263, 450, 452, 462, 533, 573, 608, 671, 696, 704, 766, 810, 839, 852, 882]. Managing [775, 803, 867]. mapped [758]. Mapper [212]. Mapping [26, 131, 180, 200, 321, 340, 354, 395, 398, 417, 567, 687, 710, 815, 837, 863]. Mappings [244]. Maps [220]. Marginal [410]. Market [652]. Market-Oriented [652]. Markov [410]. MARS [117]. MARS-M [117]. Massively [102, 130, 161, 177, 256, 314]. master [814]. Match [821]. Matching [291, 354, 747, 764].

Materials [384, 425, 522]. mathematical [805, 889]. mathematics [726]. Matrices [459, 618]. Matrix [26, 154, 240, 330, 369, 444, 470, 473, 541, 613, 712, 752, 849, 886, 894]. Matrix-based [849]. matrix-vector [886, 894]. Maximal [676]. maximum [826]. Maze [28]. MCGLS [542]. MCGS [289]. MCHF [157]. Means [447, 487, 707]. Measurement [641, 741]. Mechanically [522]. Mechanism [385, 426, 735, 807]. Mechanisms [374, 766]. Medicine [229]. Medium [239, 529]. Meiko [191]. Membership [409]. Memoriam [94, 211]. Memories [73, 355, 578]. Memory [33, 73, 74, 91, 115, 119, 144, 146, 157, 162, 167, 173, 175, 198, 202, 224, 225, 230, 250, 258, 259, 275, 283, 291, 299, 302, 303, 316, 324, 326, 329, 334, 344, 380, 400, 421, 442, 476, 500, 512, 530, 542, 613, 632, 640, 679, 711, 721, 744, 796, 846]. merging [781]. Mesh [247, 282, 317, 318, 492, 528, 536, 626, 715, 813]. Mesh-Connected [318]. Meshes [62, 98, 478, 680]. Message [169, 197, 226, 238, 269, 275, 287, 332, 416, 560, 573, 577, 703, 709, 828]. Message-Passing [169, 197, 287, 560]. messaging [761]. Meta [447, 624]. Meta-coordination [447]. Meta-Schedulers [624]. Metacomputing [224, 431]. Metareasoning [636]. Metastability [313]. Method [230, 252, 258, 261, 270, 291, 322, 413, 461, 682, 687, 688, 706, 722, 832, 834, 871, 891]. methodologies [767]. Methodology [505, 568, 684, 690]. Methods [255, 319, 344, 401, 467, 542, 558, 584, 622, 695, 785, 795, 836, 843]. metric [792]. Metrics [67, 411]. MIB [606]. Michael [25]. Microcontroller [305]. Microkernels [171]. microprocessor [718]. Microscopic [196]. Microscopy [124]. Microtasking [47]. migration [776]. MIMD [111, 123, 146, 153, 213, 218, 283, 323]. MIN [159]. MIN-Based [159]. Minimal [270]. Minimalist [305]. Minimize [149].

Minimizing [282, 306]. Minimum [114, 486]. Mining [481, 676, 719]. Minisupercomputer [139]. Minkowski [411]. mismatch [858]. misses [794]. Mobile [320, 356, 389, 439, 495, 544, 573, 579, 847, 901]. Mode [365, 798, 805]. Model [40, 144, 202, 317, 329, 343, 357, 414, 432, 451, 454, 464, 576, 596, 608, 625, 626, 669, 686, 698, 734, 797, 805, 818, 823, 834, 839, 842, 865, 902]. Modeling [124, 129, 214, 279, 441, 451, 494, 508, 536, 617, 689, 697, 724, 726, 790, 827]. Modelling [271, 627, 644]. Models [335, 375, 449, 499, 550, 637, 806, 856]. modern [894]. Modified [289]. Modular [408]. Modularization [587]. Molecular [69, 232]. Monitor [193]. Monitoring [262, 481, 623, 737]. Monte [69, 107, 126, 213, 335, 413, 806]. morphological [727]. most [887]. moving [775]. MP [16, 47, 79, 104, 112, 118, 122, 128, 151]. MP/24 [22]. MP/416 [80]. MP/Y [118, 151]. MPI [367, 471, 627, 777, 899]. MPI-Based [471]. MPLD [185]. MRI [234]. Multi [48, 206, 249, 288, 319, 342, 379, 417, 575, 724, 791, 805]. multi-algorithm [48]. multi-cluster [724]. Multi-Dimensional [249, 288, 319, 379, 575, 791]. Multi-Level [342, 417]. multi-mode [805]. Multi-Port [379]. Multi-Ring [206]. Multicast [317, 336, 347, 453, 538, 601, 649, 728, 739, 773]. Multicasting [549, 572, 616]. Multicomputer [28, 155, 401]. Multicomputers [49, 81, 180, 281, 291, 344, 380, 400, 421, 632, 721, 744, 870]. Multidatabase [454]. Multidimensional [83, 203]. Multiflow [138, 142]. multilevel [804]. Multilinear [106]. Multimedia [385, 426, 585, 779, 812]. Multipath [549, 809]. Multipath-Based [549]. Multiphase [345]. Multiple [538, 551, 554, 624, 638, 767, 791, 838]. multiple-choice [791]. Multiplication [240, 473, 886, 894]. Multiprecise [57]. Multiprocessing [214]. Multiprocessor [115, 144, 172, 176, 241, 476, 498, 780, 791]. Multiprocessors [33, 159, 168, 198, 205, 218, 259, 269, 315, 316, 321, 323, 334, 338, 463, 618, 679]. multiresolution [699]. MultiRing [465]. Multistage [4, 603, 649]. Multitasking [58]. Multithreaded [482]. Multithreading [466, 577, 796, 851]. MUMPS [432]. Mutual [310, 609, 877].

nanoarchitectures [885]. natural [801]. Nautilus [654]. Navier [58, 156, 191, 218, 219, 323, 500]. navigation [884]. near [821]. near-duplicate [821]. Nearest [411]. Nearest-Neighbor [411]. Neighbor [411, 807]. Neighboring [98]. neocortex [842]. Nested [56, 180, 329, 391, 457, 655]. Nests [507]. NestStep [329]. Network [113, 206, 226, 227, 256, 313, 315, 338, 357, 365, 402, 462, 465, 483, 502, 538, 619, 649, 674, 698, 713, 734, 736, 741, 751, 755, 765, 797, 823, 828, 838, 869]. Networked [396]. Networks [4, 53, 285, 317, 320, 342, 362, 376, 379, 401, 433, 439, 440, 484, 537, 547, 549, 554, 572, 579, 603, 609, 616, 629, 651, 664, 668, 670, 678, 700, 705, 720, 724, 732, 739, 773, 779, 783, 802, 803, 807, 810, 813, 816, 825, 829, 840, 847, 887, 901, 902]. Neural [113, 433, 720]. Neutron [129, 552]. Newtonian [871]. no [85]. Node [387, 428, 661, 775]. Nodes [155]. noninterruptible [851]. Nonlinear [433, 529, 584, 874]. nonparametric [867]. nonserial [717]. Nonsymmetric [289]. Nonuniform [202, 603]. Normal [16, 97]. Normalized [557, 843]. Notion [21]. Novel [649, 666, 669, 846]. NT [367]. NUMA [259, 316, 338]. Number [16, 179]. numbers [693]. Numerical [8, 96, 346, 407, 726]. NuMesh [208].

O [203, 308, 664, 746, 751, 818, 831, 835, 844, 855, 896]. O-Conscious [403]. Object [264, 267, 285, 364, 504, 565, 669, 736]. Object-Oriented [264, 285]. Objects [349, 396, 399, 562, 775]. oblivious [711, 717]. Observations [144]. Ocean [144, 214, 432]. OCSP [771]. ODE [358]. OFDM [809]. OFDMA [850]. Off [261]. Off-Line [261]. Offs [244, 276]. OGRO [771]. oilfield [747]. Oldroyd [407]. Oldroyd-B [407]. on-chip [825]. On-Line [261, 760, 848]. one [48, 661, 902]. one-dimensional [48]. Online [234, 433, 776]. only [661]. onto [144, 261, 567, 568]. OpenMP [388, 418, 429, 531, 740, 843]. OpenRTE [762]. Operating [170, 172, 176]. operation [902]. Operations [43, 154, 252, 287]. opportunistic [779]. Opportunities [266]. Optical [212, 283, 311–313, 458, 499, 651, 695, 765, 865]. Optimal [120, 169, 280, 281, 392, 459, 460, 558, 645, 765, 773, 886, 890]. Optimization [79, 155, 301, 322, 339, 353, 391, 471, 477, 519, 522, 524, 548, 595, 835, 873, 879]. Optimized [208, 729]. Optimizing [34, 382, 423, 664, 772, 868]. option [848]. Optoelectronic [314]. Order [76, 433, 510, 596, 689]. Ordering [459]. Organisation [368]. Organization [8, 631, 882]. Organized [283]. Oriented [264, 285, 336, 459, 652, 830]. Oscillatory [313, 358]. Other [75, 365]. OTIS [840]. OTIS-Hypercube [840]. Otorhinolaryngological [472]. Overhead [286, 302, 306]. overheads [741]. Overlapped [306]. Overlapping [321]. Overlapping-Cluster [321]. overload [867]. Overview [137].

p [865]. P2P [777, 869, 873, 881]. P2P-MPI [777]. Package [57, 157]. Packet [377, 385, 426]. PACT [489]. Paging [198]. pairs [725]. Panconnectivity [768, 898]. pancycle [898]. pancycle-connectivity [898]. pancyclicity [768]. Panda [174, 203]. Papers [381, 422, 656]. paradigm [844, 855]. Paradigms [197]. Paragon [214, 268]. Parallel [13, 15, 29, 34, 35, 75, 77, 78, 105– 107, 114, 120, 125, 129, 130, 132, 133, 155, 156, 167, 174, 187, 195, 196, 201, 206, 210, 215, 218, 219, 230, 232, 234, 244, 246, 251, 252, 256, 261, 262, 266, 268, 269, 271, 280, 291–293, 298, 314, 318, 321, 324, 328, 333, 335, 341, 346, 348, 356, 361, 364, 366, 369, 389, 390, 420, 421, 459, 460, 464, 466, 475, 510, 511, 521, 523, 533, 536, 539, 540, 545, 546, 556, 557, 560, 576, 598, 612, 637, 641, 646, 651, 689, 691, 707, 719, 781, 826, 903]. Parallel [11, 43, 53, 98, 102, 123, 124, 126, 137, 161, 169, 170, 177, 191, 200, 213, 245, 247, 249, 270, 278, 288, 290, 299, 303, 306, 325, 327, 344, 358, 406, 408, 432, 444, 465, 471, 473, 500, 506, 520, 525, 535, 542, 552, 558, 559, 581, 584, 596, 633, 639, 644, 664, 665, 694, 708, 712, 715, 727, 740, 743, 747, 752, 754, 767, 777, 778, 790, 804, 835– 837, 843, 853, 856, 859, 865, 870, 885]. Parallelism [40, 56, 95, 135, 163, 258, 329, 785, 849, 857, 904]. Parallelization [36, 103, 197, 237, 264, 323, 330, 516, 548, 655, 861]. Parallelizing [254, 300, 303, 532, 543, 700, 733, 806, 819, 820]. Parameterizable [355]. Parameterized [357, 399, 865]. PARAMICS [196]. Part [151]. Partial [47, 298, 742]. Particle [150, 213, 521, 626]. Particle-in-Cell [150, 213]. Particular [261]. Partition [400]. partitioned [765]. Partitioning [180, 253, 300, 322, 417, 421, 512, 528, 578, 618, 638, 643, 690, 693, 701, 704, 793]. Pascal [85]. Passing [169, 197, 226, 238, 269, 275, 287, 332, 560, 577, 709, 828]. passive [765]. Path [461, 547, 616, 718, 884, 896]. Path-Based [547]. Pathology [233]. Paths [247, 619, 670, 725]. Pattern [603]. Patterns [295, 856]. PC [250, 363, 665, 793, 899]. PC-Based [250]. PDE [107, 154]. Peer [668, 732, 803, 858]. Peer-exchange [858]. peer-to-peer [668, 732, 803, 858]. pen [834]. pen-based [834]. Penalty [282]. Perfect [463, 537, 708, 755]. Performance [5, 7, 20, 22, 34, 48, 80, 91, 107, 117, 118, 122, 129, 133, 145, 178, 193, 202, 214, 215, 222, 226, 227, 236, 241, 252, 254, 262, 270, 273, 274, 276, 277, 279, 294, 299, 332, 345, 360, 364– 366, 369, 381, 387, 388, 412, 414, 422, 428, 429, 437, 444, 445, 471, 479–482, 523, 527, 531, 539, 559, 576, 603, 605, 612, 621–623, 625– 627, 634, 639, 643, 644, 646, 648, 653, 682, 686, 706, 716, 718, 728, 730, 746, 751, 753, 764, 782, 796, 808, 831, 859, 894]. performance [9, 151, 656, 662, 703, 737, 754, 779, 780, 802, 827, 835, 879, 896]. Performance-based [662, 754, 859]. performance-cost [879]. Period [645]. Permutation [579]. Permutations [618]. Perspective [137]. Perturbation [153, 433]. Perturbations [641]. Perturbed [506]. Pervasive [573]. Petaflops [502]. Phenomenon [146, 869]. PHR [581]. Phylogenetics [412]. Physical [155, 220]. physics [86]. Pipeline [576, 601, 866]. Pipelined [6, 20, 27, 41, 103, 240, 296, 314, 434, 551, 614, 865]. Place [278]. placement [664, 772, 875]. Placements [537]. platform [725, 727]. Platforms [219, 234, 261, 590, 623, 764]. Point [154, 506]. Poisson [16]. Policy [682, 882, 887]. Pollution [531]. polyadic [717]. Polygon [392]. Polynomial [52, 59, 575]. Polynomial-Time [575]. Polypeptides [126]. pool [881]. popularity [669]. Port [379, 902]. Portable [308]. Porting [144]. Postal [357]. Power [411, 490, 571, 638, 779, 802]. PPM [779, 802, 829]. Practical [713]. PRAMs [296]. pre [700]. pre-compiler [700]. Preallocation [158]. Precise [391]. Precision [188]. Preconditioners [444]. Preconditioning [469, 557, 562, 764]. Preconditionings [59]. predecessor [799]. Predicting [252, 276]. Prediction [161, 274, 332, 393, 523, 622, 707, 740, 827, 874]. predictions [817]. Predictor [553].

Prefetch [677]. Prefetching [503, 818, 846]. Prefix [78, 280, 291, 460, 551, 572, 778]. Presence [414, 647]. preventing [705]. Preventive [608]. Price [276]. Price/Performance [276]. Pricing [482, 848]. principle [792]. priority [120, 723]. PRISM [608]. Probabilistic [296, 817]. Probe [374]. Probing [439]. Problem [69, 114, 257, 260, 310, 404, 411, 477, 586, 708, 742, 745, 791, 805, 840]. Problems [200, 358, 392, 470, 506, 520, 542, 694, 726]. Process [340, 410, 637, 874]. Process-to-Processor [340]. processes [745]. Processing [137, 210, 252, 361, 489, 510, 540, 556, 576, 612, 665, 727, 885]. Processor [158, 184, 312, 340, 352, 457, 477, 487, 488, 586, 596, 602, 687, 689, 798, 815, 824, 849, 851]. Processors [6, 53, 474, 593, 736, 767, 811]. Procurement [550]. Produced [596, 689]. Profile [866]. Profile-based [866]. Program [36, 90, 104, 232, 328, 333, 334]. Programmable [84, 188, 811]. Programming [43, 44, 77, 106, 174, 202, 224, 248, 395, 418, 637, 646, 654, 717, 786]. Programs [33–35, 153, 169, 231, 262, 324, 327, 332, 400, 464, 516, 535, 627, 792]. project [762, 805]. Projections [318]. Prolog [35]. Propagating [529]. Properties [553]. Prospects [86, 283]. protection [832]. Protein [67, 630, 740]. Protocol [159, 239, 376, 377, 402, 442, 579, 668, 741, 783, 881]. Protocols [309, 648, 692, 703, 897, 901]. Prototype [388, 429, 688]. Prototyping [658, 689]. Providers [652]. Provisioning [585]. Proxies [642]. Pthread [740]. PTOOL [77]. Publication [652]. Publisher [39, 46]. Pulses [529]. Purpose [47, 895]. Pyramid [674].

QoS [514, 585, 868, 875]. QoS-Aware [585, 875]. QR [63, 130]. Quadrant [688]. Quadratic [13, 371]. Quality [361, 808]. Quantitative [195, 509]. Quantization [113]. quantum [128]. Quasi [270]. Quasi-Minimal [270]. Queue [120, 476, 596, 689]. quirk [71].

radiation [87]. Radiosity [581]. Radix [299]. Radix- [299]. RAID0.5 [832]. ramifications [818]. Random [16, 97]. randomization [821]. Randomized [410, 659]. Ranking [461]. Rapid [494, 658]. Rate [639, 794, 802]. Ray [246, 536]. Ray-Tracing [536]. Rayleigh [69]. RCC [566]. RCC-Based [566]. Reading [339]. Real [162, 239, 467, 498, 533, 539, 546, 568, 625, 696, 789]. Real-Time [239, 467, 498, 533, 539, 546, 568, 625, 696, 789]. Realignment [613]. Reality [612]. realization [725]. Realizing [651]. Receive [554]. Recognition [261, 562]. Reconfigurable [68, 184, 187, 206, 240, 296, 349, 352, 354– 356, 389, 394– 397, 434, 445, 458, 465, 489, 490, 492–495, 551, 565, 567, 568, 593, 690, 729, 769, 778, 862, 865]. Reconfigurations [353]. Reconstruct [251]. Reconstruction [220, 691, 857]. Recovery [302, 360, 640]. Recurrence [510]. Recurrences [76]. Recursions [237]. Recursive [357]. Redistribution [255, 282, 319, 380, 419, 541, 543, 645, 687, 815, 837]. redistributions [733, 828]. Reducing [333, 417, 608, 752, 794]. Reduction [492, 738]. Reductions [303, 383, 424]. redundancy [787]. References [371]. Refinement [528, 715]. Region [201, 465]. Registers [326]. registry [761]. Regular [256, 651]. reinforcement [776]. related [835]. relational [800]. Relationship [162, 514]. Relationships [336]. relative [669]. Relaxed [413]. Release [442]. Reliability [87, 625, 751, 816]. reliable [661, 703]. Remapping [541]. Remappings [261]. remote [748]. remoting [736]. Renaming [359]. Rendering [292, 344, 421, 721]. repeated [738].

Replacement [642]. Replay [169]. replica [875]. replicas [897]. Replicated [265, 501]. replication [822, 897]. report [15]. Reproducibility [346]. reprogrammable [725]. Request [267]. rerouting [705]. research [749]. Resident [403, 660]. Residual [270, 868]. resilient [762]. Resolution [119, 723]. Resource [238, 263, 511, 585, 696, 704, 734, 759, 766, 801, 805, 814, 817, 839, 845, 873, 878, 882]. resource-constrained [805]. Resources [635, 769]. Response [384, 425]. restriction [781]. Restructuring [90]. Results [324, 505]. Retargetable [602]. Retransmission [438]. Retransmissions [437]. Retrieval [233, 363, 669, 870]. Reuses [326]. Reversal [414]. review [17]. Reviewers [880]. Rhombic [321]. Ring [206, 281]. Rings [310]. Risk [608]. Risk-Reducing [608]. RLC [641]. RNA [161]. RNAi [774]. Road [196]. Robust [258, 788, 809]. Robustness [306, 759]. Rollback [640]. Root [188]. Routed [317, 478, 508, 549, 616]. Router [601]. routines [70]. Routing [28, 342, 402, 404, 414, 462, 572, 579, 647, 670, 680, 705, 753, 838, 901]. Row [404, 582, 710]. rules [719]. Run [175, 328, 330, 331, 397, 493, 516]. Run-Time [175, 328, 330, 331, 397, 493, 516]. Runtime [282, 723, 786]. Russian [117].

Sample [277]. SAN [279]. Sandbox [623]. Sandbox-Based [623]. Sandwich [207]. Satellite [524]. Scalability [648, 788, 852]. Scalable [168, 176, 218, 262, 502, 525, 667, 715, 786, 800, 810]. Scalar [73]. Scalarization [570]. Scale [345, 471, 472, 500, 648, 665]. scaleable [891]. Scaling [232, 752, 842, 866]. Scenario [834]. Scenario-focused [834]. Scheduled [208]. scheduler [879]. Schedulers [624]. Schedules [614]. Scheduling [42, 111, 138, 258, 316, 359, 366, 419, 505, 513, 561, 633, 638, 682, 687, 713, 728, 730, 733, 750, 754, 760, 767, 773, 779, 780, 799, 805, 813, 814, 817, 827, 829, 902, 903]. Scheme [278, 302, 320, 341, 351, 530, 633, 721, 723, 784, 846, 850]. Schemes [29, 416, 421, 605, 632, 638, 705, 740, 744, 787, 858]. Schrödinger [584]. SCI [363]. SCI-Based [363]. Science [381, 422, 468, 656]. Scientific [79, 195, 468, 519, 521, 527, 658, 748, 795, 876]. SCMP [560]. SCS [79]. SCS-40 [79]. SDD [22]. Search [126, 525, 597, 659, 662, 668, 731, 774, 800]. Second [179]. Secondary [161, 740]. Section [479]. Sectored [241]. Secure [368, 609, 681, 809, 812, 872]. Security [608, 653, 767]. Segment [823]. Segment-based [823]. Segmentation [201, 465]. Segmented [464]. Seismic [536]. Selection [443, 509]. Selective [373]. Self [420, 489, 633, 647]. Self-Index [420]. Self-Reconfigurable [489]. Self-Scheduling [633]. Self-Similar [647]. Semantic [547]. Semaphore [173]. semiconductor [9]. Send [554]. Sensitivity [332]. sensor [783, 810, 847]. SEP [256]. Separation [247]. Sequence [67, 231]. sequences [659, 738]. Sequential [324]. Serial [114, 655]. Server [501, 607, 664, 772, 893, 896]. ServerNet(R) [415]. ServerNet^{T M} [279]. Service [604, 631, 652, 704, 758, 812, 839, 876, 879]. Services [607, 652, 748, 872, 877]. Sessions [452]. Set [114, 145, 475, 578, 602]. Sets [403, 411, 743]. Several [392]. SFS [524]. SHA [580]. SHA-2 [580]. Shanks [211]. Shared [115, 144, 146, 167, 173, 198, 224, 225, 230, 250, 259, 275, 283, 302, 324, 329, 334, 442, 452, 476, 512, 530, 577, 640, 670, 675, 679, 846]. Shared- [167]. Shared-Memory [115, 144, 146, 173, 230, 283, 476]. sharing [817]. Shear [421]. Shear-Warp [421]. Shielding [438]. Shift [457]. Shift-Variant [457]. Shortest [247, 725, 884]. SI [897]. Side [35]. Sidney [94]. Sieve [13]. Signed [461]. Signed-Distance [461]. SIMD [149, 318]. Similar [647]. Simple [629]. Simplex [107, 891]. Simulated [404, 477]. Simulating [98, 544]. Simulation [69, 124, 196, 284, 309, 333, 377, 384, 407, 410, 425, 522, 531, 765, 767, 790]. Simulation-based [309]. Simulations [107, 213, 296, 345, 406, 532, 648, 806]. Simulator [602]. simultaneous [796]. Single [395, 404, 524, 560, 579, 582, 710]. Single-Chip [560]. Single-Hop [579]. Single-Row [404, 582, 710]. Singularly [506]. sinks [847]. Size [280, 460, 492, 509, 669, 738]. sized [693]. Sizes [443]. Skeletons [248]. Skewed [400]. Skewing [570]. slave [814]. Slotnick [1]. small [869]. smart [747]. Smooth [808, 871]. SMP [510, 577, 637]. SMPs [614]. snapshot [757]. SNMP [606]. SoC [496]. SoC-Architecture [496]. social [803]. sockets [716]. SoCs [643]. Software [47, 65, 110, 186, 273, 349, 356, 381, 389, 396, 422, 449, 530, 550, 587, 625, 640, 684, 701, 833, 848]. Solid [124, 680]. Solution [92, 291, 361, 470, 557, 617, 778, 871]. Solution-Adaptive [291]. Solutions [392, 575, 801]. Solve [310, 598, 804]. Solver [123, 156, 167, 191, 323, 358, 520]. Solvers [107, 218, 282, 523, 558]. Solving [14, 52, 129, 269, 413, 519, 619]. Some [16, 694]. Sort [49, 292, 344, 458]. Sort-First [292]. Sort-Last-Sparse [344]. Sorting [54, 215, 278, 420]. Source [245, 339]. Source-to-Source [339]. Space [292, 499, 576, 659, 729, 732, 792, 864, 886]. Space-Time [576]. Spacecraft [305]. Spaces [452]. Spacing [385, 426]. spanning [860]. Sparse [107, 130, 330, 344, 348, 459, 469, 541, 613, 618, 632, 744, 752, 843, 894]. Special [135, 170, 210, 222, 236, 362, 479, 519, 527, 556, 756, 872, 904]. Specific [486, 494]. Specification [351, 654, 722]. Spectral [520, 871]. Spectrum [277]. Speed [84, 355, 362, 629, 681, 741]. Speeds [554]. speedups [718]. Spherical [699]. spin [806, 885]. spin-wave [885]. Splash [186]. Split [584, 775]. Split-Step [584]. SPM [788]. SPMD [418]. Spot [508]. SPT [838]. Spyder [184]. Square [98, 188]. Squared [289]. Squares [520, 542]. Stability [346]. Stabilizing [298]. Stage [666]. Standard [580]. Star [549, 616, 755, 890]. stars [765]. Startup [709]. state [838]. stateful [736]. Statement [300]. Statement-Level [300]. Static [111, 237, 333]. Statistical [461]. Status [7, 86]. Steiner [742]. Stellar [150]. stencil [711]. Stencils [154]. Step [584]. Stereocorrelation [206]. Stochastic [393, 647]. Stokes [58, 156, 191, 218, 219, 323, 500]. storage [751, 776, 854, 875]. storage-efficient [875]. Strassen [92]. Strategies [246, 549, 617]. Strategy [403, 677, 708, 773, 822, 828, 892, 900]. stream [811, 849]. streaming [808, 896]. streamization [849]. striping [703]. Structure [157, 161, 176, 378, 476, 630, 659, 740]. Structured [145, 168, 528, 626, 715, 732, 869]. Structuring [263]. Studies [522, 788]. Study [67, 195, 201, 241, 244, 248, 284, 331, 347, 401, 441, 559, 606, 658, 692, 743, 780]. subchannel [850]. Subscripts [327, 371]. Subsequence [260, 586]. Subsurface [345]. Success [562]. Sum [551]. super [860]. Superblock [141]. Supercomputer [7, 9, 73, 91, 178]. Supercomputers [12, 65, 80, 97, 110, 118, 119, 122, 151, 231]. Supercomputing [31, 229, 726, 872, 904]. Superconductor [125]. Superlinear [539]. supernode [830]. Superscalar [141, 184]. Supply [638]. Support [173, 175, 259, 328, 330, 383, 424, 516, 656, 660, 667, 786, 895]. supported [739]. Supporting [262, 287, 317, 716]. Supports [348, 736]. SURE [184]. Surface [274]. Surfaces [251]. Survey [295, 673, 731]. Survivability [402]. Survivable [607, 670]. Suspension [447].

Sweep [552]. Sweep3D [657]. Swirling [406]. Switch [499, 664]. switch-based [664]. Switched [379, 484, 753]. Switching [401, 493, 736]. Symbolic [254, 475]. Symmetric [679, 804]. Symmetry [102]. Symposium [381, 422, 656]. Synchronization [360]. Synchronizer [313]. Synchronous [270, 393, 645, 870]. Synthesis [350, 457, 488, 886]. Synthetic [153]. Synthetic-Perturbation [153]. System [173, 176, 226, 240, 307, 320, 339, 344, 368, 402, 434, 472, 493, 513, 551, 581, 617, 662, 676, 814, 817, 845, 850, 865, 867, 868]. system-generated [817]. System-Level [307]. Systems [14, 52, 92, 170, 172, 236, 256, 263, 289, 296, 298, 302, 304, 308, 309, 331, 336, 354, 356, 381, 389, 394, 395, 397, 399, 413, 422, 442, 454, 467, 471, 476, 485, 491, 495, 501, 504, 511, 530, 533, 564, 587, 589, 598, 605, 621, 644, 645, 648, 667, 677, 684, 695, 697, 706, 718, 722, 724, 745, 746, 751, 776, 789, 799, 804, 809, 814, 831, 843, 846, 856, 858, 865, 873, 889]. Systolic [83, 260, 557, 688, 725, 886].

T3D [215]. Table [534]. tactics [803]. tag [705]. tailed [698, 797, 823]. Tailored [356, 389]. Target [261, 774]. Targeted [465]. Task [258, 466, 505, 780]. Tasks [261, 366, 492, 750, 789]. TCP [375, 437–439, 639, 784, 802]. TCP-Probing [439]. TDMA [813]. Technique [113, 141, 404, 506, 541, 687, 815]. Techniques [68, 76, 267, 295, 300, 301, 315, 328, 376, 400, 470, 490, 519, 522, 528, 870]. Technologies [365, 747, 756]. technology [9, 740, 749]. Template [354]. terminal [742]. Terms [31, 447]. Test [384, 425, 575, 722]. Testing [309, 455, 833, 836]. text [719, 870]. TH [779, 802, 829]. TH-PPM [779, 802, 829]. their [652]. Thematic [212]. Theoretic [434]. Theory [32, 433, 694]. threading [740]. Threads [175]. Three [156, 171, 213, 218, 323, 371, 630]. Three-Dimensional [213, 218, 323, 630].

Throughput [146, 375, 741, 814]. Tier [514]. Tile [443, 509]. Tiled [614, 825]. Tiling [403, 685]. Time [175, 177, 239, 240, 258, 328, 330, 331, 393, 397, 467, 493, 498, 516, 533, 539, 546, 568, 575, 576, 625, 696, 789, 805, 809]. time-critical [177]. time-domain [809]. time-varying [809]. Timing [467]. Toeplitz [598, 804]. Tolerance [287, 307, 310, 619, 784]. Tolerant [177, 304, 305, 308, 366, 378, 478, 498, 607, 680, 753, 856, 890, 900]. tolerate [892]. tolerating [705]. Tool [294, 340, 387, 388, 428, 429, 623]. Toolkit [386, 427, 737, 859]. Tools [77, 328, 331, 334]. Toolset [294]. Top [387, 428]. Top-down [387, 428]. Topological [458]. Topologies [279]. Topology [321, 637]. Tori [484]. Toroidal [537]. Torus [379, 828]. Toruses [98]. Touching [247]. Touchstone [166]. TRACE [138, 142]. Tracing [104, 169, 246, 536]. Tracking [115, 521, 565]. Trade [244, 276]. Trade-Offs [244, 276]. tradeoff [759]. Tradeoffs [198, 440]. trading [876]. Traffic [196, 239, 326, 414, 508, 603, 634, 647, 698, 797, 823]. Transaction [454, 533]. Transcoding [642]. Transcoding-Enabled [642]. Transfer [241, 365]. Transfers [377]. Transform [5, 474, 482, 545, 610, 811]. Transformation [303, 582, 660, 662, 710]. Transformations [192, 324]. Transforming [507]. Transforms [81, 83, 249, 341, 699]. Transient [440]. Transmission [639]. Transparent [679, 762]. Transport [129, 335, 373, 552, 626, 653, 703]. Tree [498, 610, 616, 739, 742]. Tree-Based [616]. Trees [453]. Trends [170, 172]. Triangular [282]. Tribology [532]. Tridiagonal [444, 558]. TRLE [721]. Truncated [410]. Trustworthy [748, 833]. Tulip [650]. Tuned [145]. Tuning [153, 273, 662, 669]. tuple [732]. Turbulent [406]. Turn [317, 686]. Twenty [179]. Twenty-Second [179]. Two [197, 218, 506, 769]. Two- [218]. two-dimensional [769]. Two-Point [506]. Tycho [761]. Type [584, 603].

U [812]. U-home [812]. U-multimedia [812]. Ubiquitous [585, 809]. ULSI [87]. Ultrahigh [122]. Ultrahigh-performance [122]. Ultralightweight [175]. uncertain [868]. Unibus [763]. Unidirectional [597]. uniform [16]. Unimodular [192]. Unit [571]. Units [73, 690]. Universal [496]. Unroll [443]. unstable [874]. Unstructured [123, 167, 291, 292, 471, 552, 626]. update [838]. updates [847]. UpdateSearch [320]. Updating [330]. upon [606]. Use [73, 77, 261, 712]. User [175, 377, 383, 424, 716, 834]. User-defined [383, 424]. User-Level [175, 716]. Using [77, 91, 92, 133, 193, 220, 263, 271, 273, 310, 341, 346, 362, 399, 410, 421, 432, 433, 443, 445, 458, 466, 470, 472, 477, 492, 524, 526, 531, 553, 558, 570, 572, 580, 602, 608, 638, 639, 655, 659, 662, 722, 740, 758, 760, 769, 771, 800, 817, 818, 822, 827, 881]. Utility [271]. Utilization [511, 606, 629, 769, 814, 826]. utilizing [718]. UWB [779, 802, 829].

V [85]. V-Pascal [85]. Validation

[467, 602, 771]. Value [470, 506, 553]. Valued [575]. Variable [188, 575]. Variable-Precision [188]. Variables [97, 371]. Variant [457]. Various [536, 605]. varying [809]. Vector [75, 76, 79, 113, 118, 119, 151, 230, 284, 326, 886, 894]. Vectorized [16, 61, 69]. Vectorizing [85, 154]. vehicle [884]. Verification [351, 889]. version [11]. Vertical [293]. very [48]. VF [145]. Via [52, 481, 550]. Video [639, 713]. view [757]. Virtual [173, 185, 329, 384, 396, 425, 612, 631, 667, 761, 882]. virtualized [900]. Visibility [590]. Vision [245]. visual [737]. Visualisation [273]. VLIW [141, 142]. VoD [513]. voltage [752, 767]. Voltages [638]. Volume [292, 318, 344, 421, 534, 721].

Volunteer [337]. VPN [681]. VR [472]. VR-Based [472]. vs [275].

WAN [274]. WAN-Based [274]. Warp [421]. WASMII [185]. Watermarking [610]. Wave [374, 885]. Wavefront [269]. Wavelength [651]. Wavelet [249, 341, 610]. WDM [670]. web [337, 559, 650, 669, 714, 839]. WebCom [337]. Webs [223]. weights [822]. Well [376]. Well-Known [376]. WHILE [103]. Whitney [276]. whole [662]. Wide [365, 629, 761]. Wide-Area [365, 761]. Width [78]. Windows [367]. Wireless [376, 438, 440, 567, 609, 653, 673, 783, 807, 813, 829, 847]. without [640]. Work [451, 694]. Work-efficient [694]. Workbench [408]. workflow [845]. workflow-based [845]. working [669]. Workload [193, 282, 645, 671, 692, 827, 841]. Workloads [484]. workshop [15]. Workstation [366]. Workstations [220, 226, 266, 390, 538, 554, 700]. world [869]. Wormhole [297, 317, 414, 478, 484, 508, 549, 616, 753]. Wormhole-Routed [317, 478, 508, 549, 616]. Wormhole-Switched [484, 753]. Write [347].

X [16, 22, 47, 79, 80, 104, 118, 151]. X-MP [16, 22, 47, 79, 80, 104, 118, 151]. X-MP/24 [22]. X-MP/416 [80]. X-MP/Y-MP [118, 151]. XA [178]. XD1 [842]. XML [800]. XPP [489].

Y-MP [112, 118, 122, 128, 151]. YAARC [794]. yield [87].

References

Anonymous:1987:DDS

[1] Anonymous. Dedication: Daniel L. Slotnick, 1931 to 1985. The Journal of Supercomputing, 1(1):5–6, March 1987. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                             URL http://www.springerlink.com/
                                             openurl.asp?genre=article&issn=
                                             0920-8542&volume=1&issue=1&spage=
                                             5.

Anonymous:1987:E

[2] Anonymous. Editorial. The Journal of Supercomputing, 1(1):7–8, March 1987. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                             URL http://www.springerlink.com/
                                             openurl.asp?genre=article&issn=
                                             0920-8542&volume=1&issue=1&spage=
                                             7.

Anonymous:1987:CA

[3] Anonymous. Contributing authors. The Journal of Supercomputing, 1(1): 9–12, March 1987. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                             //www.springerlink.com/openurl.
                                             asp?genre=article&issn=0920-8542&
                                             volume=1&issue=1&spage=9.

Siegel:1987:IMC

[4] Howard Jay Siegel, William Tsun-Yuk Hsu, and Menkae Jeng. An introduction to the multistage cube family of interconnection networks. The Journal of Supercomputing, 1(1):13–42, March 1987. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                             URL http://www.springerlink.com/
                                             openurl.asp?genre=article&issn=
                                             0920-8542&volume=1&issue=1&spage=
                                             13.

Bailey:1987:HPF

[5] David H. Bailey. A high-performance fast Fourier transform algorithm for the Cray-2. The Journal of Supercomputing, 1(1):43–60, March 1987. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=1&spage=43.

Irwin:1987:DPP

[6] Mary Jane Irwin and Robert Michael Owens. Digit pipelined processors. The Journal of Supercomputing, 1 (1):61–86, March 1987. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=1&spage=61.

Martin:1987:SPE

[7] Joanne L. Martin and Dieter Mueller-Wichards. Supercomputer performance evaluation: Status and directions. The Journal of Supercomputing, 1(1):87–104, March 1987. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=1&spage=87.

Bell:1987:DOL

[8] J. L. Bell and G. S. Patterson, Jr. Data organization in large numerical computations. The Journal of Supercomputing, 1(1):105–136, March 1987. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=1&spage=105.

Osburn:1987:CAS

[9] Carlton M. Osburn and Arnold Reisman. Challenges in advanced semiconductor technology for high-performance and supercomputer applications. The Journal of Supercomputing, 1(2):149– 189, June 1987. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=2&spage=149.

Buell:1987:FAC

[10] Duncan A. Buell. Factoring: Algorithms, computations, and computers. The Journal of Supercomputing, 1(2):191–216, June 1987. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=2&spage=191.

Wunderlich:1987:PVC

[11] M. C. Wunderlich and H. C. Williams. A parallel version of the continued fraction integer factoring algorithm. The Journal of Supercomputing, 1(2):217–230, June 1987. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=1&issue=2&spage=
                                        217.

Resnikoff:1987:CEC

[12] H. L. Resnikoff. Cost-effectiveness of concurrent supercomputers. The Journal of Supercomputing, 1(2):231– 262, June 1987. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=2&spage=231.

Caron:1988:PIQ

[13] Thomas R. Caron and Robert D. Silverman. Parallel implementation of the quadratic sieve. The Journal of Supercomputing, 1(3):273–290, April 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=1&issue=3&spage=
273.

Grimes:1988:SSL

[14] Roger. G. Grimes. Solving systems of large dense linear equations. The Journal of Supercomputing, 1 (3):291–299, April 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=3&spage=291.

Buell:1988:LEP

[15] D. A. Buell, D. A. Carlson, Yuan chieh Chow, K. Culik, N. Deo, R. Finkel, E. N. Houstis, E. M. Jacobson, Z. M. Kedem, J. S. Kowalik, P. J. Kuekes, J. L. Martin, G. A. Michael, N. S. Ostlund, J. Potter, D. K. Pradhan, M. J. Quinn, G. W. Stewart, Q. F. Stout, L. Watson, and J. Webb. Letter to the Editor: Parallel algorithms and architectures: report of a workshop. The Journal of Supercomputing, 1(3):301–325, April 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=3&spage=301.

Petersen:1988:SVR

[16] W. P. Petersen. Some vectorized random number generators for uniform, normal, and Poisson distributions for CRAY X-MP. The Journal of Supercomputing, 1(3):327–335, April 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=3&spage=327.

Riganati:1988:BR

[17] John Riganati. Book review. The Journal of Supercomputing, 1(3):337– 339, April 1988. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=3&spage=337.

Anonymous:1988:CAa

[18] Anonymous. Contributing authors. The Journal of Supercomputing, 1 (3):341–342, April 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=3&spage=341.

Anonymous:1988:Ea

[19] Anonymous. Editorial. The Journal of Supercomputing, 1(3):??, April 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=

0920-8542&volume=1&issue=3&spage=
??.

Topham:1988:DPC

[20] Nigel P. Topham, Amos Omondi, and Roland N. Ibbett. On the design and performance of conventional pipelined architectures. The Journal of Supercomputing, 1(4):353–393, August 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=1&issue=4&spage=353.

Kruskal:1988:NG

[21] Clyde P. Kruskal and Carl H. Smith. On the notion of granularity. The Journal of Supercomputing, 1(4):395–408, August 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=1&issue=4&spage=
395.

Anderson:1988:PCC

[22] Richard E. Anderson, Roger G. Grimes, and Horst D. Simon. Performance comparison of the CRAY X-MP/24 with SDD and the CRAY-2. The Journal of Supercomputing, 1(4):409–419, August 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=1&issue=4&spage=
409.

Anonymous:1988:CAb

[23] Anonymous. Contributing authors. The Journal of Supercomputing, 1 (4):421–424, August 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=1&issue=4&spage=421.

Anonymous:1988:Eb

[24] Anonymous. Editorial. The Journal of Supercomputing, 1(4):??, August 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=1&issue=4&spage=
                                        ??.

Koehler:1988:DMJ

[25] James F. Koehler. Dedication: Michael J. Kascic, 1941–1987. The Journal of Supercomputing, 1(4):??, August 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=1&issue=4&spage=
                                        ??.

Cherkassky:1988:EMI

[26] Vladimir Cherkassky and Ross Smith. Efficient mapping and implementation of matrix algorithms on a hypercube. The Journal of Supercomputing, 2(1):7–27, September 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=1&spage=7.

Topham:1988:CFA

[27] Nigel P. Topham, Amos Omondi, and Roland N. Ibbett. Context flow: An alternative to conventional pipelined architectures. The Journal of Supercomputing, 2(1):29–53, September 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=2&issue=1&spage=
29.

Won:1988:MRH

[28] Youngju Won and Sartaj Sahni. Maze routing on a hypercube multicomputer. The Journal of Supercomputing, 2(1):55–79, September 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=1&spage=55.

Lakshmivarahan:1988:NHH

[29] S. Lakshmivarahan and Sudarshan K. Dhall. A new hierarchy of hypercube interconnection schemes for parallel computers. The Journal of Supercomputing, 2(1):81–108, September 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=1&spage=81.

Anonymous:1988:CAc

[30] Anonymous. Contributing authors. The Journal of Supercomputing, 2(1): 109–112, September 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=1&spage=109.

Anonymous:1988:GTS

[31] Anonymous. Glossary of terms (supercomputing). The Journal of Supercomputing, 2(1):113–123, September 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=2&issue=1&spage=
                                        113.

Banerjee:1988:IFT

[32] Utpal Banerjee. An introduction to a formal theory of dependence analysis. The Journal of Supercomputing, 2(2):133–149, October 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=2&spage=133.

Callahan:1988:CPD

[33] David Callahan and Ken Kennedy. Compiling programs for distributedmemory multiprocessors. The Journal of Supercomputing, 2(2):151–169, October 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=2&issue=2&spage=
                                        151.

Chen:1988:CPP

[34] Marina Chen, Young-Il Choo, and Jingke Li. Compiling parallel programs by optimizing performance. The Journal of Supercomputing, 2(2):171– 207, October 1988. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=2&spage=171.

Kale:1988:PEP

[35] Laxmikant V. Kale, David A. Padua, and David C. Sehr. OR parallel execution of Prolog programs with side effects. The Journal of Supercomputing, 2(2):209–223, October 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=2&spage=209.

Li:1988:PPI

[36] Zhiyuan Li and Pen-Chung Yew. Program parallelization with interprocedural analysis. The Journal of Supercomputing, 2(2):225–244, October 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=2&spage=225.

Anonymous:1988:CAd

[37] Anonymous. Contributing authors. The Journal of Supercomputing, 2(2): 245–250, October 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=2&spage=245.

Anonymous:1988:Ec

[38] Anonymous. Editorial. The Journal of Supercomputing, 2(2):??, October 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=2&issue=2&spage=
??.

Anonymous:1988:PAa

[39] Anonymous. Publisher’s announcement. The Journal of Supercomputing, 2(2):??, October 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=2&spage=??.

Mou:1988:AMD

[40] Zhijing G. Mou and Paul Hudak. An algebraic model for divide-and-conquer and its parallelism. The Journal of Supercomputing, 2(3):257–278, November 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=2&issue=3&spage=
                                        257.

Nicolau:1988:FGC

[41] Alexandru Nicolau, Keshav Pingali, and Alexander Aiken. Fine-grain compilation for pipelined machines. The Journal of Supercomputing, 2(3): 279–295, November 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=3&spage=279.

Polychronopoulos:1988:TAS

[42] Constantine D. Polychronopoulos. Toward auto-scheduling compilers. The Journal of Supercomputing, 2(3):297– 330, November 1988. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=2&issue=3&spage=297.

Solworth:1988:PLC

[43] Jon A. Solworth. Programming language constructs for highly parallel operations on lists. The Journal of Supercomputing, 2(3):331–347, November 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=2&issue=3&spage=
331.

Wu:1988:PAH

[44] Min-You Wu and Daniel D. Gajski. A programming aid for hypercube architectures. The Journal of Supercomputing, 2(3):349–372, November 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=3&spage=349.

Anonymous:1988:CAe

[45] Anonymous. Contributing authors. The Journal of Supercomputing, 2(3): 373–376, November 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=3&spage=373.

Anonymous:1988:PAb

[46] Anonymous. Publisher’s announcement. The Journal of Supercomputing, 2(3):??, November 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=3&spage=??.

Bieterman:1988:MGP

[47] Michael Bieterman. Microtasking general purpose partial differential equation software on the CRAY X-MP. The Journal of Supercomputing, 2(4): 381–413, December 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=4&spage=381.

Armstrong:1988:MAA

[48] Jim Armstrong. A multi-algorithm approach to very high performance onedimensional FFTs. The Journal of Supercomputing, 2(4):415–433, December 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=2&issue=4&spage=
                                        415.

Won:1988:BSH

[49] Youngju Won and Sartaj Sahni. A balanced bin sort for hypercube multicomputers. The Journal of Supercomputing, 2(4):435–448, December 1988. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=2&issue=4&spage=435.

Anonymous:1988:CAf

[50] Anonymous. Contributing authors. The Journal of Supercomputing, 2(4): 449–450, December 1988. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=2&issue=4&spage=449.

Anonymous:1989:E

[51] Anonymous. Editorial. The Journal of Supercomputing, 3(1):3, March 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=3&issue=1&spage=
3.

Allison:1989:GIS

[52] D. C. S. Allison, Amal Chakraborty, and Layne T. Watson. Granularity issues for solving polynomial systems via globally convergent algorithms on a hypercube. The Journal of Supercomputing, 3(1):5–20, March 1989. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=1&spage=5.

Oruc:1989:CNC

[53] A. Yavuz Oruç and Seth Schneider. Coset networks as connectors in parallel processors. The Journal of Supercomputing, 3(1):21–39, March 1989. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=1&spage=21.

Won:1989:HHS

[54] Youngju Won and Sartaj Sahni. Hypercube-to-host sorting. The Journal of Supercomputing, 3(1):41–61, March 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=3&issue=1&spage=
                                        41.

Anonymous:1989:CAa

[55] Anonymous. Contributing authors. The Journal of Supercomputing, 3 (1):63–65, March 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=1&spage=63.

Burke:1989:AGN

[56] Michael Burke, Ron Cytron, Jeanne Ferrante, and Wilson Hsieh. Automatic generation of nested, fork-join parallelism. The Journal of Supercomputing, 3(2):71–88, July 1989. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=2&spage=71.

Buell:1989:MIA

[57] Duncan A. Buell and Robert L. Ward. A multiprecise integer arithmetic package. The Journal of Supercomputing, 3(2):89–107, July 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=2&spage=89.

Fatoohi:1989:MNS

[58] R. A. Fatoohi. Multitasking a Navier– Stokes algorithm on the CRAY-2.

The Journal of Supercomputing, 3 (2):109–124, July 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=2&spage=109.

Wong:1989:APP

[59] Yau Shu Wong and Hong Jiang. Approximate polynomial preconditionings applied to biharmonic equations. The Journal of Supercomputing, 3 (2):125–145, July 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=2&spage=125.

Anonymous:1989:CAb

[60] Anonymous. Contributing authors. The Journal of Supercomputing, 3 (2):147–148, July 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=2&spage=147.

Appel:1989:VGC

[61] Andrew W. Appel and Aage Bendiksen. Vectorized garbage collection. The Journal of Supercomputing, 3(3): 151–160, September 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=3&spage=151.

Boxer:1989:DCG

[62] Laurence Boxer and Russ Miller. Dynamic computational geometry on meshes and hypercubes. The Journal of Supercomputing, 3(3):161–191, September 1989. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=3&spage=161.

Bischof:1989:ABQ

[63] Christian H. Bischof. Adaptive blocking in the QR factorization. The Journal of Supercomputing, 3(3):193– 208, September 1989. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=3&spage=193.

Woo:1989:LEH

[64] Jinwoon Woo and Sartaj Sahni. Letter to the Editor: Hypercube computing: Connected components. The Journal of Supercomputing, 3(3):209– 234, September 1989. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=3&spage=209.

Ranai:1989:CAB

[65] K. Ranai and B. Srinivasan. Complexity analysis of benchmark software for supercomputers. The Journal of Supercomputing, 3(3):235–242, September 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=3&issue=3&spage=
235.

Anonymous:1989:CAc

[66] Anonymous. Contributing authors. The Journal of Supercomputing, 3(3): 243–245, September 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=3&spage=243.

Lander:1989:SPS

[67] Eric Lander, Jill P. Mesirov, and Washington Taylor, IV. Study of protein sequence comparison metrics on the Connection Machine CM-2. The Journal of Supercomputing, 3(4): 255–269, December 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=3&issue=4&spage=255.

Gupta:1989:CTR

[68] Rajiv Gupta and Mary Lou Soffa. Compilation techniques for a reconfigurable LIW architecture. The Journal of Supercomputing, 3(4):271–304, December 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=3&issue=4&spage=
271.

Pryor:1989:VMC

[69] Daniel V. Pryor and Patrick J. Burns. Vectorized Monte Carlo molecular aerodynamics simulation of the Rayleigh problem. The Journal of Supercomputing, 3(4):305–330, December 1989. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=4&spage=305.

Louter-Nool:1989:LEL

[70] Margreet Louter-Nool. Letter to the Editor: LINPACK routines based on level 2 BLAS. The Journal of Supercomputing, 3(4):331–349 (or 331–339??), December 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=3&issue=4&spage=
                                        331.

Lee:1989:QCC

[71] Jon Lee. A quirk of the CRAY CFT77 compiler: IF (logical) in lieu of IF (arithmetic). The Journal of Supercomputing, 3(4):351–355, December 1989. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=4&spage=351.

Anonymous:1989:CAd

[72] Anonymous. Contributing authors. The Journal of Supercomputing, 3(4): 357–359, December 1989. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=3&issue=4&spage=357.

Sohi:1990:UIM

[73] Gurindar S. Sohi and Wei-Chung Hsu. The use of intermediate memories for low-latency memory access in supercomputer scalar units. The Journal of Supercomputing, 4(1):5–21, March 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=4&issue=1&spage=5.

Bailey:1990:FEH

[74] David H. Bailey. FFTs in external or hierarchical memory. The Journal of Supercomputing, 4(1):23–35, March 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=1&spage=
23.

Guzzi:1990:CFO

[75] Mark D. Guzzi, David A. Padua, Jay Hoeflinger, and Duncan H. Lawrie. Cedar Fortran and other vector and parallel Fortran dialects. The Journal of Supercomputing, 4(1):37–62, March 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=1&spage=
37.

Tanaka:1990:CTF

[76] Yoshikazu Tanaka, Kyouko Iwasawa, Yukio Umetani, and Shizuo Gotou. Compiling techniques for first-order linear recurrences on a vector computer. The Journal of Supercomputing, 4(1):63–82, March 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=1&spage=63.

Henderson:1990:UDD

[77] Leslie Ann Henderson, Robert E. Hiromoto, Olaf M. Lubeck, and Margaret L. Simmons. On the use of diagnostic dependence-analysis tools in parallel programming: Experiences using PTOOL. The Journal of Supercomputing, 4(1):83–96, March 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=1&spage=83.

Carlson:1990:LWP

[78] David A. Carlson and Binay Sugla. Limited width parallel prefix circuits. The Journal of Supercomputing, 4 (2):107–129, June 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=2&spage=107.

Pfeiffer:1990:BOS

[79] Wayne Pfeiffer, Arnold Alagar, Anke Kamrath, Robert H. Leary, and Jack Rogers. Benchmarking and optimization of scientific codes on the CRAY X-MP, CRAY-2, and SCS-40 vector computers. The Journal of Supercomputing, 4(2):131–152, June 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=2&spage=131.

Simmons:1990:PCC

[80] Margaret L. Simmons and Harvey J. Wasserman. Performance comparison of the CRAY-2 and CRAY X-MP/416 supercomputers. The Journal of Supercomputing, 4(2):153–167, June 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=2&spage=
153.

Ranka:1990:CHT

[81] Sanjay Ranka and Sartaj Sahni. Computing Hough transforms on hypercube multicomputers. The Journal of Supercomputing, 4(2):169–190, June 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=2&spage=
169.

Anonymous:1990:CAa

[82] Anonymous. Contributing authors. The Journal of Supercomputing, 4 (2):191–194, June 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=4&issue=2&spage=191.

Shen:1990:SAM

[83] Weicheng Shen and A. Yavuz Oruç. Systolic arrays for multidimensional discrete transforms. The Journal of Supercomputing, 4(3):201–222, September 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=3&spage=
201.

Lee:1990:HSC

[84] Peizong Lee and Zvi M. Kedem. On high-speed computing with a programmable linear array. The Journal of Supercomputing, 4(3):223– 249, September 1990. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=4&issue=3&spage=223.

Tsuda:1990:VPA

[85] Takao Tsuda and Yoshitoshi Kunieda. V-Pascal: An automatic vectorizing compiler for Pascal with no language extensions. The Journal of Supercomputing, 4(3):251–275, September 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=4&issue=3&spage=251.

Baillie:1990:LES

[86] Clive F. Baillie, Desmond A. Johnston, and Gregory W. Kilcup. Letter to the Editor: Status and prospects of the computational approach to high-energy physics. The Journal of Supercomputing, 4(3):277–300, September 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=4&issue=3&spage=277.

Reisman:1990:LEI

[87] Arnold Reisman. Letter to the Editor: Ionizing radiation effects on ULSI device yield and reliability. The Journal of Supercomputing, 4(3):301–310, September 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=3&spage=
301.

Anonymous:1990:CAb

[88] Anonymous. Contributing authors. The Journal of Supercomputing, 4(3): 311–314, September 1990. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=4&issue=3&spage=311.

Anonymous:1991:Ea

[89] Anonymous. Editorial. The Journal of Supercomputing, 4(4):319–320, January 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=4&spage=
319.

Wolfe:1991:DDP

[90] Michael Wolfe. Data dependence and program restructuring. The Journal of Supercomputing, 4(4):321–344, January 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=4&issue=4&spage=

Carlson:1991:ULM

[91] David A. Carlson. Using local memory to boost the performance of FFT algorithms on the CRAY-2 supercomputer. The Journal of Supercomputing, 4(4):345–356, January 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=4&spage=345.

Bailey:1991:USA

[92] David H. Bailey, King Lee, and Horst D. Simon. Using Strassen’s algorithm to accelerate the solution of linear systems. The Journal of Supercomputing, 4(4):357–371, January 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=4&spage=357. Ap-

peared as Technical Report RNR-90-001, NAS Systems Division, NASA Ames Laboratory, Moffat Field, CA, USA, February 1990.

Anonymous:1991:CAa

[93] Anonymous. Contributing authors. The Journal of Supercomputing, 4(4): 373–374, January 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=4&issue=4&spage=373.

Brenner:1991:SFM

[94] A. E. Brenner, J. P. Riganati, and P. B. Schneck. Sidney Fernbach — 1917–1991 — in memoriam. The Journal of Supercomputing, 5(1):5–6, June 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=1&spage=5.

Ortiz:1991:ALD

[95] Luis F. Ortiz, Ron Y. Pinter, and Shlomit S. Pinter. An array language for data parallelism: Definition, compilation, and applications. The Journal of Supercomputing, 5(1):7–29, June 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=5&issue=1&spage=
7.

MacDonald:1991:CNC

[96] Tom MacDonald. C for numerical computing. The Journal of Supercomputing, 5(1):31–48, June 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=1&spage=31.

Marsaglia:1991:NGR

[97] George Marsaglia. Normal (Gaussian) random variables for supercomputers. The Journal of Supercomputing, 5(1):49–55, June 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=1&spage=49.

Tao:1991:SPN

[98] Lixin Tao and Eva Ma. Simulating parallel neighboring communications among square meshes and square toruses. The Journal of Supercomputing, 5(1):57–71, June 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=5&issue=1&spage=57.

Woo:1991:CBC

[99] Jinwoon Woo and Sartaj Sahni. Computing biconnected components on a hypercube. The Journal of Supercomputing, 5(1):73–87, June 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=5&issue=1&spage=73.

Anonymous:1991:CAb

[100] Anonymous. Contributing authors. The Journal of Supercomputing, 5(1): 89–92, June 1991. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=5&issue=1&spage=89.

Anonymous:1991:Eb

[101] Anonymous. Editorial. The Journal of Supercomputing, 5(2–3):97, October 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=5&issue=2&spage=
                                        97.

Stiller:1991:GGC

[102] Lewis Stiller. Group graphs and computational symmetry on massively parallel architecture. The Journal of Supercomputing, 5(2–3):99–117, October 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=5&issue=2&spage=
99.

Tirumalai:1991:PWL

[103] Parthasarathy P. Tirumalai, Meng Lee, and Michael S. Schlansker. Parallelization of WHILE loops on pipelined architectures. The Journal of Supercomputing, 5(2–3):119–136, October 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=2&spage=119.

Malony:1991:TAP

[104] Allen D. Malony, John L. Larson, and Daniel A. Reed. Tracing application program execution on the Cray X-MP and Cray-2. The Journal of Supercomputing, 5(2–3):137–162, October 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=2&spage=137.

Lee:1991:EPA

[105] Yong-Fong Lee, Barbara G. Ryder, and Thomas J. Marlowe. Experiences with a parallel algorithm for data flow analysis. The Journal of Supercomputing, 5(2–3):163–188, October 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=2&spage=163.

Johnson:1991:MAP

[106] R. W. Johnson, C. H. Huang, and J. R. Johnson. Multilinear algebra and parallel programming. The Journal of Supercomputing, 5(2–3):189–217, October 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=5&issue=2&spage=
                                        189.

Cvetanovic:1991:EDP

[107] Zarka Cvetanovic, Edward G. Freedman, and Charles Nofsinger. Efficient decomposition and performance of parallel PDE, FFT, Monte Carlo simulations, simplex, and sparse solvers. The Journal of Supercomputing, 5(2– 3):219–238, October 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=5&issue=2&spage=219.

Anonymous:1991:CAc

[108] Anonymous. Contributing authors. The Journal of Supercomputing, 5(2– 3):239–244, October 1991. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=5&issue=2&spage=239.

Editorial:1992:E

[109] Guest Editorial.  Editorial.  The

Journal of Supercomputing, 5(4):249– 250, February 1992. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=5&issue=4&spage=249.

Kennedy:1992:SSF

[110] Ken Kennedy. Software for supercomputers of the future. The Journal of Supercomputing, 5(4):251–262, February 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=5&issue=4&spage=
251.

Dietz:1992:SSB

[111] Henry G. Dietz, Abderrazek Zaafrani, and Matthew T. O’Keefe. Static scheduling for barrier MIMD architectures. The Journal of Supercomputing, 5(4):263–289, February 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=4&spage=263.

Sheikh:1992:ILB

[112] Qasim Sheikh, Phuong Vu, Chao Yang, and Michael Merchant. Implementation of the level 2 and 3 BLAS on the CRAY Y-MP and the CRAY-2. The Journal of Supercomputing, 5(4):291–305, February 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=5&issue=4&spage=
291.

Ahalt:1992:IVQ

[113] Stanley C. Ahalt, Prakoon Chen, Cheng-Taou Chou, and Tzyy-Ping Jung. Implementation of a vector quantization codebook design technique based on a competitive learning artificial neural network. The Journal of Supercomputing, 5(4):307–330, February 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=4&spage=307.

Chakravarty:1992:PSH

[114] Sreejit Chakravarty and Ajay Shekhawat. Parallel and serial heuristics for the minimum set cover problem. The Journal of Supercomputing, 5(4):331–345, February 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=5&issue=4&spage=
331.

Allison:1992:HDH

[115] D. C. S. Allison, K. M. Irani, C. J. Ribbens, and L. T. Watson. Highdimensional homotopy curve tracking on a shared-memory multiprocessor. The Journal of Supercomputing, 5(4): 347–366, February 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=4&spage=347.

Anonymous:1992:CAa

[116] Anonymous. Contributing authors. The Journal of Supercomputing, 5(4): 367–372, February 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=5&issue=4&spage=367.

Dorozhevets:1992:EMM

[117] Mikhail N. Dorozhevets and Peter Wolcott. The El’brus-3 and MARS-M: Recent advances in Russian highperformance computing. The Journal of Supercomputing, 6(1):5–48, March 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=6&issue=1&spage=
5.

Hainline:1992:VPE

[118] Allen R. Hainline, Steven R. Thompson, and Lawrence L. Halcomb. Vector performance estimation for CRAY X-MP/ Y-MP supercomputers. The Journal of Supercomputing, 6(1):49–70, March 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=6&issue=1&spage=
49.

Weiss:1992:MCR

[119] Shlomo Weiss. Memory conflict resolution in vector supercomputers. The Journal of Supercomputing, 6 (1):71–85, March 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=6&issue=1&spage=71.

Deo:1992:PHO

[120] Narsingh Deo and Sushil Prasad. Parallel heap: an optimal parallel priority queue. The Journal of Supercomputing, 6(1):87–98, March 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=6&issue=1&spage=87.

Anonymous:1992:CAb

[121] Anonymous. Contributing authors. The Journal of Supercomputing, 6 (1):99–101, March 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=6&issue=1&spage=99.

Carlson:1992:UPF

[122] David A. Carlson. Ultrahigh-performance FFTS for the Cray-2 and Cray Y-MP supercomputers. The Journal of Supercomputing, 6(2):107–116, June 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=6&issue=2&spage=
                                        107.

Venkatakrishnan:1992:MIP

[123] V. Venkatakrishnan, Horst D. Simon, and Timothy J. Barth. A MIMD implementation of a parallel Euler solver for unstructured grids. The Journal of Supercomputing, 6(2):117–137, June 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=6&issue=2&spage=
                                        117.

Plimpton:1992:PSE

[124] S. J. Plimpton, J. R. Michael, and A. D. Romig, Jr. Parallel simulation of electron-solid interactions for electron microscopy modeling. The Journal of Supercomputing, 6(2):139–151, June 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=6&issue=2&spage=
139.

Geist:1992:PSC

[125] G. A. Geist, B. Ginatempo, W. A.

Shelton, and G. M. Stocks. Parallel superconductor code on the iPSC/ 860. The Journal of Supercomputing, 6(2):153–162, June 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=6&issue=2&spage=153.

Ripoll:1992:PMC

[126] Daniel R. Ripoll and Stephen J.

Thomas. A parallel Monte Carlo search algorithm for the conformational analysis of polypeptides. The Journal of Supercomputing, 6(2):163–185, June 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=6&issue=2&spage=
163.

Anonymous:1992:CAc

[127] Anonymous. Contributing authors.

The Journal of Supercomputing, 6 (2):187–190, June 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=6&issue=2&spage=187.

Frank:1992:LQH

[128] Joachim Frank and Siegfried Knecht. Lattice quantum hadrodynamics on a Cray Y-MP. The Journal of Supercomputing, 6(3–4):195–209, December 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=6&issue=3&spage=
                                        195.

Azmy:1992:PPM

[129] Yousry Y. Azmy. Performance and performance modeling of a parallel algorithm for solving the neutron transport equation. The Journal of Supercomputing, 6(3–4):211–235, December 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=6&issue=3&spage=
                                        211.

Kratzer:1992:SQF

[130] Steven G. Kratzer. Sparse QR factorization on a massively parallel computer. The Journal of Supercomputing, 6(3–4):237–255, December 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=6&issue=3&spage=237.

Chung:1992:MFE

[131] Yeh-Ching Chung and Sanjay Ranka. Mapping finite element graphs on hypercubes. The Journal of Supercomputing, 6(3–4):257–282, December 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=6&issue=3&spage=257.

Gokhale:1992:ICI

[132] Maya Gokhale and William Carlson. An introduction to compilation issues for parallel machines. The Journal of Supercomputing, 6(3–4):283–314, December 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=6&issue=3&spage=
283.

Luecke:1992:PPC

[133] Glenn R. Luecke, Jae Heon Yun, and Philip W. Smith. Performance of parallel Cholesky factorization algorithms using BLAS. The Journal of Supercomputing, 6(3–4):315–329, December 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=6&issue=3&spage=
315.

Anonymous:1992:CAd

[134] Anonymous. Contributing authors.

The Journal of Supercomputing, 6(3– 4):331–334, December 1992. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=6&issue=3&spage=331.

Draper:1993:SII

[135] Richard Draper and John Riganati. Special issue on instruction-level parallelism — introduction. The Journal of Supercomputing, 7(1–2):5, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=5.

Fisher:1993:GEI

[136] Josh Fisher and Bob Rau. Guest Editors’ introduction. The Journal of Supercomputing, 7(1–2):7, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=7.

Rau:1993:ILPa

[137] B. Ramakrishna Rau and Joseph A. Fisher. Instruction-level parallel processing: History, overview, and perspective. The Journal of Supercomputing, 7(1–2):9–50, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=9.

Lowney:1993:MTS

[138] P. Geoffrey Lowney, Stefan M. Freudenberger, Thomas J. Karzes, W. D. Lichtenstein, Robert P. Nix, John S. O’Donnell, and John C. Ruttenberg. The Multiflow Trace Scheduling compiler. The Journal of Supercomputing, 7(1–2):51–142, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=51.

Beck:1993:CMA

[139] Gary R. Beck, David W. L. Yen, and Thomas L. Anderson. The Cydra 5 minisupercomputer: Architecture and implementation. The Journal of Supercomputing, 7(1–2):143–180, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=143.

Dehnert:1993:CC

[140] James C. Dehnert and Ross A. Towle.

Compiling for the Cydra 5. The Journal of Supercomputing, 7(1–2):181–227, May 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=7&issue=1&spage=
181.

Hwu:1993:SET

[141] Wen mei W. Hwu, Scott A. Mahlke, William Y. Chen, Pohua P. Chang, Nancy J. Warter, Roger A. Bringmann, Roland G. Ouellette, Richard E. Hank, Tokuzo Kiyohara, Grant E. Haab, J. G. Holm, and D. M. Lavery. The superblock: An effective technique for VLIW and superscalar compilation. The Journal of Supercomputing, 7(1–2):229–248, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=229.

Schuette:1993:ILE

[142] Michael A. Schuette and John P. Shen.

Instruction-level experimental evaluation of the Multiflow TRACE 14/ 300 VLIW computer. The Journal of Supercomputing, 7(1–2):249–271, May 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=7&issue=1&spage=
249.

Anonymous:1993:CAa

[143] Anonymous. Contributing authors. The Journal of Supercomputing, 7 (1–2):273–282, May 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=1&spage=273.

Procassini:1993:PGO

[144] Richard J. Procassini, Scott R. Whitman, and William P. Dannevik. Porting a global ocean model onto a sharedmemory multiprocessor: Observations and guidelines. The Journal of Supercomputing, 7(3):287–321, September 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=7&issue=3&spage=
287.

Ling:1993:SHP

[145] Per Ling. A set of high-performance level 3 BLAS structured and tuned for the IBM 3090 VF and implemented in Fortran 77. The Journal of Supercomputing, 7(3):323–355, September 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=7&issue=3&spage=323.

Glenn:1993:BTP

[146] Raymond R. Glenn, Daniel V. Pryor, John M. Conroy, and Theodore Johnson. A bistability throughput phenomenon in a shared-memory MIMD machine. The Journal of Supercomputing, 7(3):357–375, September 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=3&spage=357.

Anonymous:1993:CAb

[147] Anonymous. Contributing authors.

The Journal of Supercomputing, 7(3): 377–379, September 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=3&spage=377.

Anonymous:1993:E

[148] Anonymous. Editorial. The Journal of Supercomputing, 7(4):385– 386, December 1993. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=7&issue=4&spage=385.

Knobe:1993:ADA

[149] Kathleen Knobe and Venkataraman Natarajan. Automatic data allocation to minimize communication on SIMD machines. The Journal of Supercomputing, 7(4):387–415, December 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=7&issue=4&spage=387.

Hohlfeld:1993:IPC

[150] R. G. Hohlfeld, N. F. Comins, D. Shalit, P. A. Shorey, and R. C. Giles. Implementation of particle-in-cell stellar dynamics codes on the Connection Machine-2. The Journal of Supercomputing, 7(4):417–436, December 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=7&issue=4&spage=417.

Thompson:1993:VPE

[151] Steven R. Thompson, Allen R. Hainline, and Lawrence L. Halcomb. Vector performance estimation for Cray X-MP/Y-MP supercomputers, part 2. The Journal of Supercomputing, 7(4): 437–467, December 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=7&issue=4&spage=437.

Anonymous:1993:CAc

[152] Anonymous. Contributing authors. The Journal of Supercomputing, 7(4): 469–471, December 1993. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=7&issue=4&spage=469.

Lyon:1994:SPT

[153] Gordon Lyon, Robert Snelick, and Raghu Kacker. Synthetic-perturbation tuning of MIMD programs. The Journal of Supercomputing, 8(1):5–28, March 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=8&issue=1&spage=
5.

Blom:1994:VMO

[154] J. G. Blom and J. G. Verwer. Vectorizing matrix operations arising from PDE discretization on 9-point stencils. The Journal of Supercomputing, 8(1):29–51, March 1994. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=1&spage=29.

Mansour:1994:PPO

[155] Nashat Mansour and Geoffrey C. Fox.

Parallel physical optimization algorithms for allocating data to multicomputer nodes. The Journal of Supercomputing, 8(1):53–80, March 1994. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=1&spage=53.

Fatoohi:1994:ANS

[156] R. A. Fatoohi. Adapting a Navier– Stokes solver for three parallel machines. The Journal of Supercomputing, 8(2):91–115, June 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=2&spage=91.

FroeseFischer:1994:DMI

[157] Charlotte Froese Fischer, Ming Tong, Murry Bentley, Zuchang Shen, and C. Ravimohan. A distributed-memory implementation of the MCHF atomic structure package. The Journal of Supercomputing, 8(2):117–134, June 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=8&issue=2&spage=
117.

Elsesser:1994:PPL

[158] Gary W. Elsesser, Viet N. Ngo, Sourav Bhattacharya, and Wei-Tek Tsai. Processor preallocation and load balancing of DOALL loops. The Journal of Supercomputing, 8(2):135–161, June 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=8&issue=2&spage=
135.

Yousif:1994:CCP

[159] Mazin S. Yousif, Chita R. Das, and Matthew J. Thazhuthaveetil. A cache coherence protocol for MIN-Based multiprocessors. The Journal of Supercomputing, 8(2):163–185, June 1994. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=2&spage=163.

Anonymous:1994:CAa

[160] Anonymous. Contributing authors.

The Journal of Supercomputing, 8 (2):187–189, June 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=2&spage=187.

Shapiro:1994:MPG

[161] Bruce A. Shapiro and Joseph Navetta.

A massively parallel genetic algorithm for RNA secondary structure prediction. The Journal of Supercomputing, 8(3):195–207, November 1994. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=3&spage=195.

Robbins:1994:RBA

[162] Kay A. Robbins and Steven Robbins.

Relationship between average and real memory behavior. The Journal of Supercomputing, 8(3):209–232, November 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=8&issue=3&spage=
209.

Lee:1994:EEP

[163] Yong-Fong Lee and Barbara G. Ryder. Effectively exploiting parallelism in data flow analysis. The Journal of Supercomputing, 8(3):233–262, November 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        0920-8542&volume=8&issue=3&spage=
                                        233.

Boals:1994:IHA

[164] Alfred J. Boals, Ajay K. Gupta, and Naveed A. Sherwani. Incomplete hypercubes: Algorithms and embeddings. The Journal of Supercomputing, 8(3): 263–294, November 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=8&issue=3&spage=263.

Anonymous:1994:CAb

[165] Anonymous. Contributing authors. The Journal of Supercomputing, 8(3): 295–298, November 1994. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=8&issue=3&spage=295.

Thakur:1995:CEC

[166] Rajeev Thakur, Ravi Ponnusamy, Alok Choudhary, and Geoffrey Fox. Complete exchange on the CM-5 and Touchstone Delta. The Journal of Supercomputing, 8(4):305–328, ???? 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=8&issue=4&spage=305.

Mavriplis:1995:IPU

[167] D. J. Mavriplis, Raja Das, Joel Saltz, and R. E. Vermeland. Implementation of a parallel unstructured Euler solver on shared- and distributed-memory architectures. The Journal of Supercomputing, 8(4):329–344, ???? 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=4&spage=329.

Farkas:1995:SCC

[168] K. Farkas, Z. Vranesic, and M. Stumm.

Scalable cache consistency for hierarchically structured multiprocessors. The Journal of Supercomputing, 8 (4):345–369, ???? 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=4&spage=345.

Netzer:1995:OTR

[169] R. H. B. Netzer and B. P. Miller.

Optimal tracing and replay for debugging message-passing parallel programs. The Journal of Supercomputing, 8(4):371–388, ???? 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=8&issue=4&spage=371.

Wu:1995:SIT

[170] Min-You Wu, Wolfgang Schröder-Preikschat, R. Draper, and J. Riganati. Special issue on trends in parallel operating systems. The Journal of Supercomputing, 9(1–2):5–6, March 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=1&spage=5.

Tanenbaum:1995:CTM

[171] Andrew S. Tanenbaum. A comparison of three microkernels. The Journal of Supercomputing, 9(1–2):7–22, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=9&issue=1&spage=
                                        7.

Tripathi:1995:TMD

[172] Anand R. Tripathi and Neeran M. Karnik. Trends in multiprocessor and distributed operating systems designs. The Journal of Supercomputing, 9(1–2):23–49, March 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=9&issue=1&spage=23.

Ramachandran:1995:DSS

[173] Mahendra Ramachandran and Mukesh Singhal. Decentralized semaphore support in a virtual shared-memory system. The Journal of Supercomputing, 9(1–2):51–70, March 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=9&issue=1&spage=51.

Assenmacher:1995:PPP

[174] H. Assenmacher, T. Breitbach, P. Buhler, V. Hubsch, H. Peine, and R. Schwarz. Parallel programming in Panda. The Journal of Supercomputing, 9(1–2):71–90, March 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=1&spage=71.

Shu:1995:RTS

[175] Wei Shu. Run-time support for user-level ultralightweight threads on distributed-memory computers. The Journal of Supercomputing, 9(1–2): 91–103, March 1995. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=1&spage=91.

Unrau:1995:HCS

[176] Ronald C. Unrau, Orran Krieger, Benjamin Gamsa, and Michael Stumm. Hierarchical clustering: a structure for scalable multiprocessor operating system design. The Journal of Supercomputing, 9(1–2):105–134, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=1&spage=
105.

Ahmad:1995:MPF

[177] Ishfaq Ahmad. A massively parallel fault-tolerant architecture for timecritical computing. The Journal of Supercomputing, 9(1–2):135–162, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=1&spage=

Wadleigh:1995:HPF

[178] Kevin R. Wadleigh, Gary B. Gostin, and John Liu. High-performance FFT algorithms for the Convex C4/XA supercomputer. The Journal of Supercomputing, 9(1–2):163–178, March 1995. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=9&issue=1&spage=163.

Trevisan:1995:CCT

[179] Vilmar Trevisan and João B. Carvalho. The composite character of the twentysecond Fermat number. The Journal of Supercomputing, 9(1–2):179–182, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=9&issue=1&spage=
                                        179.

Sheu:1995:PMN

[180] Jang-Ping Sheu and Tzung-Shi Chen. Partitioning and mapping of nested loops for linear array multicomputers. The Journal of Supercomputing, 9(1–2):183–202, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=9&issue=1&spage=183.

Wu:1995:CA

[181] Min-You Wu. Contributing authors. The Journal of Supercomputing, 9(1– 2):203–211, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=9&issue=1&spage=203.

Anonymous:1995:Ea

[182] Anonymous. Editorial. The Journal of Supercomputing, 9(1–2):??, March 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=1&spage=
??.

Buell:1995:CCM

[183] Duncan A. Buell and Kenneth L. Pocek.

Custom computing machines: an introduction. The Journal of Supercomputing, 9(3):219–229, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=3&spage=
219;   http://www.wkap.nl/oasis.
htm/95683.

Iseli:1995:SSS

[184] Christian Iseli and Eduardo Sanchez.

Spyder: A SURE (SUperscalar and REconfigurable) processor. The Journal of Supercomputing, 9(3):231–252, September 1995. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=3&spage=231;   http:
//www.wkap.nl/oasis.htm/95686.

Ling:1995:WMD

[185] Xiaoping Ling and Hideharu Amano.

WASMII: An MPLD with datadriven control on a virtual hardware. The Journal of Supercomputing, 9(3):253–276, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=3&spage=
253;   http://www.wkap.nl/oasis.
htm/95688.

Arnold:1995:SSE

[186] Jeffrey M. Arnold. The Splash 2 software environment. The Journal of Supercomputing, 9(3):277–290, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=3&spage=
277;   http://www.wkap.nl/oasis.
htm/95690.

Gokhale:1995:DPC

[187] Maya Gokhale and Brian Schott. Data-parallel C on a reconfigurable logic array. The Journal of Supercomputing, 9(3):291–313, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=3&spage=
291;   http://www.wkap.nl/oasis.
htm/95691.

Louie:1995:VPS

[188] Marianne E. Louie and Milos D. Ercegovac. A variable-precision square root implementation for field programmable gate arrays. The Journal of Supercomputing, 9(3):315–336, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=3&spage=
315;   http://www.wkap.nl/oasis.
htm/95692.

Anonymous:1995:CAa

[189] Anonymous. Contributing authors.

The Journal of Supercomputing, 9(3): 337–340, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=3&spage=337.

Anonymous:1995:Eb

[190] Anonymous. Editorial. The Journal of Supercomputing, 9(4):345– 346, December 1995. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=4&spage=345.

Prestin:1995:PNS

[191] Michael Prestin and Leonid Shtilman. A parallel Navier–Stokes solver: The Meiko implementation. The Journal of Supercomputing, 9(4):347–364, December 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=4&spage=
347;   http://www.wkap.nl/oasis.
htm/103754.

Ramanujam:1995:BUT

[192] J. Ramanujam. Beyond unimodular transformations. The Journal of Supercomputing, 9(4):365–389, December 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=9&issue=4&spage=
365;   http://www.wkap.nl/oasis.
htm/103755.

Gao:1995:WCU

[193] Hui Gao and John L. Larson. Workload characterization using the Cray Hardware Performance Monitor. The Journal of Supercomputing, 9(4):391– 412, December 1995. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=4&spage=391;   http:
//www.wkap.nl/oasis.htm/103756.

Anonymous:1995:CAb

[194] Anonymous. Contributing authors. The Journal of Supercomputing, 9(4): 413–414, December 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=9&issue=4&spage=413.

Cypher:1996:QSP

[195] Robert Cypher, Alex Ho, Smaragda Konstantinidou, and Paul Messina. A quantitative study of parallel scientific applications with explicit communication. The Journal of Supercomputing, 10(1):5–24, March 1996. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&

volume=10&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/113484.

Cameron:1996:PPM

[196] Gordon D. B. Cameron and Gordon I. D. Duncan. PARAMICS — parallel microscopic simulation of road traffic. The Journal of Supercomputing, 10(1):25–53, March 1996. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=1&spage=25;   http:
//www.wkap.nl/oasis.htm/113486.

Bae:1996:CDM

[197] Seungjo Bae and Sanjay Ranka.

A comparison of different messagepassing paradigms for the parallelization of two irregular applications. The Journal of Supercomputing, 10 (1):55–85, March 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=1&spage=55;   http:
//www.wkap.nl/oasis.htm/113487.

Burger:1996:PTD

[198] Douglas C. Burger, Rahmat S. Hyder, Barton P. Miller, and David A. Wood. Paging tradeoffs in distributedshared-memory multiprocessors. The Journal of Supercomputing, 10(1):87– 104, March 1996. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=1&spage=87;   http:
//www.wkap.nl/oasis.htm/113488.

Anonymous:1996:CAa

[199] Anonymous. Contributing authors. The Journal of Supercomputing, 10 (1):105–??, ???? 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=1&spage=105.

Ou:1996:FPM

[200] Chao-Wei Ou, Sanjay Ranka, and Geoffrey Fox. Fast and parallel mapping algorithms for irregular problems. The Journal of Supercomputing, 10(2):119–140, June 1996. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=2&spage=119;   http:
                                        //www.wkap.nl/oasis.htm/116934.

Bader:1996:PAI

[201] David A. Bader, Joseph JáJá, David Harwood, and Larry S. Davis. Parallel algorithms for image enhancement and segmentation by region growing, with an experimental study. The Journal of Supercomputing, 10 (2):141–168, June 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=2&spage=141;   http:
                                        //www.wkap.nl/oasis.htm/116935.

Nieplocha:1996:GAN

[202] Jarosław Nieplocha, Robert J. Harrison, and Richard J. Littlefield. Global arrays: a nonuniform memory access programming model for high-performance computers. The Journal of Supercomputing, 10(2):169–189, June 1996. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=2&spage=169;   http:
//www.wkap.nl/oasis.htm/116936.

Seamons:1996:MAP

[203] Kent E. Seamons and Marianne Winslett. Multidimensional array I/O in Panda 1.0. The Journal of Supercomputing, 10(2):191–211, June 1996. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=2&spage=191;   http:
//www.wkap.nl/oasis.htm/116937.

Anonymous:1996:CAb

[204] Anonymous. Contributing authors.

The Journal of Supercomputing, 10 (2):213–??, ???? 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=2&spage=213.

Abdelrahman:1996:LHC

[205] Tarek S. Abdelrahman. Latency hiding on COMA multiprocessors. The Journal of Supercomputing, 10(3):225– 242, September 1996. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=3&spage=225;   http:
//www.wkap.nl/oasis.htm/120598.

Arabnia:1996:PSR

[206] Hamid R. Arabnia and Suchendra M. Bhandarkar. Parallel stereocorrelation on a reconfigurable multi-ring network. The Journal of Supercomputing, 10(3): 243–269, September 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=3&spage=243;   http:
                                        //www.wkap.nl/oasis.htm/120599.

Houlahan:1996:HSA

[207] Joanne F. Houlahan, Lenore J. Cowen, and Gerald M. Masson. Hypercube sandwich approach to conferencing. The Journal of Supercomputing, 10(3): 271–283, September 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=3&spage=271;   http:
                                        //www.wkap.nl/oasis.htm/120600.

Shoemaker:1996:NAO

[208] David Shoemaker, Frank Honoré, Chris Metcalf, and Steve Ward. NuMesh: An architecture optimized for scheduled communication. The Journal of Supercomputing, 10(3):285– 302, September 1996. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=3&spage=285;   http:
                                        //www.wkap.nl/oasis.htm/120601.

Anonymous:1996:CAc

[209] Anonymous. Contributing authors. The Journal of Supercomputing, 10(3):

303–305, September 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=3&spage=303.

Arabnia:1996:SIP

[210] H. R. Arabnia. A special issue on parallel and distributed processing. The Journal of Supercomputing, 10(3):U2–U3, ???? 1996. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=3&spage=U2.

Draper:1996:DSM

[211] R. Draper and J. Riganati. Daniel Shanks, 1917–1996 — in memoriam. The Journal of Supercomputing, 10 (3):U3, ???? 1996. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=3&spage=U3.

Fallah-Adl:1997:FAE

[212] Hassan Fallah-Adl, Joseph JáJá, and Shunlin Liang. Fast algorithms for estimating aerosol optical depth and correcting thematic mapper imagery. The Journal of Supercomputing, 10(4): 315–329, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=4&spage=315;   http:
//www.wkap.nl/oasis.htm/134566.

Wang:1997:TDE

[213] J. Wang, P. Liewer, and E. Huang. Three-dimensional electromagnetic particlein-cell with Monte Carlo collision simulations on three MIMD parallel computers. The Journal of Supercomputing, 10(4):331–348, December 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=4&spage=331;   http:
                                        //www.wkap.nl/oasis.htm/134568.

Ahmad:1997:MOC

[214] Ishfaq Ahmad, Ka-Cheong Leung, and Hsiao-Ming Hsu. Multiprocessing ocean circulation: Modeling, implementation, and performance on the Intel Paragon. The Journal of Supercomputing, 10(4): 349–369, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=4&spage=349;   http:
                                        //www.wkap.nl/oasis.htm/134569.

Dixon:1997:HPS

[215] Brandon Dixon and John Swallow. High-performance sorting algorithms for the CRAY T3D parallel computer. The Journal of Supercomputing, 10(4): 371–395, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=10&issue=4&spage=371;   http:
                                        //www.wkap.nl/oasis.htm/134570.

Anonymous:1997:CAa

[216] Anonymous. Contributing authors.

The Journal of Supercomputing, 10(4): 397–400, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=10&issue=4&spage=397.

Draper:1997:FE

[217] Richard Draper and John Riganati.

Farewell editorial. The Journal of Supercomputing, 11(1):5–6, March 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/141468.

Averbuch:1997:HST

[218] A. Averbuch, L. Ioffe, M. Israeli, and L. Vozovoi. Highly scalable two- and three-dimensional Navier–Stokes parallel solvers on MIMD multiprocessors. The Journal of Supercomputing, 11(1):7–39, March 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/141469.

Jayasimha:1997:EAP

[219] D. N. Jayasimha, M. E. Hayder, and S. K. Pillay. An evaluation of architectural platforms for parallel Navier–Stokes computations. The Journal of Supercomputing, 11(1): 41–60, March 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=1&spage=41;   http:
                                        //www.wkap.nl/oasis.htm/141470.

Bhandarkar:1997:CRP

[220] Suchendra M. Bhandarkar and Salem Machaka. Chromosome reconstruction from physical maps using a cluster of workstations. The Journal of Supercomputing, 11(1):61–86, March 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=1&spage=61;   http:
                                        //www.wkap.nl/oasis.htm/141471.

Anonymous:1997:CAb

[221] Anonymous. Contributing authors. The Journal of Supercomputing, 11 (1):87–??, ???? 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=1&spage=87.

Hariri:1997:ESI

[222] Salim Hariri. Editorial: Special issue on high performance distributed computing. The Journal of Supercomputing, 11(2):99, October 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=2&spage=99;   http:
                                        //www.wkap.nl/oasis.htm/149820.

Chandy:1997:WAD

[223] K. Mani Chandy, Joseph Kiniry, Adam Rifkin, and Daniel Zimmerman. Webs of archived distributed computations for asynchronous collaboration. The Journal of Supercomputing, 11(2):101– 118, October 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=2&spage=101;   http:
//www.wkap.nl/oasis.htm/149821.

Nieplocha:1997:SMP

[224] Jarek Nieplocha and Robert J. Harrison. Shared memory programming in metacomputing environments: The global array approach. The Journal of Supercomputing, 11(2):119– 136, October 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=2&spage=119;   http:
//www.wkap.nl/oasis.htm/149823.

Silva:1997:CDS

[225] Luis M. Silva and João Gabriel Silva.

Checkpointing distributed shared memory. The Journal of Supercomputing, 11(2):137–158, October 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=2&spage=137;   http:
//www.wkap.nl/oasis.htm/149824.

Park:1997:HPM

[226] Sung-Yong Park and Salim Hariri.

A high performance message passing system for Network of Workstations. The Journal of Supercomputing, 11 (2):159–180, October 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=2&spage=159;   http:
//www.wkap.nl/oasis.htm/149826.

Sarkar:1997:ANI

[227] Prasenjit Sarkar and Mary Bailey. Adapting the network interface for highperformance computing: The CNI approach. The Journal of Supercomputing, 11(2):181–200, October 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=2&spage=181;   http:
//www.wkap.nl/oasis.htm/149828.

Anonymous:1997:CAc

[228] Anonymous. Contributing authors. The Journal of Supercomputing, 11 (2):201–204, October 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=2&spage=201;   http:
//www.wkap.nl/oasis.htm/149830.

Andrews:1997:SMI

[229] Phil Andrews. Supercomputing in medicine — introduction. The Journal of Supercomputing, 11(3):211– 212, November 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&

volume=11&issue=3&spage=211;   http:
//www.wkap.nl/oasis.htm/144233.

McQueen:1997:SMP

[230] David M. McQueen and Charles S. Peskin. Shared-memory parallel vector implementation of the immersed boundary method for the computation of blood flow in the beating mammalian heart. The Journal of Supercomputing, 11(3):213–236, November 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=3&spage=213;   http:
//www.wkap.nl/oasis.htm/144235.

Ropelewski:1997:IGS

[231] Alexander J. Ropelewski, Hugh B.

Nicholas, Jr., and David W. Deerfield, II. Implementation of genetic sequence alignment programs on supercomputers. The Journal of Supercomputing, 11(3):237–253, November 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=3&spage=237;   http:
//www.wkap.nl/oasis.htm/144236.

Crowley:1997:AIS

[232] Michael F. Crowley, Tom A. Darden, Thomas E. Cheatham, III, and David W. Deerfield, II. Adventures in improving the scaling and accuracy of a parallel molecular dynamics program. The Journal of Supercomputing, 11(3):255–278, November 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=3&spage=255;   http:
                                        //www.wkap.nl/oasis.htm/144238.

Wetzel:1997:CAP

[233] Arthur W. Wetzel. Computational aspects of pathology image classification and retrieval. The Journal of Supercomputing, 11(3):279– 293, November 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=3&spage=279;   http:
                                        //www.wkap.nl/oasis.htm/144239.

Goddard:1997:OAF

[234] Nigel H. Goddard, Greg Hood, Jonathan D. Cohen, William F. Eddy, Christopher R. Genovese, Douglas C. Noll, and Leigh E. Nystrom. Online analysis of functional MRI datasets on parallel platforms. The Journal of Supercomputing, 11(3):295– 318, November 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=3&spage=295;   http:
                                        //www.wkap.nl/oasis.htm/144241.

Anonymous:1997:CAd

[235] Anonymous. Contributing authors. The Journal of Supercomputing, 11(3): 319–324, November 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=11&issue=3&spage=319;   http:
                                        //www.wkap.nl/oasis.htm/150546.

Graham:1997:ESI

[236] Peter Graham and Brian J. d’Auriol.

Editorial: Special issue — highperformance computing systems. The Journal of Supercomputing, 11(4): 331, December 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=331;   http:
//www.wkap.nl/oasis.htm/134882.

Lengauer:1997:SPL

[237] Christian Lengauer, Sergei Gorlatch, and Christoph A. Herrmann. The static parallelization of loops and recursions. The Journal of Supercomputing, 11(4):333–353, December 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=333;   http:
//www.wkap.nl/oasis.htm/147009.

Brune:1997:HMP

[238] Matthias Brune, Jörn Gehring, and Alexander Reinefeld. Heterogeneous message passing and a link to resource management. The Journal of Supercomputing, 11(4):355– 369, December 1997. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=355;   http:
//www.wkap.nl/oasis.htm/147011.

Chan:1997:EFM

[239] Edward Chan, Daoxu Chen, Victor Lee, Jiannong Cao, and Chan-Hee Lee.

On the effectiveness of the FDDI-M medium access protocol for real-time traffic. The Journal of Supercomputing, 11(4):371–389, December 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=371;   http:
//www.wkap.nl/oasis.htm/147012.

Li:1997:CTB

[240] Keqin Li. Constant time Boolean matrix multiplication on a linear array with a reconfigurable pipelined bus system. The Journal of Supercomputing, 11(4):391–403, December 1997. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=391;   http:
//www.wkap.nl/oasis.htm/147013.

Liu:1997:PSB

[241] Kuang-Chih Liu and Chung-Ta King. A performance study on bounteous transfer in multiprocessor sectored caches. The Journal of Supercomputing, 11(4): 405–420, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=405;   http:
//www.wkap.nl/oasis.htm/147014.

Anonymous:1997:CAe

[242] Anonymous. Contributing authors. The Journal of Supercomputing, 11(4): 421–422, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=11&issue=4&spage=421;   http:
//www.wkap.nl/oasis.htm/153600.

Arabnia:1998:E

[243] Hamid R. Arabnia. Editorial. The Journal of Supercomputing, 12(1–2):5– 6, January/February 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/158087.

Armstrong:1998:PIC

[244] James B. Armstrong, Muthucumaru Maheswaran, Mitchell D. Theys, Howard Jay Siegel, Mark A. Nichols, and Kenneth H. Casey. Parallel image correlation: Case study to examine trade-offs in algorithmto-machine mappings. The Journal of Supercomputing, 12(1–2):7–35, January/February 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/158088.

Wallace:1998:DSP

[245] A. M. Wallace, G. J. Michaelson, N. Scaife, and W. J. Austin. A dual source, parallel architecture for computer vision. The Journal of Supercomputing, 12(1–2):37–56, January/February 1998. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=12&issue=1&spage=37;   http:
                                        //www.wkap.nl/oasis.htm/158089.

Heirich:1998:CAL

[246] Alan Heirich and James Arvo. A competitive analysis of load balancing strategies for parallel ray tracing. The Journal of Supercomputing, 12(1– 2):57–68, January/February 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=12&issue=1&spage=57;   http:
                                        //www.wkap.nl/oasis.htm/158090.

Shi:1998:PMA

[247] Hongchi Shi, Paul Gader, and Hongzheng Li. Parallel mesh algorithms for grid graph shortest paths with application to separation of touching chromosomes. The Journal of Supercomputing, 12(1– 2):69–83, January/February 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=12&issue=1&spage=69;   http:
                                        //www.wkap.nl/oasis.htm/158091.

Gorlatch:1998:PDC

[248] Sergei Gorlatch. Programming with divide-and-conquer skeletons: a case study of FFT. The Journal of Supercomputing, 12(1–2):85–97, January/February 1998. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=12&issue=1&spage=85;   http:
                                        //www.wkap.nl/oasis.htm/158093.

Yang:1998:CGP

[249] Linda Yang and Manavendra Misra.

Coarse-grained parallel algorithms for multi-dimensional wavelet transforms. The Journal of Supercomputing, 12(1– 2):99–118, January/February 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=1&spage=99;   http:
//www.wkap.nl/oasis.htm/158096.

Houzet:1998:PBS

[250] Dominique Houzet and Abdelkrim Fatni. PC-based shared memory architecture and language. The Journal of Supercomputing, 12(1–2):119– 136, January/February 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=1&spage=119;   http:
//www.wkap.nl/oasis.htm/158097.

Mabin:1998:PAR

[251] Frederic H. Mabin and Catherine Mongenet. A parallel algorithm to reconstruct bounding surfaces in 3D images. The Journal of Supercomputing, 12(1–2):137–155, January/February 1998. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=1&spage=137;   http:
//www.wkap.nl/oasis.htm/158099.

Johasz:1998:AMP

[252] Zoltan Johasz. An analytical method for predicting the performance of parallel image processing operations. The Journal of Supercomputing, 12(1–2): 157–174, January/February 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=1&spage=157;   http:
//www.wkap.nl/oasis.htm/158100.

Ayed:1998:AHC

[253] Moez Ayed and Jean-Luc Gaudiot. Analysis of a heuristic for code partitioning. The Journal of Supercomputing, 12(3):191–226, May 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=3&spage=191;   http:
//www.wkap.nl/oasis.htm/163887.

Fahringer:1998:ESA

[254] Thomas Fahringer. Efficient symbolic analysis for parallelizing compilers and performance estimators. The Journal of Supercomputing, 12 (3):227–252, May 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=3&spage=227;   http:
//www.wkap.nl/oasis.htm/163888.

Hsu:1998:EMA

[255] Ching-Hsien Hsu and Yeh-Ching Chung. Efficient methods for kr ⟶ r and r ⟶ kr array redistribution. The Journal of Supercomputing, 12 (3):253–276, May 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=3&spage=253;   http:
//www.wkap.nl/oasis.htm/163889.

Latifi:1998:SFD

[256] Shahram Latifi and Pradip K. Srimani. SEP: a fixed degree regular network for massively parallel systems. The Journal of Supercomputing, 12(3):277–291, May 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=3&spage=277;   http:
//www.wkap.nl/oasis.htm/163891.

Aluru:1998:DIH

[257] Srinivas Aluru, John Gustafson, G. M.

Prabhu, and Fatih E. Sevilgen. Distribution-independent hierarchical algorithms for the N-body problem. The Journal of Supercomputing, 12 (4):303–323, October 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=12&issue=4&spage=303;   http:
//www.wkap.nl/oasis.htm/183399.

Darbha:1998:RCT

[258] Sekhar Darbha and Santosh Pande.

A robust compile time method for scheduling task parallelism on distributed memory machines. The Journal of Supercomputing, 12(4):325– 347, October 1998. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&

                                        volume=12&issue=4&spage=325;   http:
                                        //www.wkap.nl/oasis.htm/183400.

Abdelrahman:1998:CSA

[259] Tarek S. Abdelrahman and Thomas N. Wong. Compiler support for array distribution on NUMA shared memory multiprocessors. The Journal of Supercomputing, 12(4):349–371, October 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=12&issue=4&spage=349;   http:
                                        //www.wkap.nl/oasis.htm/183401.

Lin:1998:ESA

[260] Yen-Chun Lin and Jyh-Chian Chen. An efficient systolic algorithm for the longest common subsequence problem. The Journal of Supercomputing, 12 (4):373–385, October 1998. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=12&issue=4&spage=373;   http:
                                        //www.wkap.nl/oasis.htm/183402.

Budenske:1998:MLU

[261] John R. Budenske, Ranga S. Ramanujan, and Howard Jay Siegel. A method for the on-line use of off-line derived remappings of iterative automatic target recognition tasks onto a particular class of heterogeneous parallel platforms. The Journal of Supercomputing, 12(4):387–406, October 1998. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&

volume=12&issue=4&spage=387;   http:
//www.wkap.nl/oasis.htm/183403.

Li:1999:SSP

[262] Kei-Chun Li and Kang Zhang. Supporting scalable performance monitoring and analysis of parallel programs. The Journal of Supercomputing, 13(1):5–31, January 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/193732.

Eckert:1999:IRM

[263] C. Eckert and M. Pizka. Improving resource management in distributed systems using language-level structuring concepts. The Journal of Supercomputing, 13(1):33–55, January 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=1&spage=33;   http:
//www.wkap.nl/oasis.htm/193735.

Omori:1999:OOF

[264] Youichi Omori, Akira Fukuda, and Kazuki Joe. An object-oriented framework for loop parallelization. The Journal of Supercomputing, 13(1):57– 69, January 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=1&spage=57;   http:
//www.wkap.nl/oasis.htm/193737.

Fu:1999:LCR

[265] Ada Waichee Fu and Siu-Cheung Chau. Locating corruptions in a replicated file in a distributed environment. The Journal of Supercomputing, 13(1):71–97, January 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=1&spage=71;   http:
                                        //www.wkap.nl/oasis.htm/193739.

Hamdi:1999:PCE

[266] Mounir Hamdi, Yi Pan, B. Hamidzadeh, and F. M. Lim. Parallel computing on an Ethernet cluster of workstations: Opportunities and constraints. The Journal of Supercomputing, 13 (2):111–132, March 1999. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=2&spage=111;   http:
                                        //www.wkap.nl/oasis.htm/200574.

Galdamez:1999:EBT

[267] Pablo Galdámez, Declan Murphy, José M. Bernabéu-Aubán, and Francesc D. Muñoz-Escoó. Event-based techniques to debug an object request broker. The Journal of Supercomputing, 13 (2):133–149, March 1999. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=2&spage=133;   http:
                                        //www.wkap.nl/oasis.htm/200576.

Chung:1999:PDJ

[268] Soon M. Chung and Arindam Chatterjee. Parallel distributive join algorithm on the Intel Paragon. The Journal of Supercomputing, 13(2):151– 169, March 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=2&spage=151;   http:
//www.wkap.nl/oasis.htm/200577.

Claver:1999:PWA

[269] José M. Claver. Parallel wavefront algorithms solving Lyapunov equations for the Cholesky factor on message passing multiprocessors. The Journal of Supercomputing, 13(2):171– 189, March 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=2&spage=171;   http:
//www.wkap.nl/oasis.htm/200579.

Yang:1999:PPA

[270] Tianruo Yang and Hai-Xiang Lin. Parallel performance analysis of the improved quasi-minimal residual method on bulk synchronous parallel architectures. The Journal of Supercomputing, 13(2):191–210, March 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=2&spage=191;   http:
//www.wkap.nl/oasis.htm/200581.

Coelho:1999:MUB

[271] P. J. Coelho, P. A. Novo, and M. G. Carvalho. Modelling of a utility boiler using parallel computing. The Journal of Supercomputing, 13 (2):211–232, March 1999. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=2&spage=211;   http:
                                        //www.wkap.nl/oasis.htm/200583.

Clement:1999:E

[272] Mark J. Clement and Xian-He Sun. Editorial. The Journal of Supercomputing, 13(3):247, May 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=3&spage=247;   http:
                                        //www.wkap.nl/oasis.htm/207335.

Brorsson:1999:PTS

[273] Mats Brorsson and Martin Kral. Performance tuning software DSM applications using visualisation. The Journal of Supercomputing, 13(3): 249–265, May 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=3&spage=249;   http:
                                        //www.wkap.nl/oasis.htm/206580.

Clement:1999:PSP

[274] Mark J. Clement, Glenn M. Judd, Bryan S. Morse, and J. Kelly Flanagan. Performance surface prediction for WAN-based clusters. The Journal of Supercomputing, 13(3): 267–281, May 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=3&spage=267;   http:
//www.wkap.nl/oasis.htm/206581.

Luo:1999:SMV

[275] Yong Luo. Shared memory vs. message passing: The COMOPS benchmark experiment. The Journal of Supercomputing, 13(3):283–301, May 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=3&spage=283;   http:
//www.wkap.nl/oasis.htm/206582.

Becker:1999:PPP

[276] Jeffrey C. Becker, Bill Nitzberg, Rob F. Van Der Wijngaart, and Maurice Yarrow. Predicting price/ performance trade-offs for Whitney: a commodity computing cluster. The Journal of Supercomputing, 13(3): 303–319, May 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=13&issue=3&spage=303;   http:
//www.wkap.nl/oasis.htm/206583.

Gustafson:1999:CBS

[277] John L. Gustafson and Rajat Todi.

Conventional benchmarks as a sample of the performance spectrum. The Journal of Supercomputing, 13 (3):321–342, May 1999. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=13&issue=3&spage=321;   http:
                                        //www.wkap.nl/oasis.htm/206584.

Zheng:1999:EGP

[278] S. Q. Zheng, Balaji Calidas, and Yanjun Zhang. An efficient general in-place parallel sorting scheme. The Journal of Supercomputing, 14 (1):5–17, July 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=14&issue=1&spage=5;   http:
                                        //www.wkap.nl/oasis.htm/231954.

Avresky:1999:PMS

[279] D. R. Avresky, V. Shurbanov, R. Horst, W. Watson, L. Young, and D. Jewett. Performance modeling of ServerNet^{T M} SAN topologies. The Journal of Supercomputing, 14(1):19–37, July 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=14&issue=1&spage=19;   http:
                                        //www.wkap.nl/oasis.htm/231955.

Lin:1999:NCD

[280] Yen-Chun Lin and Chao-Cheng Shih. A new class of depth-size optimal parallel prefix circuits. The Journal of Supercomputing, 14(1):39–52, July 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=14&issue=1&spage=39;   http:
                                        //www.wkap.nl/oasis.htm/231956.

Sibai:1999:OCH

[281] Fadi N. Sibai. Optimal clustering of hierarchical hyper-ring multicomputers. The Journal of Supercomputing, 14(1):53–76, July 1999. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=1&spage=53;   http:
//www.wkap.nl/oasis.htm/231957.

Wang:1999:MCP

[282] Dianqin Wang and Eleanor Chu. Minimizing communication penalty of triangular solvers by runtime mesh configuration and workload redistribution. The Journal of Supercomputing, 14(1):77–95, July 1999. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=1&spage=77;   http:
//www.wkap.nl/oasis.htm/231958.

Frietman:1999:POI

[283] Edward E. E. Frietman, Ramon J.

Ernst, Roy Crosbie, and Masao Shimoji. Prospects for optical interconnects in distributed, shared-memory organized MIMD architectures. The Journal of Supercomputing, 14(2):107– 128, September/October 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=2&spage=107;   http:
//www.wkap.nl/oasis.htm/233855.

Espasa:1999:SSD

[284] Roger Espasa and Mateo Valero. A simulation study of decoupled vector architectures. The Journal of Supercomputing, 14(2):129–152, September/October 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=2&spage=124;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=2&spage=129;   http:
//www.wkap.nl/oasis.htm/233856.

Poggi:1999:DOO

[285] Agostino Poggi and Giuliano Armano. A distributed object-oriented language for heterogeneous computer networks. The Journal of Supercomputing, 14(2): 153–169, September/October 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=2&spage=153;   http:
//www.wkap.nl/oasis.htm/233857.

Royo:1999:LCO

[286] Dolors Royo, Antonio González, and Miguel Valero-García. Low communication overhead Jacobi algorithms for eigenvalues computation on hypercubes. The Journal of Supercomputing, 14(2): 171–193, September/October 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=2&spage=171;   http:
//www.wkap.nl/oasis.htm/233858.

Ouyang:1999:SCE

[287] Jinsong Ouyang and Piyush Maheshwari. Supporting cost-effective fault tolerance in distributed message-passing applications with file operations. The Journal of Supercomputing, 14(3):207– 232, November/December 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=3&spage=207;   http:
//www.wkap.nl/oasis.htm/237512.

Yip:1999:EPA

[288] Hoi-Man Yip, Ishfaq Ahmad, and Ting-Chuen Pong. An efficient parallel algorithm for computing the Gaussian convolution of multi-dimensional image data. The Journal of Supercomputing, 14(3):233–255, November/December 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=3&spage=233;   http:
//www.wkap.nl/oasis.htm/237514.

Maheswaran:1999:MMC

[289] Muthucumaru Maheswaran, Kevin J.

Webb, and Howard Jay Siegel. MCGS: a modified conjugate gradient squared algorithm for nonsymmetric linear systems. The Journal of Supercomputing, 14(3):257–280, November/December 1999. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=14&issue=3&spage=257;   http:
//www.wkap.nl/oasis.htm/237516.

Olariu:2000:DIP

[290] Stephan Olariu, Ivan Stojmenovic, and Albert Zomaya. On the dynamic initialization of parallel computers. The Journal of Supercomputing, 15(1):5–24, January 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=1&spage=5;   http:
                                        //www.wkap.nl/oasis.htm/243234.

Chung:2000:PCM

[291] Yeh-Ching Chung, Ching-Jung Liao, and Don-Lin Yang. A prefix code matching parallel load-balancing method for solution-adaptive unstructured finite element graphs on distributed memory multicomputers. The Journal of Supercomputing, 15(1):25– 49, January 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=1&spage=25;   http:
                                        //www.wkap.nl/oasis.htm/243235.

Kutluca:2000:ISD

[292] Hüuseyin Kutluca, Tah¨ sin M. Kurç, and Cevdet Aykanat. Image-space decomposition algorithms for sort-first parallel volume rendering of unstructured grids. The Journal of Supercomputing, 15(1):51–93, January 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=1&spage=51;   http:
                                        //www.wkap.nl/oasis.htm/243236.

Imamura:2000:ECC

[293] Toshiyuki Imamura. An estimation of complexity and computational costs for vertical block-cyclic distributed parallel LU factorization. The Journal of Supercomputing, 15 (1):95–110, January 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=15&issue=1&spage=95;   http:
//www.wkap.nl/oasis.htm/243237.

Bourgeois:2000:CPC

[294] J. Bourgeois, F. Spies, M. J. Zemerly, and T. Delaitre. Chronos: a performance characterization tool inside the EDPEPPS toolset. The Journal of Supercomputing, 15(2):123– 140, February 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=15&issue=2&spage=123;   http:
//www.wkap.nl/oasis.htm/250901.

Schnekenburger:2000:LBC

[295] Thomas Schnekenburger. Load balancing in CORBA: a survey of concepts, patterns, and techniques. The Journal of Supercomputing, 15(2):141– 161, February 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=15&issue=2&spage=141;   http:
//www.wkap.nl/oasis.htm/250902.

Li:2000:EDP

[296] Keqin Li, Yi Pan, and Si Qing Zheng. Efficient deterministic and probabilistic simulations of PRAMs on linear arrays with reconfigurable pipelined bus systems. The Journal of Supercomputing, 15(2):163–181, February 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=2&spage=163;   http:
                                        //www.wkap.nl/oasis.htm/250903.

Latifi:2000:WBH

[297] Shahram Latifi, Myung Hoon Lee, and Pradip K. Srimani. Wormhole broadcast in hypercubes. The Journal of Supercomputing, 15(2):183– 192, February 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=2&spage=183;   http:
                                        //www.wkap.nl/oasis.htm/250904.

Benner:2000:PPS

[298] Peter Benner, Maribel Castillo, Enrique S. Quintana-Ortií, and Vicente Hernández. Parallel partial stabilizing algorithms for large linear control systems. The Journal of Supercomputing, 15(2):193–206, February 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=2&spage=193;   http:
                                        //www.wkap.nl/oasis.htm/250905.

Takahashi:2000:HPR

[299] Daisuke Takahashi and Yasumasa Kanada. High-performance radix-2, 3 and 5 parallel 1-D complex FFT algorithms for distributedmemory parallel computers. The Journal of Supercomputing, 15(2):207– 228, February 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=15&issue=2&spage=207;   http:
//www.wkap.nl/oasis.htm/250906.

Shih:2000:SLC

[300] Kuei-Ping Shih, Jang-Ping Sheu, and Chua-Huang Huang. Statement-level communication-free partitioning techniques for parallelizing compilers. The Journal of Supercomputing, 15(3):243– 269, February 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=15&issue=3&spage=243;   http:
//www.wkap.nl/oasis.htm/254453.

Rauber:2000:DAD

[301] Thomas Rauber and Gudula Rünger.

Deriving array distributions by optimization techniques. The Journal of Supercomputing, 15(3):271– 293, February 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=15&issue=3&spage=271;   http:
//www.wkap.nl/oasis.htm/254454.

Park:2000:LOL

[302] Taesoon Park and Heon Y. Yeom. A low overhead logging scheme for fast recovery in distributed shared memory systems. The Journal of Supercomputing, 15(3):295–320, February 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=3&spage=295;   http:
                                        //www.wkap.nl/oasis.htm/254455.

Wu:2000:ITP

[303] Jan-Jan Wu. An interleaving transformation for parallelizing reductions for distributed-memory parallel machines. The Journal of Supercomputing, 15(3):321–339, February 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=15&issue=3&spage=321;   http:
                                        //www.wkap.nl/oasis.htm/254456.

Avresky:2000:EFT

[304] D. R. Avresky, B. W. Johnson, and F. F. Lombardi. Embedded fault-tolerant systems — introduction. The Journal of Supercomputing, 16(1–2):5–6, May 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=16&issue=1&spage=5;   http:
                                        //www.wkap.nl/oasis.htm/257738.

Caldwell:2000:MFT

[305] Douglas W. Caldwell and David A. Rennels. A minimalist fault-tolerant microcontroller design for embedded spacecraft computing. The Journal of Supercomputing, 16(1–2): 7–25, May 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/257739.

Somani:2000:ARM

[306] Arun K. Somani and Allen M. Sansano.

Achieving robustness and minimizing overhead in parallel algorithms through overlapped communication/ computation. The Journal of Supercomputing, 16(1–2):27–52, May 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=1&spage=27;   http:
//www.wkap.nl/oasis.htm/257740.

Haines:2000:ALF

[307] Joshua Haines, Vijay Lakamraju, Israel Koren, and C. Mani Krishna. Application-level fault tolerance as a complement to system-level fault tolerance. The Journal of Supercomputing, 16(1–2):53–68, May 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=1&spage=53;   http:
//www.wkap.nl/oasis.htm/257741.

Lyubashevskiy:2000:FTF

[308] Igor Lyubashevskiy and Volker Strumpen.

Fault-tolerant file-I/O for portable checkpointing systems. The Journal of Supercomputing, 16(1–2):69– 92, May 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=16&issue=1&spage=69;   http:
                                        //www.wkap.nl/oasis.htm/257742.

Alvarez:2000:SBT

[309] Guillermo A. Alvarez and Flaviu Cristian. Simulation-based testing of communication protocols for dependable embedded systems. The Journal of Supercomputing, 16(1–2): 93–116, May 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=16&issue=1&spage=93;   http:
                                        //www.wkap.nl/oasis.htm/257743.

Makki:2000:ULR

[310] Kia Makki, John Dell, Niki Pissinou, W. Melody Moh, and Xiaohua Jia. Using logical rings to solve the distributed mutual exclusion problem with fault tolerance issues. The Journal of Supercomputing, 16(1–2):117–132, May 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=16&issue=1&spage=117;   http:
                                        //www.wkap.nl/oasis.htm/257744.

Hinton:2000:IFA

[311] H. Scott Hinton. Introduction to focus articles on optical computing. The Journal of Supercomputing, 16 (3):147, July 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=3&spage=147;   http:
//www.wkap.nl/oasis.htm/262634.

Liu:2000:DPA

[312] Guoping Liu, Kyungsook Y. Lee, and Harry F. Jordan. n-dimensional processor arrays with optical dBuses. The Journal of Supercomputing, 16 (3):149–163, July 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=3&spage=149;   http:
//www.wkap.nl/oasis.htm/262635.

Love:2000:OMO

[313] Carl E. Love and Harry F. Jordan. Oscillatory metastability in optical network synchronizer circuits. The Journal of Supercomputing, 16 (3):165–176, July 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=3&spage=165;   http:
//www.wkap.nl/oasis.htm/262636.

Fey:2000:DPA

[314] Dietmar Fey and Marko Degenkolb.

Digit pipelined arithmetic for 3-D massively parallel optoelectronic circuits. The Journal of Supercomputing, 16(3):177–196, July 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&

volume=16&issue=3&spage=177;   http:
//www.wkap.nl/oasis.htm/262637.

Kim:2000:IBI

[315] Sunil Kim and Alexander V. Veidenbaum. On interaction between interconnection network design and latency hiding techniques in multiprocessors. The Journal of Supercomputing, 16(3):197–216, July 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=3&spage=197;   http:
//www.wkap.nl/oasis.htm/262638.

Koita:2000:MCS

[316] Takahiro Koita, Tetsuro Katayama, Keizo Saisho, and Akira Fukuda. Memory conscious scheduling for clusterbased NUMA multiprocessors. The Journal of Supercomputing, 16(3): 217–235, July 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=3&spage=217;   http:
//www.wkap.nl/oasis.htm/262640.

Fan:2000:TGM

[317] Kuo-Pao Fan and Chung-Ta King. Turn grouping for multicast in wormholerouted mesh networks supporting the turn model. The Journal of Supercomputing, 16(3):237–260, July 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=16&issue=3&spage=237;   http:
//www.wkap.nl/oasis.htm/262641.

Li:2000:PAV

[318] Hongzheng Li and Hongchi Shi. A parallel algorithm for volume projections on SIMD mesh-connected computers. The Journal of Supercomputing, 17(1):5–22, August 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/266088.

Hsu:2000:EMM

[319] Ching-Hsien Hsu, Yeh-Ching Chung, and Chyi-Ren Dow. Efficient methods for multi-dimensional array redistribution. The Journal of Supercomputing, 17(1):23–46, August 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=1&spage=23;   http:
//www.wkap.nl/oasis.htm/266089.

Gupta:2000:UND

[320] Sandeep K. S. Gupta and Pradip K.

Srimani. UpdateSearch: a new dynamic channel allocation scheme for mobile networks that can adjust to system loads. The Journal of Supercomputing, 17(1):47–65, August 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=1&spage=47;   http:
//www.wkap.nl/oasis.htm/266090.

Hoganson:2000:MPA

[321] Kenneth Hoganson. Mapping parallel application communication topology to rhombic overlapping-cluster multiprocessors. The Journal of Supercomputing, 17(1):67–90, August 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=1&spage=67;   http:
//www.wkap.nl/oasis.htm/266091.

Yang:2000:DDO

[322] Don-Lin Yang, Yeh-Ching Chung, Chih-Chang Chen, and Ching-Jung Liao. A dynamic diffusion optimization method for irregular finite element graph partitioning. The Journal of Supercomputing, 17(1):91–110, August 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=1&spage=91;   http:
//www.wkap.nl/oasis.htm/266093.

Averbuch:2000:EPT

[323] A. Averbuch, B. Epstein, L. Ioffe, and I. Yavneh. Efficient parallelization of a three-dimensional Navier– Stokes solver on MIMD multiprocessors. The Journal of Supercomputing, 17(2):123–142, January 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=2&spage=123;   http:
//www.wkap.nl/oasis.htm/271242.

Ben-Asher:2000:BRA

[324] Yosi Ben-Asher and Esti Stein. Basic results in automatic transformations of shared memory parallel programs into sequential programs. The Journal of Supercomputing, 17(2):143– 165, January 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=2&spage=143;   http:
//www.wkap.nl/oasis.htm/271244.

Melab:2000:PAG

[325] N. Melab, E.-G. Talbi, and S. Petiton.

A parallel adaptive Gauss–Jordan algorithm. The Journal of Supercomputing, 17(2):167–185, January 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=2&spage=167;   http:
//www.wkap.nl/oasis.htm/271245.

Chang:2000:IMT

[326] Chih-Yung Chang, Tzung-Shi Chen, and Jang-Ping Sheu. Improving memory traffic by assembly-level exploitation of reuses for vector registers. The Journal of Supercomputing, 17 (2):187–204, January 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=2&spage=187;   http:
//www.wkap.nl/oasis.htm/271248.

Shih:2000:EAG

[327] Kuei-Ping Shih, Jang-Ping Sheu, and Chih-Yung Chang. Efficient address generation for affine subscripts in data-parallel programs. The Journal of Supercomputing, 17(2):205– 227, January 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=17&issue=2&spage=205;   http:
                                        //www.wkap.nl/oasis.htm/271249.

DiMartino:2000:ITT

[328] Beniamino Di Martino and Antonino Mazzeo. Introduction: Techniques and tools for parallel and distributed program analysis, development and run-time support. The Journal of Supercomputing, 17(3):243– 244, November 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=17&issue=3&spage=243;   http:
                                        //www.wkap.nl/oasis.htm/277524.

Kessler:2000:NNP

[329] Christoph W. Keßler. NestStep: Nested parallelism and virtual shared memory for the BSP model. The Journal of Supercomputing, 17(3):245– 262, November 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=17&issue=3&spage=245;   http:
                                        //www.wkap.nl/oasis.htm/277525.

Bandera:2000:CRT

[330] Gerardo Bandera, Manuel Ujaldón, and Emilio L. Zapata. Compile and runtime support for the parallelization of sparse matrix updating algorithms. The Journal of Supercomputing, 17(3): 263–276, November 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=3&spage=263;   http:
//www.wkap.nl/oasis.htm/277526.

Wismuller:2000:IRT

[331] Roland Wismüller and Thomas Ludwig. Interoperable run-time tools for distributed systems — a case study. The Journal of Supercomputing, 17(3): 277–289, November 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=3&spage=277;   http:
//www.wkap.nl/oasis.htm/277527.

Girona:2000:SPP

[332] Sergi Girona and Jesús Labarta. Sensitivity of performance prediction of message passing programs. The Journal of Supercomputing, 17(3):291– 298, November 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=3&spage=291;   http:
//www.wkap.nl/oasis.htm/277529.

Aversa:2000:RPP

[333] Rocco Aversa, Beniamino Di Martino, Nicola Mazzocca, and Umberto Villano. Reducing parallel program simulation complexity by static analysis. The Journal of Supercomputing, 17(3): 299–310, November 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&

volume=17&issue=3&spage=299;   http:
//www.wkap.nl/oasis.htm/277530.

Chapman:2000:PDT

[334] B. Chapman, J. Merlin, D. Pritchard, F. Bodin, Y. Mevel, T. Sørevik, and L. Hill. Program development tools for clusters of shared memory multiprocessors. The Journal of Supercomputing, 17(3):311–322, November 2000. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=17&issue=3&spage=311;   http:
//www.wkap.nl/oasis.htm/277531.

Alme:2001:DDM

[335] Henry J. Alme, Garry H. Rodrigue, and George B. Zimmerman. Domain decomposition models for parallel Monte Carlo transport. The Journal of Supercomputing, 18(1):5–23, January 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/319612.

Tsaur:2001:ACR

[336] Woei-Jiunn Tsaur and Shi-Jinn Horng. Auditing causal relationships of group multicast communications in grouporiented distributed systems. The Journal of Supercomputing, 18(1):25– 45, January 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=1&spage=25;   http:
//www.wkap.nl/oasis.htm/319614.

Morrison:2001:WWB

[337] John P. Morrison, James J. Kennedy, and David A. Power. WebCom: a Web based volunteer computer. The Journal of Supercomputing, 18 (1):47–61, January 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=1&spage=47;   http:
//www.wkap.nl/oasis.htm/319615.

Hsiao:2001:ENL

[338] Hung-Chang Hsiao and Chung-Ta King. Exploiting network locality for CC-NUMA multiprocessors. The Journal of Supercomputing, 18(1):63– 87, January 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=1&spage=63;   http:
//www.wkap.nl/oasis.htm/319617.

Pereira:2001:SSI

[339] P. Pereira, L. Heutte, and Y. Lecourtier.

Source-to-source instrumentation for the optimization of an automatic reading system. The Journal of Supercomputing, 18(1):89–104, January 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=1&spage=89;   http:
//www.wkap.nl/oasis.htm/319618.

Shen:2001:AIG

[340] Hong Shen, Sam Lor, and Piyush Maheshwari. An architecture-independent graphical tool for automatic contentionfree process-to-processor mapping. The Journal of Supercomputing, 18(2):115– 139, February 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=2&spage=115;   http:
//www.wkap.nl/oasis.htm/320462.

Gonzalez:2001:PCW

[341] Patricia González, José C. Cabaleiro, and Tomás F. Pena. Parallel computation of wavelet transforms using the lifting scheme. The Journal of Supercomputing, 18(2):141– 152, February 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=2&spage=141;   http:
//www.wkap.nl/oasis.htm/320463.

Hadim:2001:MLC

[342] M. B. Hadim and I. Sakho. The multi-level communication: Efficient routing for interconnection networks. The Journal of Supercomputing, 18 (2):153–177, February 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=2&spage=153;   http:
//www.wkap.nl/oasis.htm/320464.

Cha:2001:HBH

[343] Hojung Cha and Dongho Lee. H-BSP: a Hierarchical BSP computation model. The Journal of Supercomputing, 18(2):179–200, February 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=2&spage=179;   http:
//www.wkap.nl/oasis.htm/320465.

Yang:2001:ECM

[344] Don-Lin Yang, Jen-Chih Yu, and Yeh-Ching Chung. Efficient compositing methods for the sort-last-sparse parallel volume rendering system on distributed memory multicomputers. The Journal of Supercomputing, 18 (2):201–220, February 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=2&spage=201;   http:
//www.wkap.nl/oasis.htm/320466.

Elmroth:2001:HPC

[345] Erik Elmroth, Chris Ding, and Yu-Shu Wu. High performance computations for large scale simulations of subsurface multiphase fluid and heat flow. The Journal of Supercomputing, 18(3):235–258, March 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=18&issue=3&spage=235;   http:
//www.wkap.nl/oasis.htm/323811.

He:2001:UAA

[346] Yun He and Chris H. Q. Ding. Using accurate arithmetics to improve numerical reproducibility and stability in parallel applications. The Journal of Supercomputing, 18(3):259–277, March 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=18&issue=3&spage=259;   http:
                                        //www.wkap.nl/oasis.htm/323815.

Hsiao:2001:ADS

[347] Hung-Chang Hsiao and Chung-Ta King. An application-driven study of multicast communication for write invalidation. The Journal of Supercomputing, 18(3):279–304, March 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=18&issue=3&spage=279;   http:
                                        //www.wkap.nl/oasis.htm/323816.

Chang:2001:PSS

[348] Rong-Guey Chang, Tyng-Ruey Chuang, and Jenq Kuen Lee. Parallel sparse supports for array intrinsic functions of Fortran 90. The Journal of Supercomputing, 18(3):305–339, March 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=18&issue=3&spage=305;   http:
                                        //www.wkap.nl/oasis.htm/323817.

Plaks:2001:ERH

[349] Toomas P. Plaks and Graham M. Megson. Engineering of reconfigurable hardware/software objects. The Journal of Supercomputing, 19 (1):5–6, May 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&

volume=19&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/333394.

Susanto:2001:FAD

[350] Kong Woei Susanto and Tom Melham.

Formally analyzed dynamic synthesis of hardware. The Journal of Supercomputing, 19(1):7–22, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/333395.

Bowen:2001:ASV

[351] Jonathan P. Bowen and He Jifeng.

An approach to the specification and verification of a hardware compilation scheme. The Journal of Supercomputing, 19(1):23–39, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=23;   http:
//www.wkap.nl/oasis.htm/333396.

Fimmel:2001:DPA

[352] Dirk Fimmel and Renate Merker. Design of processor arrays for reconfigurable architectures. The Journal of Supercomputing, 19(1):41–56, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=41;   http:
//www.wkap.nl/oasis.htm/333397.

Teich:2001:ODH

[353] Jürgen Teich, Sándor P. Fekete, and Jörg Schepers. Optimization of dynamic hardware reconfigurations. The Journal of Supercomputing, 19(1): 57–75, May 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=57;   http:
//www.wkap.nl/oasis.htm/333398.

Liang:2001:DBA

[354] Xuejun Liang, Jack Jean, and Karen Tomko. Data buffering and allocation in mapping generalized template matching on reconfigurable systems. The Journal of Supercomputing, 19(1):77–91, May 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=77;   http:
//www.wkap.nl/oasis.htm/333404.

McEwan:2001:HSR

[355] Alistair A. McEwan and Jonathan Saul. A high speed reconfigurable firewall based on parameterizable FPGAbased content addressable memories. The Journal of Supercomputing, 19 (1):93–103, May 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=93;   http:
//www.wkap.nl/oasis.htm/333400.

Becker:2001:PDRa

[356] Jürgen Becker and Manfred Glesner. A parallel dynamically reconfigurable architecture designed for flexible applicationtailored hardware/software systems in future mobile communication. The Journal of Supercomputing, 19(1): 105–127, May 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=1&spage=105;   http:
//www.wkap.nl/oasis.htm/333401.

Wu:2001:PNR

[357] Jie Wu and Yuanyuan Yang. The postal network: a recursive network for parameterized communication model. The Journal of Supercomputing, 19 (2):143–161, June 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=2&spage=143;   http:
//www.wkap.nl/oasis.htm/336451.

Vigo-Aguiar:2001:POS

[358] J. Vigo-Aguiar and L. M. Quintales. A parallel ODE solver adapted to oscillatory problems. The Journal of Supercomputing, 19(2):163– 171, June 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=2&spage=163;   http:
//www.wkap.nl/oasis.htm/336452.

Wang:2001:BIS

[359] L. Wang and Ted C. Yang. On the boosting of instruction scheduling by renaming. The Journal of Supercomputing, 19(2):173–197, June 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=19&issue=2&spage=173;   http:
                                        //www.wkap.nl/oasis.htm/337134.

Bhalla:2001:PED

[360] Subhash Bhalla. The performance of an efficient distributed synchronization and recovery algorithm. The Journal of Supercomputing, 19(2): 199–219, June 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=19&issue=2&spage=199;   http:
                                        //www.wkap.nl/oasis.htm/336453.

Akl:2001:ISQ

[361] Selim G. Akl and Stefan D. Bruda. Improving a solution’s quality through parallel processing. The Journal of Supercomputing, 19(2):221– 233, June 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=19&issue=2&spage=221;   http:
                                        //www.wkap.nl/oasis.htm/336454.

Vedantham:2001:SIG

[362] Sundararajan Vedantham, Rajkumar Buyya, and Mahbub Hassan. Special issue guest editorial: Cluster computing using high-speed networks. The Journal of Supercomputing, 19 (3):247–249, July 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&

volume=19&issue=3&spage=247;   http:
//www.wkap.nl/oasis.htm/338200.

Chung:2001:IRS

[363] Sang-Hwa Chung, Hyuk-Chul Kwon, Kwang Ryel Ryu, Yoojin Chung, Hankook Jang, and Cham-Ah Choi. Information retrieval on an SCI-based PC cluster. The Journal of Supercomputing, 19(3):251–265, July 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=3&spage=251;   http:
//www.wkap.nl/oasis.htm/338202.

Kim:2001:OCH

[364] Hyeong-Do Kim and Chang-Sung Jeong. Object clustering for high performance parallel computing. The Journal of Supercomputing, 19(3): 267–283, July 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=3&spage=267;   http:
//www.wkap.nl/oasis.htm/338203.

Hawick:2001:ATM

[365] K. A. Hawick and H. A. James. Asynchronous transfer mode and other network technologies for wide-area and high-performance cluster computing. The Journal of Supercomputing, 19 (3):285–297, July 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=19&issue=3&spage=285;   http:
//www.wkap.nl/oasis.htm/338204.

Kwok:2001:FTP

[366] Yu-Kwong Kwok. Fault-tolerant parallel scheduling of tasks on a heterogeneous high-performance workstation cluster. The Journal of Supercomputing, 19(3):299–314, July 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=19&issue=3&spage=299;   http:
                                        //www.wkap.nl/oasis.htm/338206.

Takeda:2001:AME

[367] K. Takeda, N. K. Allsopp, J. C. Hardwick, P. C. Macey, D. A. Nicole, S. J. Cox, and D. J. Lancaster. An assessment of MPI environments for Windows NT. The Journal of Supercomputing, 19(3):315–323, July 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=19&issue=3&spage=315;   http:
                                        //www.wkap.nl/oasis.htm/338207.

Antonopoulos:2001:AOS

[368] Nick Antonopoulos and Alex Shafarenko. An active organisation system for customised, secure agent discovery. The Journal of Supercomputing, 20(1):5–35, August 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=20&issue=1&spage=5;   http:
                                        //www.wkap.nl/oasis.htm/350763.

Al-Ayyoub:2001:PPM

[369] A. Al-Ayyoub, M. Ould-Khaoua, and K. Day. On the performance of parallel matrix factorisation on the hypermesh. The Journal of Supercomputing, 20(1):37–53, August 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=20&issue=1&spage=37;   http:
//www.wkap.nl/oasis.htm/350764.

Quintana-Orti:2001:EAB

[370] Enrique S. Quintana-Ortí, Gregorio Quintana-Ortí, Maribel Castillo, and Vicente Hernández. Efficient algorithms for the block Hessenberg form. The Journal of Supercomputing, 20 (1):55–66, August 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=20&issue=1&spage=55;   http:
//www.wkap.nl/oasis.htm/350765.

Chang:2001:CFA

[371] Weng-Long Chang, Chih-Ping Chu, and Jia-Hwa Wu. Communicationfree alignment for array references with linear subscripts in three loop index variables or quadratic subscripts. The Journal of Supercomputing, 20 (1):67–83, August 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=20&issue=1&spage=67;   http:
//www.wkap.nl/oasis.htm/350766.

Tsaoussidis:2001:EC

[372] Vassilis Tsaoussidis. Editorial comments. The Journal of Supercomputing, 20(2):99–100, September 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/28/1/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/28/1/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=2&spage=
99.

Batsiolas:2001:SIE

[373] I. Batsiolas and Ioanis Nikolaidis. Selective idling: Experiments in transport layer energy conservation. The Journal of Supercomputing, 20(2):101– 114, September 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/28/2/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/28/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=2&spage=
101.

Tsaoussidis:2001:WPC

[374] Vassilis Tsaoussidis, A. Lahanas, and C. Zhang. The wave & probe communication mechanisms. The Journal of Supercomputing, 20(2):115– 135, September 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/28/3/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/28/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=2&spage=
115.

Mitzenmacher:2001:TMC

[375] Michael Mitzenmacher and Rajmohan Rajaraman. Towards more complete models of TCP latency and throughput. The Journal of Supercomputing, 20(2):137–160, September 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/28/4/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/28/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=2&spage=
137.

Langendorfer:2001:EWK

[376] Peter Langendörfer, Rolf Kraemer, and Hartmut König. Evaluation of wellknown protocol implementation techniques for application in wireless networks. The Journal of Supercomputing, 20(2):161–173, September 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/28/5/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/28/5/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=2&spage=
161.

Markovski:2001:SAP

[377] Velibor Markovski, Fei Xue, and Ljiljana Trajković. Simulation and analysis of packet loss in user datagram protocol transfers. The Journal of Supercomputing, 20(2):175– 196, September 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/28/6/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/28/6/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=2&spage=
175.

Wang:2001:LCF

[378] Dajin Wang. A low-cost fault-tolerant structure for the hypercube. The Journal of Supercomputing, 20(3):203– 216, November 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/29/1/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/29/1/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=3&spage=
203.

Wang:2001:CSB

[379] San-Yuan Wang, Yu-Chee Tseng, Sze-Yao Ni, and Jang-Ping Sheu. Circuitswitched broadcasting in multi-port multi-dimensional torus networks. The Journal of Supercomputing, 20(3):217– 241, November 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/29/2/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/29/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=3&spage=
217.

Guo:2001:FED

[380] Minyi Guo and Ikuo Nakata. A framework for efficient data redistribution on distributed memory multicomputers. The Journal of Supercomputing, 20(3):243–265, November 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/29/3/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/29/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=20&issue=3&spage=
243.

Oldehoeft:2001:ISH

[381] Rod Oldehoeft. Introduction to software for high-performance systems papers from the Symposium of The Los Alamos Computer Science Institute. The Journal of Supercomputing, 23(1):5–6, May 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=5.

Cooper:2001:AOC

[382] Keith D. Cooper, Devika Subramanian, and Linda Torczon. Adaptive optimizing compilers for the 21st Century. The Journal of Supercomputing, 23(1):7–22, May 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=7.

Deitz:2001:HLL

[383] Steven J. Deitz, Bradford L. Chamberlain, and Lawrence Snyder. High-level language support for user-defined reductions. The Journal of Supercomputing, 23(1):23–37, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=23.

Cummings:2001:VTF

[384] Julian Cummings, Michael Aivazis, Ravi Samtaney, Raul Radovitzky, Sean Mauch, and Dan Meiron. A virtual test facility for the simulation of dynamic response in materials. The Journal of Supercomputing, 23(1):39–50, May 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=39.

Feng:2001:PSE

[385] Annette C. Feng, Apu C. Kapadia, Wu-Chun Feng, and Geneva G. Belford. Packet spacing: An enabling mechanism for delivering multimedia content in computational grids. The Journal of Supercomputing, 23 (1):51–66, May 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=51.

Feng:2001:MTD

[386] Wu-Chun Feng, Mark K. Gardner, and Jeffrey R. Hay. The MAGNeT toolkit: Design, implementation and evaluation. The Journal of Supercomputing, 23(1):67–79, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=67.

Mellor-Crummey:2001:HTT

[387] John Mellor-Crummey, Robert J.

Fowler, Gabriel Marin, and Nathan Tallent. HPCVIEW: a tool for topdown analysis of node performance. The Journal of Supercomputing, 23 (1):81–104, May 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=81.

Anonymous:2001:EDP

[388] Anonymous. Erratum: Design and prototype of a performance tool interface for OpenMP. The Journal of Supercomputing, 23(1):105–128, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=105. See

[429].

Becker:2001:PDRb

[389] Jürgen Becker, Ahmad Alsolaim, Manfred Glesner, and Janusz Starzyk. A parallel dynamically reconfigurable architecture designed for flexible applicationtailored hardware/software systems in future mobile communication. The Journal of Supercomputing, 23(1): 132, May 2001. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=132.

Chung:2002:APD

[390] Soon M. Chung and Arindam Chatterjee. An adaptive parallel distributive join algorithm on a cluster of workstations. The Journal of Supercomputing, 21(1):5–35, January 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/30/1/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/30/1/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=1&spage=
5.

Loechner:2002:PDL

[391] Vincent Loechner, Benoît Meister, and Philippe Clauss. Precise data locality optimization of nested loops. The Journal of Supercomputing, 21 (1):37–76, January 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/30/2/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/30/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=1&spage=
37.

Myoupo:2002:OBS

[392] Jean-Frédéric Myoupo, David Semé, and Ivan Stojmenovic. Optimal BSR solutions to several convex polygon problems. The Journal of Supercomputing, 21(1):77–90, January 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/30/3/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/30/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=1&spage=
77.

Xu:2002:SPE

[393] Cheng-Zhong Xu, Le Yi Wang, and Ngo-Tai Fong. Stochastic prediction of execution time for dynamic bulk synchronous computations. The Journal of Supercomputing, 21(1):91– 103, January 2002. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/30/4/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/30/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=1&spage=
91.

Plaks:2002:GEF

[394] Toomas P. Plaks. Guest Editor’s foreword: Engineering of reconfigurable systems. The Journal of Supercomputing, 21(2):115–116, February 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/31/1/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/31/1/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=

                                        0920-8542&volume=21&issue=2&spage=
                                        115.

Bohm:2002:MSA

[395] W. Böhm, J. Hammes, B. Draper, M. Chawathe, C. Ross, R. Rinker, and W. Najjar. Mapping a single assignment programming language to reconfigurable systems. The Journal of Supercomputing, 21(2):117–130, February 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/31/2/abstract.htm;
                                        http://ipsapp008.lwwonline.com/
                                        content/getfile/5189/31/2/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0920-8542&volume=21&issue=2&spage=
                                        117.

Ha:2002:BVF

[396] Yajun Ha, Serge Vernalde, Partrick Schaumont, Marc Engels, Rudy Lauwereins, and Hugo De Man. Building a virtual framework for networked reconfigurable hardware and software objects. The Journal of Supercomputing, 21(2):131–144, February 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/31/3/abstract.htm;
                                        http://ipsapp008.lwwonline.com/
                                        content/getfile/5189/31/3/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0920-8542&volume=21&issue=2&spage=
                                        131.

Eisenring:2002:FRT

[397] Michael Eisenring and Marco Platzner. A framework for run-time reconfigurable systems. The Journal of Supercomputing, 21(2):145–159, February 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/31/4/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/31/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=2&spage=
145.

Plaks:2002:CAM

[398] Toomas P. Plaks. Configuring of algorithms in mapping into hardware. The Journal of Supercomputing, 21 (2):161–177, February 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/31/5/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/31/5/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=2&spage=
161.

Koch:2002:CAC

[399] Andreas Koch. Compilation for adaptive computing systems using complex parameterized hardware objects. The Journal of Supercomputing, 21 (2):179–190, February 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/31/6/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/31/6/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=2&spage=

Chen:2002:SDP

[400] Tzung-Shi Chen and Chih-Yung Chang. Skewed data partition and alignment techniques for compiling programs on distributed memory multicomputers. The Journal of Supercomputing, 21 (2):191–211, February 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/31/7/abstract.htm;
                                        http://ipsapp008.lwwonline.com/
                                        content/getfile/5189/31/7/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0920-8542&volume=21&issue=2&spage=
                                        191.

Min:2002:CSS

[401] Geyong Min and Mohamed Ould-Khaoua. A comparative study of switching methods in multicomputer networks. The Journal of Supercomputing, 21(3):227–238, March 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/32/1/abstract.htm;
                                        http://ipsapp008.lwwonline.com/
                                        content/getfile/5189/32/1/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0920-8542&volume=21&issue=3&spage=
                                        227.

Chowdhury:2002:DDD

[402] Abdur Chowdhury, Ophir Frieder, and Peng-Jun Wan. On the design, development, deployment, and network survivability analysis of the dynamic routing system protocol. The Journal of Supercomputing, 21(3):239– 256, March 2002. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/32/2/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/32/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=3&spage=
239.

Kandemir:2002:CTS

[403] Mahmut Kandemir, Alok Choudhary, and J. Ramanujam. An I/O-conscious tiling strategy for disk-resident data sets. The Journal of Supercomputing, 21(3):257–284, March 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/32/3/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/32/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=21&issue=3&spage=
257.

Salleh:2002:ESA

[404] Shaharuddin Salleh, Bahrom Sanugi, Hishamuddin Jamaluddin, Stephan Olariu, and Albert Y. Zomaya. Enhanced simulated annealing technique for the single-row routing problem. The Journal of Supercomputing, 21 (3):285–302, March 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/32/4/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/32/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=

                                        0920-8542&volume=21&issue=3&spage=
                                        285.

Anonymous:2002:GEE

[405] Anonymous. Guest Editors’ editorial. The Journal of Supercomputing, 22(1):5–6, May 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/33/1/abstract.htm;
                                        http://ipsapp008.lwwonline.com/
                                        content/getfile/5189/33/1/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0920-8542&volume=22&issue=1&spage=
                                        5.

Stone:2002:PSS

[406] Christopher Stone and Suresh Menon. Parallel simulations of swirling turbulent flames. The Journal of Supercomputing, 22(1):7–28, May 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/33/2/abstract.htm;
                                        http://ipsapp008.lwwonline.com/
                                        content/getfile/5189/33/2/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0920-8542&volume=22&issue=1&spage=
                                        7.

Meng:2002:NSO

[407] Sha Meng, Xin Kai Li, and Gwynne Evans. Numerical simulation of Oldroyd-B fluid in a contraction channel. The Journal of Supercomputing, 22(1):29–43, May 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp008.lwwonline.com/content/
                                        getfile/5189/33/3/abstract.htm;

http://ipsapp008.lwwonline.com/
content/getfile/5189/33/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
29.

Wong:2002:MCP

[408] Kwai L. Wong and A. J. Baker. A modular collaborative parallel CFD workbench. The Journal of Supercomputing, 22(1):45–53, May 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/33/4/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/33/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
45.

Pascoe:2002:CGM

[409] James S. Pascoe, Roger J. Loader, and Vaidy S. Sunderam. Collaborative group membership. The Journal of Supercomputing, 22(1):55–68, May 2002. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/33/5/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/33/5/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
55.

Rodionov:2002:PSU

[410] Alexei S. Rodionov, Hyunseung Choo, and Hee Y. Youn. Process simulation using randomized Markov chain and truncated marginal distribution.

The Journal of Supercomputing, 22 (1):69–85, May 2002. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/33/6/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/33/6/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
69.

Gavrilova:2002:NNP

[411] Marina L. Gavrilova. On a nearestneighbor problem under Minkowski and power metrics for large data sets. The Journal of Supercomputing, 22(1):87–98, May 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/
getfile/5189/33/7/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/33/7/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
87.

Moret:2002:HPA

[412] Bernard M. E. Moret, David A. Bader, and Tandy Warnow. Highperformance algorithm engineering for computational phylogenetics. The Journal of Supercomputing, 22(1): 99–111, May 2002. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/33/8/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/33/8/fulltext.
pdf;   http://www.springerlink.

com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
99.

Tan:2002:SSL

[413] Chih Jeng Kenneth Tan. Solving systems of linear equations with relaxed Monte Carlo method. The Journal of Supercomputing, 22(1): 113–123, May 2002. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.lwwonline.com/content/
getfile/5189/33/9/abstract.htm;
http://ipsapp008.lwwonline.com/
content/getfile/5189/33/9/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=22&issue=1&spage=
113.

Sarbazi-Azad:2002:PMA

[414] Hamid Sarbazi-Azad, Mohamed Ould-Khaoua, and Lewis M. Mackenzie. A performance model of adaptive wormhole routing in k-ary n-cubes in the presence of digit-reversal traffic. The Journal of Supercomputing, 22(2):139– 159, June 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/34/1/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/34/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=2&spage=139.

Shurbanov:2002:FCS

[415] Vladimir Shurbanov, Dimiter Avresky, Pankaj Mehra, and W. Watson. Flow control in ServerNet(R) clusters. The Journal of Supercomputing, 22(2):161– 173, June 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/34/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/34/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=2&spage=161.

Ahn:2002:EGC

[416] Jinho Ahn, Sung-Gi Min, Chong-Sun Hwang, and Heonchang Yu. Efficient garbage collection schemes for causal message logging with independent checkpointing. The Journal of Supercomputing, 22(2):175–196, June 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/34/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/34/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=2&spage=175.

Chang:2002:RCC

[417] Chih-Yung Chang, Jang-Ping Sheu, and Hsi-Chiuen Chen. Reducing cache conflicts by multi-level cache partitioning and array elements mapping. The Journal of Supercomputing, 22(2):197– 219, June 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/34/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/34/4/fulltext.pdf;

http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=2&spage=197.

Wallcraft:2002:CCA

[418] Alan J. Wallcraft. A comparison of Co-Array Fortran and OpenMP Fortran for SPMD programming. The Journal of Supercomputing, 22(3):231– 250, July 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/36/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/36/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=3&spage=231.

Yook:2002:SGA

[419] Hyun-Gyoo Yook and Myong-Soon Park. Scheduling GEN BLOCK array redistribution. The Journal of Supercomputing, 22(3):251–267, July 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/36/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/36/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=3&spage=251.

Bahig:2002:PSI

[420] Hazem M. Bahig, Sameh S. Daoud, and Mahmoud K. A. Khairat. Parallel self-index integer sorting. The Journal of Supercomputing, 22(3):269– 275, July 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/36/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/36/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=3&spage=269.

Lin:2002:PSW

[421] Ching-Feng Lin, Don-Lin Yang, and Yeh-Ching Chung. Parallel shear-warp factorization volume rendering using efficient 1-D and 2-D partitioning schemes for distributed memory multicomputers. The Journal of Supercomputing, 22(3): 277–302, July 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/36/4/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/36/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=22&issue=3&spage=277.

Oldehoeft:2002:ISH

[422] Rod Oldehoeft. Introduction to Software for High-Performance Systems Papers from the Symposium of The Los Alamos Computer Science Institute. The Journal of Supercomputing, 23(1): 5–6, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/37/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/37/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=5.

Cooper:2002:AOC

[423] Keith D. Cooper, Devika Subramanian, and Linda Torczon. Adaptive optimizing compilers for the 21st Century. The Journal of Supercomputing, 23(1): 7–22, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/37/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/37/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=7.

Deitz:2002:HLL

[424] Steven J. Deitz, Bradford L. Chamberlain, and Lawrence Snyder. Highlevel language support for user-defined reductions. The Journal of Supercomputing, 23(1):23–37, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/37/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/37/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=23.

Cummings:2002:VTF

[425] Julian Cummings, Michael Aivazis, Ravi Samtaney, Raul Radovitzky, Sean Mauch, and Dan Meiron. A virtual test facility for the simulation of dynamic response in materials. The Journal of Supercomputing, 23(1):39– 50, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.

                                        kluweronline.com/content/getfile/
                                        5189/37/4/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/37/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=39.

Feng:2002:PSE

[426] Annette C. Feng, Apu C. Kapadia, Wu-Chun Feng, and Geneva G. Belford. Packet spacing: An enabling mechanism for delivering multimedia content in computational grids. The Journal of Supercomputing, 23(1):51– 66, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        5189/37/5/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/37/5/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=51.

Feng:2002:MTD

[427] Wu-Chun Feng, Mark K. Gardner, and Jeffrey R. Hay. The MAGNeT toolkit: Design, implementation and evaluation. The Journal of Supercomputing, 23(1): 67–79, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        5189/37/6/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/37/6/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=67.

Mellor-Crummey:2002:HTT

[428] John Mellor-Crummey, Robert J.

Fowler, Gabriel Marin, and Nathan Tallent. HPCVIEW: a tool for top-down analysis of node performance. The Journal of Supercomputing, 23(1):81– 104, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/37/7/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/37/7/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=81.

Mohr:2002:DPP

[429] Bernd Mohr, Allen D. Malony, Sameer Shende, and Felix Wolf. Design and prototype of a performance tool interface for OpenMP. The Journal of Supercomputing, 23(1):105–128, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/37/8/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/37/8/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=1&spage=105. See

erratum [388].

Anonymous:2002:E

[430] Anonymous. Erratum. The Journal of Supercomputing, 23(1):132, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/37/9/abstract.htm;   http://

                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/37/9/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=1&spage=132.

Gray:2002:CMI

[431] Paul A. Gray and Vaidy S. Sunderam. Collaborative metacomputing with IceT. The Journal of Supercomputing, 23(2):139–166, September 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        5189/38/1/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/38/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=2&spage=139.

deKloe:2002:FIP

[432] Jos de Kloe, Aad van der Steen, Hakan Öksüzoğlu, and Henk Dijkstra. A fully implicit parallel ocean model using MUMPS. The Journal of Supercomputing, 23(2):167–183, September 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        5189/38/2/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/38/2/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=2&spage=167.

Maldonado:2002:OHO

[433] G. Ivan Maldonado and Naveen Kondapalli. Online higher-order error correction of nonlinear diffusion generalized perturbation theory using neural networks. The Journal of Supercomputing, 23(2):185–192, September 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/38/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/38/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=2&spage=185.

Datta:2002:EGT

[434] Amitava Datta. Efficient graphtheoretic algorithms on a linear array with a reconfigurable pipelined bus system. The Journal of Supercomputing, 23(2):193–211, September 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/38/4/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/38/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=2&spage=193.

Anonymous:2002:CA

[435] Anonymous. Contributing authors.

The Journal of Supercomputing, 23(2): 213–214, September 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.kluweronline.com/content/
getfile/5189/38/5/abstract.htm;
 http://ipsapp008.kluweronline.
com/content/getfile/5189/38/5/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=

                                        0920-8542&volume=23&issue=2&spage=
                                        213.

Langendorfer:2002:EC

[436] Peter Langendörfer. Editorial comments. The Journal of Supercomputing, 23(3):223–224, November 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        5189/39/1/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/5189/39/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=3&spage=223.

Pan:2002:TPB

[437] Jianping Pan, Jon W. Mark, and Sherman X. Shen. TCP performance and behaviors with local retransmissions. The Journal of Supercomputing, 23(3):225– 244, November 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/39/2/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/39/2/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=23&issue=3&spage=225.

Langendorfer:2002:STW

[438] Peter Langendörfer, Michael Methfessel, Horst Frankenfeldt, Irina Babanskaja, Irina Matthaei, and Rolf Kraemer. Shielding TCP from wireless link errors: Retransmission effort and fragmentation. The Journal of Supercomputing, 23(3):245–260, November 2002. CODEN JOSUED.

ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/39/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/39/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=3&spage=245.

Lahanas:2002:EET

[439] A. Lahanas and V. Tsaoussidis. Experimental evaluation of TCP-probing in mobile networks. The Journal of Supercomputing, 23(3):261–279, November 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/39/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/39/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=3&spage=261.

Vardalis:2002:EFT

[440] D. Vardalis and V. Tsaoussidis. Efficiency/fairness tradeoffs in networks with wireless components and transient congestion. The Journal of Supercomputing, 23(3):281–296, November 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/39/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/39/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=23&issue=3&spage=281.

Aversa:2003:CSA

[441] R. Aversa, N. Mazzocca, and U. Villano. A case study of application analytical modeling in heterogeneous computing environments: Cholesky factorization in a NOW. The Journal of Supercomputing, 24(1):5–24, January 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/40/1/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/40/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=24&issue=1&spage=5.

Yeo:2003:APR

[442] Jaeheung Yeo, Heon Y. Yeom, and Taesoon Park. An asynchronous protocol for release consistent distributed shared memory systems. The Journal of Supercomputing, 24(1):25–41, January 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/40/2/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/40/2/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=24&issue=1&spage=25.

Knijnenburg:2003:CST

[443] P. M. W. Knijnenburg, T. Kisuki, and M. F. P. O’Boyle. Combined selection of tile sizes and unroll factors using iterative compilation. The Journal of Supercomputing, 24(1):43– 67, January 2003. CODEN JOSUED.

ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/40/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/40/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=1&spage=43.

Yun:2003:PPB

[444] Jae Heon Yun. Parallel performance of block ILU preconditioners for a block-tridiagonal matrix. The Journal of Supercomputing, 24(1):69–89, January 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/40/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/40/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=1&spage=69.

Damaj:2003:PAL

[445] Issam Damaj and Hassan Diab. Performance analysis of linear algebraic functions using reconfigurable computing. The Journal of Supercomputing, 24(1):91–107, January 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/40/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/40/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=1&spage=91.

Arbab:2003:I

[446] Farhad Arbab and Khalil Drira. Introduction. The Journal of Supercomputing, 24(2):119–120, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/41/1/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/41/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=24&issue=2&spage=119.

DeBosschere:2003:STM

[447] K. De Bosschere and J.-M. Jacquet. Suspension terms as a means for metacoordination in the μ log coordination framework. The Journal of Supercomputing, 24(2):121–129, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/41/2/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/41/2/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=24&issue=2&spage=121.

DeBruin:2003:BLC

[448] Hans De Bruin. BCOOPL: a language for controlling component interactions. The Journal of Supercomputing, 24(2):131–139, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/41/3/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/

getfile/5189/41/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=131.

Inverardi:2003:SAC

[449] Paola Inverardi and Henry Muccini.

Software architectures and coordination models. The Journal of Supercomputing, 24(2):141–149, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/41/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=141.

Pellegrini:2003:CMD

[450] Marie-Claude Pellegrini and Michel Riveill. Component management in a dynamic architecture. The Journal of Supercomputing, 24(2):151–159, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/41/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=151.

Tellioglu:2003:MCW

[451] H. Tellioğlu. Modeling coordinated work: Definition and application of the model “Coordinated work environment”. The Journal of Supercomputing, 24(2):161–171, February

2003.  CODEN JOSUED.  ISSN

0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/6/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/41/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=161.

Texier:2003:AMS

[452] Géraldine Texier and Noël Plouzeau. Automatic management of sessions in shared spaces. The Journal of Supercomputing, 24(2):173–181, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/7/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/41/7/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=173.

Dommel:2003:EGC

[453] H.-P. Dommel and J. J. Garcia-Luna-Aceves. Efficient group coordination in multicast trees. The Journal of Supercomputing, 24(2):183–192, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/8/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/41/8/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=183.

Barkaoui:2003:CCM

[454] Kamel Barkaoui and Rabah Benamara. On concurrency control in multidatabase systems with an extended transaction model. The Journal of Supercomputing, 24(2):193–202, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/9/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/41/9/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=193.

Rafiq:2003:CAD

[455] Omar Rafiq and Leo Cacciari. Coordination algorithm for distributed testing. The Journal of Supercomputing, 24(2):203–211, February 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/41/10/abstract.htm;   http:/
/ipsapp009.kluweronline.com/content/
getfile/5189/41/10/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=2&spage=203.

Lin:2003:F

[456] Yen-Chun Lin and Hong Shen. Foreword. The Journal of Supercomputing, 24(3):227–228, March 2003. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp009.kluweronline.com/content/
getfile/5189/42/1/abstract.htm;
http://ipsapp009.kluweronline.
com/content/getfile/5189/42/1/fulltext.

pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=24&issue=3&spage=
227.

Kittitornkun:2003:PAS

[457] Surin Kittitornkun and Yu Hen Hu. Processor array synthesis from shiftvariant deep nested do loops. The Journal of Supercomputing, 24(3):229– 249, March 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/42/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/42/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=3&spage=229.

Li:2003:MET

[458] Jie Li, Yi Pan, and Hong Shen. More efficient topological sort using reconfigurable optical buses. The Journal of Supercomputing, 24(3):251–258, March 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/42/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/42/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=3&spage=251.

Lin:2003:FOO

[459] Wen-Yang Lin. Finding optimal ordering of sparse matrices for columnoriented parallel Cholesky factorization. The Journal of Supercomputing, 24(3):259–277, March 2003. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp009.kluweronline.com/content/
getfile/5189/42/4/abstract.htm;
 http://ipsapp009.kluweronline.
com/content/getfile/5189/42/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=24&issue=3&spage=
259.

Lin:2003:CFD

[460] Yen-Chun Lin, Yao-Hsien Hsu, and Chun-Keng Liu. Constructing H4, a fast depth-size optimal parallel prefix circuit. The Journal of Supercomputing, 24(3):279–304, March 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/42/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/42/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=3&spage=279.

Lin:2003:FCP

[461] Feng-Tse Lin and Jing-Shing Yao.

Fuzzy critical path method based on signed-distance ranking and statistical confidence-interval estimates. The Journal of Supercomputing, 24(3):305– 325, March 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/42/6/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/42/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=24&issue=3&spage=305.

Sum:2003:AEA

[462] John Sum, Hong Shen, G. Young, Jie Wu, and Chi-Sing Leung. Analysis on extended ant routing algorithms for network routing and management. The Journal of Supercomputing, 24(3):327– 340, March 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/42/7/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/42/7/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=24&issue=3&spage=327.

Jan:2003:EAP

[463] Gene Eu Jan and Yuan-Shin Hwang. An efficient algorithm for perfect load balancing on hypercube multiprocessors. The Journal of Supercomputing, 25(1): 5–15, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/43/1/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/43/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=25&issue=1&spage=5.

Hwang:2003:SAE

[464] Gwan-Hwan Hwang, Cheng-Wei Chen, Jenq Kuen Lee, and Roy Dz-Ching Ju. Segmented alignment: An enhanced model to align data parallel programs of HPF. The Journal of Supercomputing, 25(1):17–41, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.

kluweronline.com/content/getfile/
5189/43/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/43/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=1&spage=17.

Wani:2003:PER

[465] M. Arif Wani and Hamid R. Arabnia. Parallel edge-region-based segmentation algorithm targeted at reconfigurable MultiRing network. The Journal of Supercomputing, 25(1):43– 62, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/43/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/43/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=1&spage=43.

Kwok:2003:EHC

[466] Yu-Kwong Kwok. On exploiting heterogeneity for cluster based parallel multithreading using task duplication. The Journal of Supercomputing, 25(1):63– 72, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/43/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/43/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=1&spage=63.

Cha:2003:EAT

[467] Hojung Cha, Rhan Ha, and Jane W. S. Liu. Experimental analysis of timing validation methods for distributed real-time systems. The Journal of Supercomputing, 25(1):73– 94, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/43/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/43/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=1&spage=73.

Gravvanis:2003:GEE

[468] George A. Gravvanis. Guest Editor’s editorial computational science and scientific computing. The Journal of Supercomputing, 25(2):107–108, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/1/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=107.

Huckle:2003:FSA

[469] Thomas Huckle. Factorized sparse approximate inverses for preconditioning. The Journal of Supercomputing, 25(2): 109–117, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/3/fulltext.pdf;
http://www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=109.

Gravvanis:2003:SBV

[470] George A. Gravvanis. On the solution of boundary value problems by using fast generalized approximate inverse banded matrix techniques. The Journal of Supercomputing, 25(2):119– 129, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=119.

Shires:2003:OPF

[471] Dale Shires and Ram Mohan. Optimization and performance of a Fortran 90 MPI-based unstructured code on large-scale parallel systems. The Journal of Supercomputing, 25(2):131– 141, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=131.

Gerndt:2003:LSC

[472] Andreas Gerndt, Thomas Van Reimersdahl, Torsten Kuhlen, Christian Bischof, Ingolf Hörschler, Matthias Meinke, and Wolfgang Schröder. Large-scale CFD data handling in a VR-based otorhinolaryngological CAS-system using a Linux-cluster. The Journal of Supercomputing, 25(2):143–154, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=143.

Santos:2003:PCM

[473] Eunice E. Santos. Parallel complexity of matrix multiplication. The Journal of Supercomputing, 25(2):155– 175, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/6/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=155.

Gavrilova:2003:CED

[474] Marina L. Gavrilova and Muhammad H. Alsuwaiyel. Computing the Euclidean distance transform on a linear array of processors. The Journal of Supercomputing, 25(2):177–185, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/44/7/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/44/7/fulltext.pdf;
http://www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=25&issue=2&spage=177.

Guo:2003:SCS

[475] Minyi Guo, Yi Pan, and Zhen Liu.

Symbolic communication set generation for irregular parallel applications. The Journal of Supercomputing, 25(3):199– 214, July 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/45/1/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/45/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=3&spage=199.

Zhu:2003:CQS

[476] W. Zhu. Cluster queue structure for shared-memory multiprocessor systems. The Journal of Supercomputing, 25(3): 215–236, July 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/45/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/45/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=25&issue=3&spage=215.

Onbascioglu:2003:ODD

[477] Esin Onbasçioğlu and Linet Özdamar.

Optimization of data distribution and processor allocation problem using simulated annealing. The Journal of Supercomputing, 25(3):237–253, July 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.

                                        kluweronline.com/content/getfile/
                                        5189/45/3/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/45/3/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=25&issue=3&spage=237.

Jiang:2003:FTB

[478] Zhen Jiang and Jie Wu. Faulttolerant broadcasting in 2-D wormholerouted meshes. The Journal of Supercomputing, 25(3):255–275, July 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/45/4/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/45/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=25&issue=3&spage=255.

Anonymous:2003:GEI

[479] Anonymous. Guest Editor introduction for the special section on commercial applications for highperformance computing. The Journal of Supercomputing, 26(1):5–6, August 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/46/1/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/46/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=1&spage=5.

Braun:2003:ICD

[480] Terry A. Braun, Todd E. Scheetz, Gregg Webster, Abe Clark, Edwin M. Stone, Val C. Sheffield, and Thomas L. Casavant. Identifying candidate disease genes with high-performance computing. The Journal of Supercomputing, 26(1):7– 24, August 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/46/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/46/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=1&spage=7.

Thoennes:2003:EPD

[481] Mathew S. Thoennes and Charles C.

Weems. Exploration of the performance of a data mining application via hardware based monitoring. The Journal of Supercomputing, 26(1):25– 42, August 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/46/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/46/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=1&spage=25.

Thulasiram:2003:PEM

[482] Ruppa K. Thulasiram and Parimala Thulasiraman. Performance evaluation of a multithreaded Fast Fourier Transform algorithm for derivative pricing. The Journal of Supercomputing, 26(1): 43–58, August 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/

                                        5189/46/4/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/46/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=1&spage=43.

Jin:2003:FLL

[483] Hyun-Wook Jin, Chuck Yoo, and Jin-Young Choi. Firmware-level latency analysis on a Gigabit network. The Journal of Supercomputing, 26(1):59– 75, August 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/46/5/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/46/5/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=1&spage=59.

Min:2003:CDW

[484] G. Min and M. Ould-Khaoua. Communication delay in wormhole-switched tori networks under bursty workloads. The Journal of Supercomputing, 26(1): 77–94, August 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/46/6/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/46/6/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=1&spage=77.

Plaks:2003:ECSa

[485] Toomas P. Plaks and Peter M. Athanas. Engineering of configurable systems: Guest Editors’ foreword. The Journal of Supercomputing, 26(2):107–108, September 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/47/1/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/47/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=2&spage=107.

Plessl:2003:ISA

[486] Christian Plessl and Marco Platzner.

Instance-specific accelerators for minimum covering. The Journal of Supercomputing, 26(2):109–129, September 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/47/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/47/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=2&spage=109.

Gokhale:2003:EHP

[487] Maya Gokhale, Jan Frigo, Kevin Mccabe, James Theiler, Christophe Wolinski, and Dominique Lavenier. Experience with a hybrid processor: Kmeans clustering. The Journal of Supercomputing, 26(2):131–148, September 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/47/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/47/3/fulltext.pdf;
http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=2&spage=131.

Bednara:2003:ASF

[488] Marcus Bednara and Jürgen Teich. Automatic synthesis of FPGA processor arrays from loop algorithms. The Journal of Supercomputing, 26(2):149–165, September 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/47/4/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/47/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=2&spage=149.

Baumgarte:2003:PXS

[489] V. Baumgarte, G. Ehlers, F. May, A. Nückel, M. Vorbach, and M. Weinhardt. PACT XPP — a selfreconfigurable data processing architecture. The Journal of Supercomputing, 26(2):167–184, September 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/47/5/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/47/5/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=2&spage=167.

Kretzschmar:2003:LPE

[490] Claudia Kretzschmar, Robert Siegmund, and Dietmar Müller. Low power encoding techniques for dynamically reconfigurable hardware. The Journal of Supercomputing, 26(2):185–203, September 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/47/6/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/47/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=2&spage=185.

Plaks:2003:ECSb

[491] Toomas P. Plaks and Peter M. Athanas.

Engineering of configurable systems, II. Guest Editors’ foreword. The Journal of Supercomputing, 26(3):219–220, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/48/1/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/48/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=3&spage=219.

Janson:2003:ECA

[492] Stefan Janson, Daniel Merkle, Martin Middendorf, Hossam Elgindy, and Hartmut Schmeck. On enforced convergence of ACO and its implementation on the reconfigurable mesh architecture using size reduction tasks. The Journal of Supercomputing, 26(3):221–238, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/48/2/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/48/2/fulltext.pdf;
http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=3&spage=221.

Puttegowda:2003:CSR

[493] Kiran Puttegowda, David I. Lehn, Jae H. Park, Peter Athanas, and Mark Jones. Context switching in a runtime reconfigurable system. The Journal of Supercomputing, 26(3):239–257, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/48/3/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/48/3/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=3&spage=239.

Choi:2003:DSM

[494] Seonil Choi, Ju wook Jang, Sumit Mohanty, and Viktor K. Prasanna. Domain-specific modeling for rapid energy estimation of reconfigurable architectures. The Journal of Supercomputing, 26(3):259–281, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/48/4/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/48/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=26&issue=3&spage=259.

Heysters:2003:FEE

[495] Paul Heysters, Gerard Smit, and Egbert Molenkamp. A flexible and energyefficient coarse-grained reconfigurable architecture for mobile systems. The Journal of Supercomputing, 26(3):283– 308, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/48/5/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/48/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=3&spage=283.

Siemers:2003:UCB

[496] Christian Siemers and Volker Winterstein. The universal configurable block/ machine — an approach for a configurable SoC-architecture. The Journal of Supercomputing, 26(3):309–331, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/48/6/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/48/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=3&spage=309.

Anonymous:2003:E

[497] Anonymous. Erratum. The Journal of Supercomputing, 26(3):332, November 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/48/7/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/48/7/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=26&issue=3&spage=332.

Izadi:2004:AAT

[498] Baback A. Izadi and Füsun Özgüner. An augmented k-ary tree multiprocessor with real-time fault-tolerant capability. The Journal of Supercomputing, 27(1): 5–17, January 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/49/1/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/49/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=27&issue=1&spage=5.

Er-El:2004:CMF

[499] David Er-El and Dror G. Feitelson. Communication models for a free-space optical cross-connect switch. The Journal of Supercomputing, 27(1):19–48, January 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        5189/49/2/abstract.htm;   http://
                                        ipsapp009.kluweronline.com/content/
                                        getfile/5189/49/2/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=27&issue=1&spage=19.

Peigin:2004:PLS

[500] S. Peigin, B. Epstein, T. Rubin, and S. Seror. Parallel large scale high accuracy Navier–Stokes computations on distributed memory clusters. The Journal of Supercomputing, 27(1):49– 68, January 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/

5189/49/3/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/49/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=1&spage=49.

Bhalla:2004:ABI

[501] Subhash Bhalla and Stuart E. Madnick. Asynchronous backup and initialization of a database server for replicated database systems. The Journal of Supercomputing, 27(1):69–89, January 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
5189/49/4/abstract.htm;   http://
ipsapp009.kluweronline.com/content/
getfile/5189/49/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=1&spage=69.

Katsinis:2004:SIN

[502] Constantine Katsinis and Bahram Nabet. A scalable interconnection network architecture for petaflops computing. The Journal of Supercomputing, 27(2):103–128, February 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/50/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/50/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=2&spage=103.

Zhang:2004:EHG

[503] Yi Zhang, Steve Haga, and Rajeev Barua. Execution history guided instruction prefetching. The Journal of Supercomputing, 27(2):129–147, February 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/50/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/50/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=2&spage=129.

Cheung:2004:LBA

[504] Lap sun Cheung and Yu kwok Kwok. On load balancing approaches for distributed object computing systems. The Journal of Supercomputing, 27(2):149– 175, February 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/50/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/50/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=2&spage=149.

Sinnen:2004:TSA

[505] Oliver Sinnen and Leonel Sousa. On task scheduling accuracy: Evaluation methodology and results. The Journal of Supercomputing, 27(2):177–194, February 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/50/4/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/50/4/fulltext.pdf;
http://www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=27&issue=2&spage=177.

Vigo-aguiar:2004:PBV

[506] J. Vigo-aguiar and S. Natesan. A parallel boundary value technique for singularly perturbed two-point boundary value problems. The Journal of Supercomputing, 27(2):195–206, February 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/50/5/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/50/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=2&spage=195.

Yi:2004:TCL

[507] Qing Yi, Ken Kennedy, and Vikram Adve. Transforming complex loop nests for locality. The Journal of Supercomputing, 27(3):219–264, March 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/51/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/51/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=3&spage=219.

Loucif:2004:MLD

[508] S. Loucif and M. Ould-khaoua. Modeling latency in deterministic wormholerouted hypercubes under hot-spot traffic. The Journal of Supercomputing, 27(3):265–278, March 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.kluweronline.com/content/
getfile/5189/51/2/abstract.htm;
 http://ipsapp008.kluweronline.
com/content/getfile/5189/51/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
0920-8542&volume=27&issue=3&spage=
265.

Hsu:2004:QAT

[509] Chung hsing Hsu and Ulrich Kremer. A quantitative analysis of tile size selection algorithms. The Journal of Supercomputing, 27(3):279–294, March 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/51/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/51/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=3&spage=279.

Kim:2004:PPF

[510] Hong soog Kim, Young ha Yoon, and Dong soo Han. Parallel processing of first order linear recurrence on SMP machines. The Journal of Supercomputing, 27(3):295–310, March 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
5189/51/4/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/5189/51/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=27&issue=3&spage=295.

Gonzalez:2004:ERU

[511] Teofilo F. Gonzalez. Efficient resource utilization in parallel and distributed systems. The Journal of Supercomputing, 28(1):5–6, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/52/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=1&spage=5.

Suh:2004:DPS

[512] G. E. Suh, L. Rudolph, and S. Devadas. Dynamic partitioning of shared cache memory. The Journal of Supercomputing, 28(1):7–26, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/52/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=1&spage=7.

Lee:2004:ESA

[513] Singling Lee, Hann jang Ho, and Wen wei Mai. An efficient scheduling algorithm for information delivery on VoD system. The Journal of Supercomputing, 28(1):27–41, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/52/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=1&spage=27.

Dekel:2004:IIT

[514] Eliezer Dekel and Gera Goft. ITRA: Inter-tier relationship architecture for end-to-end QoS. The Journal of Supercomputing, 28(1):43–70, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/52/A/4/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=1&spage=43.

Li:2004:ECC

[515] Yamin Li, Shietung Peng, and Wanming Chu. Efficient collective communications in Dual-Cube. The Journal of Supercomputing, 28(1):71–90, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/52/A/5/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=1&spage=71.

Chan:2004:RTS

[516] Bryan Chan and Tarek S. Abdelrahman. Run-time support for the automatic parallelization of Java programs. The Journal of Supercomputing, 28(1): 91–117, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/52/A/6/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=1&spage=91.

Anonymous:2004:CAa

[517] Anonymous. Contributing authors. The Journal of Supercomputing, 28(1):119– 123, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/52/A/7/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=1&spage=119.

Anonymous:2004:IA

[518] Anonymous. Information for authors.

The Journal of Supercomputing, 28(1): 125–127, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/52/A/8/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=1&spage=125.

Gavrilova:2004:GEE

[519] Marina L. Gavrilova and C. J. Kenneth Tan. Guest Editors’ editorial: Special issue: Iterative solving environments and optimization techniques for scientific applications. The Journal of Supercomputing, 28(2):131–133, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=2&spage=131.

Nool:2004:PIL

[520] Margreet Nool and Michael M. J.

Proot. Parallel implementation of a least-squares spectral element solver for incompressible flow problems. The Journal of Supercomputing, 28(2):135– 148, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=2&spage=135.

Cheng:2004:PPT

[521] Jing-Ru C. Cheng and Paul E. Plassmann. A parallel particle tracking framework for applications in scientific computing. The Journal of Supercomputing, 28(2):149–164, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=2&spage=149.

Gavrilova:2004:ESO

[522] Marina L. Gavrilova. Empirical studies of optimization techniques in the event-driven simulation of mechanically alloyed materials. The Journal of Supercomputing, 28(2):165–176, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/4/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=2&spage=165.

Blanco:2004:PPP

[523] V. Blanco, P. González, J. C. Cabaleiro, D. B. Heras, T. F. Pena, J. J. Pombo, and F. F. Rivera. Performance prediction for parallel iterative solvers. The Journal of Supercomputing, 28(2):177– 191, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/5/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=2&spage=177.

Rajabi:2004:ODI

[524] Mohammad A. Rajabi and J. A. Rod Blais. Optimization of DTM interpolation using SFS with single satellite imagery. The Journal of Supercomputing, 28(2):193–213, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/6/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=2&spage=193.

Ralphs:2004:LHI

[525] T. K. Ralphs, L. Ládanyi, and M. J. Saltzman. A library hierarchy for implementing scalable parallel search algorithms. The Journal of Supercomputing, 28(2):215–234, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/53/A/7/abstract.htm;
http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=2&spage=215.

Dixon:2004:UDC

[526] Matthew F. Dixon and C. J. Kenneth Tan. Using distributed computers to deterministically approximate higher dimensional convection diffusion equations. The Journal of Supercomputing, 28(2):235–253, May 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/53/A/8/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=2&spage=235.

Yang:2004:SIH

[527] Laurence Tianruo Yang. Special issue on high performance scientific and engineering applications. The Journal of Supercomputing, 28(3):263–264, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/54/A/1/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=3&spage=263.

Li:2004:HPT

[528] Xiaolin Li and Manish Parashar. Hierarchical partitioning techniques for structured adaptive mesh refinement applications. The Journal of Supercomputing, 28(3):265–278, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/

x/J/5189/I/54/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=3&spage=265.

Bourgeade:2004:DLB

[529] A. Bourgeade and B. Nkonga. Dynamic load balancing computation of pulses propagating in a nonlinear medium. The Journal of Supercomputing, 28(3):279– 294, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/54/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=3&spage=279.

Zhuang:2004:GBL

[530] Yi chang Zhuang, Tyng Yue Liang, Ce kuen Shieh, Jun qi Lee, and Laurence Tianruo Yang. A group-based load balance scheme for software distributed shared memory systems. The Journal of Supercomputing, 28(3):295– 309, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/54/A/4/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=3&spage=295.

Martin:2004:HPA

[531] María J. Martín, Marta Parada, and Ramón Doallo. High performance air pollution simulation using OpenMP. The Journal of Supercomputing, 28(3): 311–321, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.

                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/54/A/5/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=3&spage=311.

Chaudhary:2004:EPT

[532] V. Chaudhary, W. L. Hase, H. Jiang, L. Sun, and D. Thaker. Experiments with parallelizing tribology simulations. The Journal of Supercomputing, 28(3): 323–343, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/54/A/6/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=3&spage=323.

Bhalla:2004:PCC

[533] Subhash Bhalla. Parallel concurrency control activity for transaction management in real-time database systems. The Journal of Supercomputing, 28(3):345– 369, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/54/A/7/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=28&issue=3&spage=345.

Anonymous:2004:TCV

[534] Anonymous. Table of contents: Volume 28 (2004). The Journal of Supercomputing, 28(3):371–372, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/54/A/8/abstract.htm;

http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=28&issue=3&spage=371.

Niculescu:2004:DDC

[535] Virginia Niculescu. On data distributions in the construction of parallel programs. The Journal of Supercomputing, 29(1):5–25, July 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/55/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=1&spage=5.

Grunberg:2004:SRT

[536] Marc Grunberg, Stéphane Genaud, and Catherine Mongenet. Seismic raytracing and earth mesh modeling on various parallel architectures. The Journal of Supercomputing, 29(1):27– 44, July 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/55/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=1&spage=27.

Albdaiwi:2004:PDP

[537] Bader F. Albdaiwi and Marilynn L.

Livingston. Perfect distance-d placements in 2D toroidal networks. The Journal of Supercomputing, 29(1):45– 57, July 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/55/A/3/abstract.htm;

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=1&spage=45.

Wu:2004:EMM

[538] Jan jan Wu, Shih hsien Yeh, and Pangfeng Liu. Efficient multiple multicast on heterogeneous Network of Workstations. The Journal of Supercomputing, 29(1):59–88, July 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/55/A/4/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=1&spage=59.

Akl:2004:SPR

[539] Selim G. Akl. Superlinear performance in real-time parallel computation. The Journal of Supercomputing, 29(1):89– 111, July 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/55/A/5/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=1&spage=89.

Guo:2004:EPD

[540] Minyi Guo. Editorial: Parallel and distributed processing with applications. The Journal of Supercomputing, 29(2): 123, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/56/A/1/abstract.htm;
                                        http://www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=29&issue=2&spage=123.

Hsu:2004:CDR

[541] Ching-Hsien Hsu and Kun-Ming Yu. A compressed diagonals remapping technique for dynamic data redistribution on banded sparse matrix. The Journal of Supercomputing, 29(2):125– 143, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/56/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=2&spage=125.

Yang:2004:PMI

[542] Laurence Tianruo Yang and Richard P.

Brent. Parallel MCGLS and ICGLS methods for least squares problems on distributed memory architectures. The Journal of Supercomputing, 29(2):145– 156, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/56/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=2&spage=145.

Wang:2004:DCA

[543] Hui Wang, Minyi Guo, and Daming Wei. A divide-and-conquer algorithm for irregular redistribution in parallelizing compilers. The Journal of Supercomputing, 29(2):157–170, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/

                                        x/J/5189/I/56/A/4/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=2&spage=157.

Li:2004:DEA

[544] Xuhui Li, Jiannong Cao, and Yanxiang He. A direct execution approach to simulating mobile agent algorithms. The Journal of Supercomputing, 29(2):171– 184, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/56/A/5/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=2&spage=171.

Chen:2004:FEP

[545] Ling Chen, Hongjian Chen, Yi Pan, and Yixin Chen. A fast efficient parallel Hough transform algorithm on LARPBS. The Journal of Supercomputing, 29(2):185–195, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/56/A/6/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=2&spage=185.

Jia:2004:DCP

[546] Weijia Jia, Bo Han, Chuanlin Zhang, and Wanlei Zhou. Delay control and parallel admission algorithms for realtime anycast flow. The Journal of Supercomputing, 29(2):197–209, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.

kluweronline.com/IPS/content/ext/
x/J/5189/I/56/A/7/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=2&spage=197.

Lee:2004:DPB

[547] Chain-Wu Lee, Chun-Hsi Huang, Laurence Tianruo Yang, and Sanguthevar Rajasekaran. Distributed path-based inference in semantic networks. The Journal of Supercomputing, 29(2):211– 227, August 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/56/A/8/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=2&spage=211.

Peigin:2004:EPA

[548] S. Peigin and B. Epstein. Embedded parallelization approach for optimization in aerodynamic design. The Journal of Supercomputing, 29(3):243–263, September 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/58/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=29&issue=3&spage=243.

Wang:2004:MBM

[549] Neng-Chung Wang, Chih-Ping Chu, and Tzung-Shi Chen. Multipath-based multicasting strategies for wormholerouted star graph interconnection networks. The Journal of Supercomputing, 29(3):265–286, September

2004.  CODEN JOSUED.  ISSN

0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/58/A/2/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=3&spage=265.

Nasiri:2004:NAT

[550] Ramin Nasiri, Ali Moeini, and Ahmad Abdollahzadeh. A new approach towards procurement of software models via distributed business models. The Journal of Supercomputing, 29(3):287–302, September 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/58/A/3/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=3&spage=287.

Datta:2004:MAP

[551] Amitava Datta. Multiple addition and prefix sum on a linear array with a reconfigurable pipelined bus system. The Journal of Supercomputing, 29(3):303–317, September 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/58/A/4/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=29&issue=3&spage=303.

Zeyao:2004:PFS

[552] Mo Zeyao and Fu Lianxiang. Parallel flux sweep algorithm for neutron transport on unstructured grid. The Journal of Supercomputing, 30(1):5– 17, October 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/57/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=1&spage=5.

Chuang:2004:EVP

[553] Po-Jen Chuang, Young-Tzong Hsiao, and Yu-Shian Chiu. An efficient value predictor dynamically using loop and locality properties. The Journal of Supercomputing, 30(1):19–36, October 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/57/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=1&spage=19.

Lin:2004:HCF

[554] Chao Lin. Heuristic contention-free broadcast in heterogeneous networks of workstations with multiple send and receive speeds. The Journal of Supercomputing, 30(1):37–64, October 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/57/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=1&spage=37.

Anonymous:2004:CAb

[555] Anonymous. Contributing authors.

The Journal of Supercomputing, 30(1):

65, October 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/57/A/4/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=30&issue=1&spage=65.

Gravvanis:2004:GES

[556] George A. Gravvanis and Eunice E. Santos. Guest editorial: Special issue: Special issue on parallel and distributed processing and applications. The Journal of Supercomputing, 30(2):75–76, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/59/A/1/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=30&issue=2&spage=75.

Gravvanis:2004:PSS

[557] G. A. Gravvanis, K. M. Giannoutakis, M. P. Bekakos, and O. B. Efremides. Parallel and systolic solution of normalized explicit approximate inverse preconditioning. The Journal of Supercomputing, 30(2):77–96, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/59/A/2/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=30&issue=2&spage=77.

Santos:2004:OEP

[558] Eunice E. Santos. Optimal and efficient parallel tridiagonal solvers using direct methods. The Journal of Supercomputing, 30(2):97–115, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/59/A/7/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=2&spage=97.

Xu:2004:PHW

[559] Shuting Xu and Jun Zhang. A parallel hybrid Web document clustering algorithm and its performance study. The Journal of Supercomputing, 30(2):117– 131, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/59/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=2&spage=117.

Baker:2004:SSC

[560] James M. Baker, Jr., Brian Gold, Mark Bucciero, Sidney Bennett, Rajneesh Mahajan, Priyadarshini Ramachandran, and Jignesh Shah. SCMP: a single-chip message-passing parallel computer. The Journal of Supercomputing, 30(2):133– 149, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/59/A/4/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=2&spage=133.

Tabirca:2004:FGD

[561] Tatiana Tabirca, Len Freeman, Sabin Tabirca, and Laurence Tianruo Yang. Feedback guided dynamic loop scheduling: Convergence of the continuous case. The Journal of Supercomputing, 30(2):151–178, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/59/A/5/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=2&spage=151.

Maglogiannis:2004:CSF

[562] Ilias G. Maglogiannis, Elias P. Zafiropoulos, Agapios N. Platis, and George A. Gravvanis. Computing the success factors in consistent acquisition and recognition of objects in color digital images by explicit preconditioning. The Journal of Supercomputing, 30(2):179–198, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/59/A/6/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=2&spage=179.

Anonymous:2004:CAc

[563] Anonymous. Contributing authors. The Journal of Supercomputing, 30(2):199– 200, November 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/59/A/8/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=2&spage=199.

Plaks:2004:FEC

[564] Toomas P. Plaks. Foreword: Engineering of configurable systems, III. The Journal of Supercomputing, 30(3):211– 212, December 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/60/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=3&spage=211.

Ghiasi:2004:CRO

[565] Soheil Ghiasi, Hyun J. Moon, Ani Nahapetian, and Majid Sarrafzadeh. Collaborative and reconfigurable object tracking. The Journal of Supercomputing, 30(3):213–238, December 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/60/A/4/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=3&spage=213.

Smith:2004:TRB

[566] William D. Smith and Austars R.

Schnore. Towards an RCC-based accelerator for computational fluid dynamics applications. The Journal of Supercomputing, 30(3):239–261, December 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/60/A/5/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=30&issue=3&spage=239.

Rauwerda:2004:MWC

[567] Gerard K. Rauwerda, Paul M. Heysters, and Gerard J. M. Smit. Mapping wireless communication algorithms onto a reconfigurable architecture. The Journal of Supercomputing, 30(3):263–282, December 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/60/A/6/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=30&issue=3&spage=263.

Kaouane:2004:MIR

[568] Linda Kaouane, Mohamed Akil, Thierry Grandpierre, and Yves Sorel. A methodology to implement real-time applications onto reconfigurable circuits. The Journal of Supercomputing, 30(3):283– 301, December 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/60/A/2/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=30&issue=3&spage=283.

Anonymous:2004:CAd

[569] Anonymous. Contributing authors. The Journal of Supercomputing, 30(3):303– 305, December 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/60/A/3/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=30&issue=3&spage=303.

Zhao:2005:SUL

[570] Yuan Zhao and Ken Kennedy. Scalarization using loop alignment and loop skewing. The Journal of Supercomputing, 31(1):5–46, January 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/61/A/1/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=1&spage=5.

Haga:2005:DFU

[571] Steve Haga, Natasha Reeves, Rajeev Barua, and Diana Marculescu. Dynamic functional unit assignment for low power. The Journal of Supercomputing, 31(1):47–62, January 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/61/A/2/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=1&spage=47.

Wu:2005:DFM

[572] Jie Wu and Li Sheng. Deadlockfree multicasting in irregular networks using prefix routing. The Journal of Supercomputing, 31(1):63–78, January 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/5189/I/61/A/3/abstract.htm;
http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=1&spage=63.

Kamangar:2005:MAC

[573] Farhad Kamangar, David Levine, Gergely V. Záruba, and Renjith Thomas. Mobile agent Connection Establishment and Management (CEMA)

— message exchange for pervasive computing environments. The Journal of Supercomputing, 31(1):79–99, January 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/61/A/4/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=31&issue=1&spage=79.

Anonymous:2005:CAa

[574] Anonymous. Contributing authors. The Journal of Supercomputing, 31(1):101– 103, January 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/5189/I/61/A/5/abstract.htm;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=31&issue=1&spage=101.

Chang:2005:PTD

[575] Weng-Long Chang, Chih-Ping Chu, and Jia-Hwa Wu. A polynomial-time dependence test for determining integervalued solutions in multi-dimensional arrays under variable bounds. The Journal of Supercomputing, 31(2):111– 135, February 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=31&issue=2&spage=111.

Liao:2005:PEP

[576] Wei-Keng Liao, Alok Choudhary, Donald Weiner, and Pramod Varshney. Performance evaluation of a parallel pipeline computational model for space-time adaptive processing. The Journal of Supercomputing, 31(2):137– 160, February 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=2&spage=137.

Basharahil:2005:DSA

[577] Ramzi Basharahil, Brian Wims, Cheng-Zhong Xu, and Song Fu. Distributed shared arrays: An integration of message passing and multithreading on SMP clusters. The Journal of Supercomputing, 31(2):161–184, February 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=2&spage=161.

Zarandi:2005:HBS

[578] Hamid Reza Zarandi and Hamid Sarbazi-Azad. Hierarchical binary set partitioning in cache memories. The Journal of Supercomputing, 31 (2):185–202, February 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=2&spage=185.

Karimou:2005:AIP

[579] Djibo Karimou and Jean Frédéric Myoupo. An application of an initialization protocol to permutation routing in a single-hop mobile ad hoc networks. The Journal of Supercomputing, 31(3):215–226, March 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=3&spage=215.

Sklavos:2005:ISH

[580] N. Sklavos and O. Koufopavlou. Implementation of the SHA-2 hash family standard using FPGAs. The Journal of Supercomputing, 31(3):227– 248, March 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=3&spage=227.

Sinop:2005:PPH

[581] Ali Kemal Sinop, Tolga Abaci, Ümit Akkus, Attila Gürsoy, and Ugur Güdükbay. PHR: a parallel hierarchical radiosity system with dynamic load balancing. The Journal of Supercomputing, 31(3):249–263, March 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=3&spage=249.

Salleh:2005:SRT

[582] Shaharuddin Salleh, Stephan Olariu, Bahrom Sanugi, and Mohd Ismail Abd Aziz. Single-row transformation of complete graphs. The Journal of Supercomputing, 31(3):265–279, March 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=3&spage=265.

Anonymous:2005:CAb

[583] Anonymous. Contributing authors.

The Journal of Supercomputing, 31 (3):281–282, March 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=31&issue=3&spage=281.

Taha:2005:PSS

[584] Thiab R. Taha and Xiangming Xu.

Parallel split-step Fourier methods for the coupled nonlinear Schrödinger type equations. The Journal of Supercomputing, 32(1):5–23, April 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=1&spage=5.

Lee:2005:QAA

[585] Wonjun Lee, Jaideep Srivastava, and Bikash Sabata. QoS-aware admission control and dynamic resource provisioning framework in ubiquitous multimedia computing environments. The Journal of Supercomputing, 32 (1):25–50, April 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=1&spage=25.

Michailidis:2005:NPA

[586] Panagiotis D. Michailidis and Konstantinos G. Margaritis. New processor array architectures for the longest common subsequence problem. The Journal of Supercomputing, 32(1):51–69, April 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=1&spage=51.

Parsa:2005:DIF

[587] S. Parsa and O. Bushehrian. The design and implementation of a framework for automatic modularization of software systems. The Journal of Supercomputing, 32(1):71–94, April 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=1&spage=71.

Anonymous:2005:CAc

[588] Anonymous. Contributing authors. The Journal of Supercomputing, 32 (1):95, April 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=1&spage=95.

Plaks:2005:ECS

[589] Toomas P. Plaks. Engineering of configurable systems, IV: Foreword. The Journal of Supercomputing, 32 (2):103–104, May 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=2&spage=103.

Bellows:2005:HVD

[590] Peter Bellows. High-visibility debugby-design for FPGA platforms. The Journal of Supercomputing, 32(2): 105–118, May 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=2&spage=105.

Todman:2005:CHC

[591] Tim Todman, José Gabriel de F. Coutinho, and Wayne Luk. Customisable hardware compilation. The Journal of Supercomputing, 32(2):119–137, May 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=2&spage=119.

Prasanna:2005:EEC

[592] Viktor K. Prasanna. Energy-efficient computations on FPGAs. The Journal of Supercomputing, 32(2):139– 162, May 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=2&spage=139.

Gatzka:2005:ACR

[593] Stephan Gatzka and Christian Hochberger.

The AMIDAR class of reconfigurable processors. The Journal of Supercomputing, 32(2):163–181, May 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=2&spage=163.

Anonymous:2005:CAd

[594] Anonymous. Contributing authors. The Journal of Supercomputing, 32 (2):183–184, May 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=32&issue=2&spage=183.

Zhu:2005:HAO

[595] Ming Zhu, Wentong Cai, Bu-Sung Lee, and Xudong Wu. A hybrid analysis of an optimization approach for cluster applications. The Journal of Supercomputing, 32(3):191–215, June 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=32&issue=3&spage=191.

Sowa:2005:PQP

[596] Masahiro Sowa, Ben A. Abderazek, and Tsutomu Yoshinaga. Parallel queue processor architecture based on produced order computation model. The Journal of Supercomputing, 32 (3):217–229, June 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=32&issue=3&spage=217.

Al-Ayyoub:2005:DUB

[597] Abdel-Elah Al-Ayyoub. Distributed unidirectional and bidirectional heuristic search: Algorithm design and empirical assessment. The Journal of Supercomputing, 32(3):231–250, June 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=3&spage=231.

Alonso:2005:EPA

[598] P. Alonso, J. M. Badía, and A. M.

Vidal. An efficient parallel algorithm to solve block-Toeplitz systems. The Journal of Supercomputing, 32 (3):251–278, June 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=3&spage=251.

Anonymous:2005:CAe

[599] Anonymous. Contributing authors.

The Journal of Supercomputing, 32 (3):279–280, June 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=32&issue=3&spage=279.

Mun:2005:GE

[600] Youngsong Mun. Guest editorial.

The Journal of Supercomputing, 33 (1–2):5–6, July 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=5.

Bang:2005:BAM

[601] Young-Cheol Bang and Hyunseung Choo. On bandwidth adjusted multicast communications in pipeline router architecture. The Journal of Supercomputing, 33(1–2):7–18, July 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=7.

Yang:2005:EPV

[602] Hoonmo Yang and Moonkey Lee. Embedded processor validation environment using a cycle-accurate retargetable instruction-set simulator. The Journal of Supercomputing, 33 (1–2):19–32, July 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=19.

Mun:2005:PAB

[603] Youngsong Mun. Performance analysis of banyan-type multistage interconnection networks under nonuniform traffic pattern. The Journal of Supercomputing, 33(1–2):33–52, July 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=33.

Lim:2005:EDI

[604] Minyeol Lim and Eui-Nam Huh. An efficient design and implementation for grid advanced information service. The Journal of Supercomputing, 33(1–2):53–63, July 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=53.

Park:2005:NAH

[605] Gyung-Leen Park and Hee Youn Yong.

A new approach for high performance computing systems with various checkpointing schemes. The Journal of Supercomputing, 33(1–2): 65–78, July 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=65.

Ahn:2005:SUS

[606] Seongjin Ahn and Jin Chung Wook.

A study upon SNMP MIB II algorithm: For the calculation of the line utilization. The Journal of Supercomputing, 33(1–2):79–91, July 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=79.

Joon:2005:ITS

[607] Byoung Min Joon. Intrusion-tolerant server architecture for survivable services. The Journal of Supercomputing, 33(1–2):93–102, July 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=93.

Kim:2005:PPR

[608] D. S. Kim, Y. J. Jung, and T. M.

Chung. PRISM: a Preventive and Risk-Reducing Integrated Security Management model using security label. The Journal of Supercomputing, 33 (1–2):103–121, July 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=103.

Kim:2005:SMA

[609] Byung-Gi Kim and Sang-Sun Hong. Secure mutual authentication for ad hoc wireless networks. The Journal of Supercomputing, 33(1–2):123– 132, July 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=123.

Lee:2005:NIW

[610] Joong-Jae Lee, Won Kim, Na-Young Lee, and Gye-Young Kim. A new incremental watermarking based on dual-tree complex wavelet transform. The Journal of Supercomputing, 33 (1–2):133–140, July 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=133.

Anonymous:2005:CAf

[611] Anonymous. Contributing authors. The Journal of Supercomputing, 33 (1–2):141–147, July 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=1&spage=141.

Deligiannidis:2005:IPV

[612] Leonidas Deligiannidis and Robert J. K.

Jacob. Improving performance of virtual reality applications through parallel processing. The Journal of Supercomputing, 33(3):155–173, September 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=3&spage=155.

Hsu:2005:SMB

[613] Ching-Hsien Hsu. Sparse matrix block-cyclic realignment on distributed memory machines. The Journal of Supercomputing, 33(3):175–196, September 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=3&spage=175.

Athanasaki:2005:HGP

[614] Maria Athanasaki, Aristidis Sotiropoulos, Georgios Tsoukalas, Nectarios Koziris, and Panayiotis Tsanakas. Hyperplane grouping and pipelined schedules: How to execute tiled loops fast on clusters of SMPs. The Journal of Supercomputing, 33(3):197– 226, September 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=3&spage=197.

Anonymous:2005:CAg

[615] Anonymous. Contributing authors.

The Journal of Supercomputing, 33 (3):227, September 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=33&issue=3&spage=227.

Wang:2005:ETB

[616] Nen-Chung Wang and Chih-Ping Chu. An efficient tree-based multicasting algorithm on wormhole-routed star graph interconnection networks embedded with Hamiltonian path. The Journal of Supercomputing, 34(1):5– 26, October 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=1&spage=5.

Joshua:2005:MBS

[617] Rose Joshua and David H. Scuse. Modeling beliefs and solution strategies in a distributed learning system. The Journal of Supercomputing, 34 (1):27–39, October 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=1&spage=27.

Garz:2005:ABP

[618] E. M. Garz and I. García. Approaches based on permutations for partitioning sparse matrices on multiprocessors. The Journal of Supercomputing, 34(1):41–61, October 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=34&issue=1&spage=41.

Chen:2005:DDP

[619] Ching-Wen Chen and Chung-Ping Chung. Designing a disjoint paths interconnection network with fault tolerance and collision solving. The Journal of Supercomputing, 34(1):63– 80, October 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=1&spage=63.

Anonymous:2005:CAh

[620] Anonymous. Contributing authors.

The Journal of Supercomputing, 34 (1):81, October 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=1&spage=81.

Ould-Khaoua:2005:PEG

[621] Mohamed Ould-Khaoua and Geyong Min. Performance evaluation of Grid and cluster computing systems. The Journal of Supercomputing, 34(2):91– 92, November 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=91.

Bacigalupo:2005:IAD

[622] David A. Bacigalupo, Stephen A. Jarvis, Ligang He, Daniel P. Spooner, Donna N. Dillenberger, and Graham R. Nudd. An investigation into the application of different performance prediction methods to distributed enterprise applications. The Journal of Supercomputing, 34(2):93–111, November 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=93.

Cicotti:2005:DPM

[623] P. Cicotti, M. Taufer, and Andrew A. Chien. DGMonitor: a performance monitoring tool for sandboxbased desktop Grid platforms. The Journal of Supercomputing, 34(2):113– 133, November 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=113.

Jones:2005:CBA

[624] William M. Jones, Walter B. Ligon, Louis W. Pang, and Dan Stanzione. Characterization of bandwidth-aware meta-schedulers for co-allocating jobs across multiple clusters. The Journal of Supercomputing, 34(2):135– 163, November 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=135.

Jin:2005:RTP

[625] Hai Jin, Xia Xie, Yunfa Li, Zongfen Han, Zhihua Dai, and Peng Lu. A real-time performance evaluation model for distributed software with reliability constrains. The Journal of Supercomputing, 34(2):165–179, November 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=165.

Mathis:2005:GPM

[626] Mark M. Mathis and Darren J. Kerbyson. A general performance model of structured and unstructured mesh particle transport computations. The Journal of Supercomputing, 34(2):181– 199, November 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=181.

Grove:2005:CBP

[627] D. A. Grove and P. D. Coddington. Communication benchmarking and performance modelling of MPI programs on cluster computers. The Journal of Supercomputing, 34(2):201– 217, November 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=201.

Anonymous:2005:CAi

[628] Anonymous. Contributing authors.

The Journal of Supercomputing, 34(2): 219–223, November 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=2&spage=219.

Grossman:2005:SAB

[629] Robert L. Grossman, M. Mazzucco, H. Sivakumar, Y. Pan, and Q. Zhang. Simple available bandwidth utilization library for high-speed wide area networks. The Journal of Supercomputing, 34(3):231–242, December 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=34&issue=3&spage=231.

Micikevicius:2005:CCD

[630] Paulius Micikevicius and Narsingh Deo. Cluster computing for determining three-dimensional protein structure. The Journal of Supercomputing, 34(3):243–271, December 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=34&issue=3&spage=243.

Jie:2005:ISG

[631] Wei Jie, Terence Hung, and Wentong Cai. An information service for Grid virtual organization: Architecture, implementation and evaluation. The Journal of Supercomputing, 34(3): 273–290, December 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=34&issue=3&spage=273.

Lin:2005:EDD

[632] Chun-Yuan Lin, Yeh-Ching Chung, and Jen-Shiuh Liu. Efficient data distribution schemes for EKMR-based sparse arrays on distributed memory multicomputers. The Journal of Supercomputing, 34(3):291–313, December 2005. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=3&spage=291.

Yang:2005:EPL

[633] Chao-Tung Yang, Kuan-Wei Cheng, and Kuan-Ching Li. An enhanced parallel loop self-scheduling scheme for cluster environments. The Journal of Supercomputing, 34(3):315– 335, December 2005. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=34&issue=3&spage=315.

Avallone:2006:HPI

[634] Stefano Avallone, Donato Emma, Antonio Pescapè, and Giorgio Ventre. High performance Internet traffic generators. The Journal of Supercomputing, 35(1):5–26, January 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=1&spage=5.

Bradley:2006:DRC

[635] Alan Bradley, Kevin Curran, and Gerard Parr. Discovering resources in computational GRID environments. The Journal of Supercomputing, 35 (1):27–49, January 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=1&spage=27.

Pourazin:2006:CM

[636] Shahriar Pourazin and Ahmad AbdollahZadeh Barforoush. Concurrent metareasoning. The Journal of Supercomputing, 35(1):51– 64, January 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=1&spage=51.

Drosinos:2006:EPT

[637] Nikolaos Drosinos and Nectarios Koziris. The effect of process topology and load balancing on parallel programming models for SMP clusters and iterative algorithms. The Journal of Supercomputing, 35(1):65–91, January 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=1&spage=65.

Wang:2006:SPS

[638] Ling Wang, Yingtao Jiang, and Henry Selvaraj. Scheduling and partitioning schemes for low power designs using multiple supply voltages. The Journal of Supercomputing, 35(1):93– 113, January 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=1&spage=93.

Wan:2006:HDR

[639] Hon-Hing Wan and Yu-Kwong Kwok.

High data rate video transmission using parallel TCP connections: Approaches and performance evaluation. The Journal of Supercomputing, 35 (2):119–139, February 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=2&spage=119.

Park:2006:LBR

[640] Soyeon Park and Seung Ryoul Maeng.

Log-based rollback recovery without checkpoints of shared memory in software DSM. The Journal of Supercomputing, 35(2):141– 154, February 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=2&spage=141.

Akl:2006:PAE

[641] Selim G. Akl and Weiguang Yao. A parallel approach eliminates measurement perturbations in RLC circuits. The Journal of Supercomputing, 35 (2):155–164, February 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=2&spage=155.

Li:2006:ECR

[642] Keqiu Li, Hong Shen, Keishi Tajima, and Liusheng Huang. An effective cache replacement algorithm in transcoding-enabled proxies. The Journal of Supercomputing, 35(2):165– 184, February 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=2&spage=165.

Galanis:2006:PIP

[643] Michalis D. Galanis, Gregory Dimitroulakos, and Costas E. Goutis. Performance improvements from partitioning applications to FPGA hardware in embedded SoCs. The Journal of Supercomputing, 35(2):185– 199, February 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=2&spage=185.

Ould-Khaoua:2006:PME

[644] Mohamed Ould-Khaoua and Geyong Min. Performance modelling and evaluation of parallel and distributed systems. The Journal of Supercomputing, 35(3):203–204, March 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=3&spage=203.

Li:2006:OPW

[645] Keqin Li. Optimal period of workload redistribution for dynamic bulk synchronous computations in heterogeneous computing systems. The Journal of Supercomputing, 35(3):205– 226, March 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=3&spage=205.

Bahi:2006:PCP

[646] Jacques M. Bahi, Sylvain Contassot-Vivier, and Raphaël Couturier. Performance comparison of parallel programming environments for implementing AIAC algorithms. The Journal of Supercomputing, 35(3):227– 244, March 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=3&spage=227.

Min:2006:SAD

[647] Geyong Min, Mohamed Ould-Khaoua, Demetres D. Kouvatsos, and Irfan U. Awan. Stochastic analysis of deterministic routing algorithms in the presence of self-similar traffic. The Journal of Supercomputing, 35(3):245– 258, March 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=35&issue=3&spage=245.

Boukerche:2006:SPE

[648] Azzedine Boukerche and Caron Dzermajko. Scalability and performance evaluation of DDM-based aggregation/ dissaggregation protocols for large-scale distributed interactive simulations systems. The Journal of Supercomputing, 35(3):259–276, March 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=3&spage=259.

Tian:2006:NMN

[649] Hao Tian, Ajay K. Katangur, Jiling Zhong, and Yi Pan. A novel multistage network architecture with multicast and broadcast capability. The Journal of Supercomputing, 35 (3):277–300, March 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=3&spage=277.

Xu:2006:TNH

[650] Zhiyong Xu, Laxmi Bhuyan, and Yiming Hu. Tulip: a new hash based cooperative Web caching architecture. The Journal of Supercomputing, 35 (3):301–320, March 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=35&issue=3&spage=301.

Chen:2006:WAR

[651] Yawen Chen, Hong Shen, and Fangai Liu. Wavelength assignment for realizing parallel FFT on regular optical networks. The Journal of Supercomputing, 36(1):3–16, April 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=1&spage=3.

Yu:2006:MOG

[652] Jia Yu, Srikumar Venugopal, and Rajkumar Buyya. A market-oriented Grid directory service for publication and discovery of Grid service providers and their services. The Journal of Supercomputing, 36(1):17–31, April 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=1&spage=17.

Sklavos:2006:DAP

[653] N. Sklavos, P. Kitsos, K. Papadopoulos, and O. Koufopavlou. Design, architecture and performance evaluation of the wireless transport layer security. The Journal of Supercomputing, 36(1):33–50, April 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=1&spage=33.

Fuzitaki:2006:NCD

[654] Claudio Naoto Fuzitaki, Paulo Blauth Menezes, Júlio Pereira Machado, and Fernando D’Andrea. Nautilus, a concurrent diagrammatic specification and programming language. The Journal of Supercomputing, 36(1): 51–81, April 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=1&spage=51.

Parsa:2006:NAP

[655] Saeed Parsa and Shahriar Lotfi. A new approach to parallelization of serial nested loops using genetic algorithms. The Journal of Supercomputing, 36(1):83–94, April 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=1&spage=83.

Oldehoeft:2006:CSS

[656] Rod Oldehoeft. Computer science in support of high-performance applications: Papers from the 2004 LACSI Symposium. The Journal of Supercomputing, 36(2):99–100, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=2&spage=99.

Coarfa:2006:ESI

[657] Cristian Coarfa, Yuri Dotsenko, and John Mellor-Crummey. Experiences with Sweep3D implementations in Coarray Fortran. The Journal of Supercomputing, 36(2):101–121, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=2&spage=101.

Rickett:2006:RPF

[658] Christopher D. Rickett, Sung-Eun Choi, Craig E. Rasmussen, and Matthew J. Sottile. Rapid prototyping frameworks for developing scientific applications: a case study. The Journal of Supercomputing, 36(2):123–134, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=36&issue=2&spage=123.

Cooper:2006:ESS

[659] Keith D. Cooper, Alexander Grosul, Timothy J. Harvey, Steve Reeves, Devika Subramanian, Linda Torczon, and Todd Waterman. Exploring the structure of the space of compilation sequences using randomized search algorithms. The Journal of Supercomputing, 36(2):135–151, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=2&spage=135.

Krishnamoorthy:2006:LTS

[660] Sriram Krishnamoorthy, Gerald Baumgartner, Chi-Chung Lam, Jarek Nieplocha, and P. Sadayappan. Layout transformation support for the disk resident arrays framework. The Journal of Supercomputing, 36(2):153–170, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=2&spage=153.

Hendriks:2006:HBF

[661] Erik A. Hendriks and Ronald G. Minnich. How to build a fast and reliable 1024 node cluster with only one disk. The Journal of Supercomputing, 36(2):171–181, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=2&spage=171.

Qasem:2006:ATW

[662] Apan Qasem, Ken Kennedy, and John Mellor-Crummey. Automatic tuning of whole applications using direct search and a performance-based transformation system. The Journal of Supercomputing, 36(2):183–196, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=2&spage=183.

Guo:2006:F

[663] Minyi Guo, Jiannong Cao, and Laurence T. Yang. Foreword. The Journal of Supercomputing, 36(3): 199–200, June 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=3&spage=199.

Lin:2006:OSP

[664] Yih-Fang Lin, Chien-Min Wang, and Jan-Jan Wu. Optimizing I/O server placement for parallel I/O on switchbased irregular networks. The Journal of Supercomputing, 36(3):201– 217, June 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=3&spage=201.

Takizawa:2006:HPP

[665] Hiroyuki Takizawa and Hiroaki Kobayashi. Hierarchical parallel processing of large scale data clustering on a PC cluster with GPU co-processing. The Journal of Supercomputing, 36(3): 219–234, June 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=3&spage=219.

Xiao:2006:NAD

[666] Bin Xiao, Wei Chen, and Yanxiang He. A novel approach to detecting DDoS attacks at an early stage. The Journal of Supercomputing, 36 (3):235–248, June 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=3&spage=235.

Wang:2006:ASS

[667] Tianqi Wang, Cho-Li Wang, and Francis C. M. Lau. An architecture to support scalable distributed virtual environment systems on Grid. The Journal of Supercomputing, 36 (3):249–264, June 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=3&spage=249.

Kwong:2006:CAS

[668] Kin Wah Kwong and Danny H. K.

Tsang. A congestion-aware search protocol for heterogeneous peer-to-peer networks. The Journal of Supercomputing, 36(3):265–282, June 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=3&spage=265.

Wu:2006:CND

[669] Richard S. L. Wu, Allan K. Y. Wong, and Tharam S. Dillon. CACHE_{RP}: a novel dynamic cache size tuning model working with relative object popularity for fast Web information retrieval. The Journal of Supercomputing, 36(3):283–296, June 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=3&spage=283.

Ngo:2006:ABS

[670] Son Hong Ngo, Xiaohong Jiang, Vinh Trong Le, and Susumu Horiguchi. Ant-based survivable routing in dynamic WDM networks with shared backup paths. The Journal of Supercomputing, 36(3):297–307, June 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=3&spage=297.

Xavier:2006:WMC

[671] Percival Xavier, Wentong Cai, and Bu-Sung Lee. Workload management of cooperatively federated computing clusters. The Journal of Supercomputing, 36(3):309–322, June 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=36&issue=3&spage=309.

Cai:2006:CA

[672] Wentong Cai, Jiannong Cao, Wei Chen, T. S. Dillon, Minyi Guo, Susumu Horiguchi, Xiaohong Jiang, Hiroaki Kobayashi, Kin Wah Kwong, Francis C. M. Lau, Vinh Trong Le, Bu-Sung Lee, Son Hong Ngo, Hiroyuki Takizawa, Danny Tsang, Cho-Li Wang, Tianqi Wang, Allan K. Y. Wong, Richard S. L. Wu, Percival Xavier, Bin Xiao, Laurence T. Yang, and Yanxiang He. Contributing authors. The Journal of Supercomputing, 36 (3):323–330, June 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=36&issue=3&spage=323.

Ahuja:2006:SWG

[673] Sanjay P. Ahuja and Jack R. Myers. A survey on wireless Grid computing. The Journal of Supercomputing, 37(1):3–21, July 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=1&spage=3.

HoseinyFarahabady:2006:GPG

[674] M. R. HoseinyFarahabady and H. Sarbazi-Azad. The Grid-Pyramid: a generalized pyramid network. The Journal of Supercomputing, 37(1):23–45, July 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=1&spage=23.

Ohn:2006:DAC

[675] Kyungoh Ohn and Haengrae Cho. Dynamic affinity cluster allocation in a shared disks cluster. The Journal of Supercomputing, 37(1): 47–69, July 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=1&spage=47.

Luo:2006:DMM

[676] Congnan Luo, Anil L. Pereira, and Soon M. Chung. Distributed mining of maximal frequent itemsets on a data Grid system. The Journal of Supercomputing, 37(1):71–90, July 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=1&spage=71.

Wang:2006:DIE

[677] Hsiao-Hsi Wang, Kuan-Ching Li, Kuo-Jen Wang, and Ssu-Hsuan Lu. On the design and implementation of an effective prefetch strategy for DSM systems. The Journal of Supercomputing, 37(1):91–112, July 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=1&spage=91.

Sinha:2006:DBG

[678] Koushik Sinha and Pradip K. Srimani. Deterministic broadcast and gossiping algorithms for ad hoc networks. The Journal of Supercomputing, 37(2):115–144, August 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=2&spage=115.

Chang:2006:TDS

[679] Jyh-Biau Chang, Ce-Kuen Shieh, and Tyng-Yeu Liang. A transparent distributed shared memory for clustered symmetric multiprocessors. The Journal of Supercomputing, 37 (2):145–160, August 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=2&spage=145.

Youn:2006:FTR

[680] Jong-Hoon Youn, Bella Bose, and Seungjin Park. Fault-tolerant routing algorithm in meshes with solid faults. The Journal of Supercomputing, 37 (2):161–177, August 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=2&spage=161.

Kakarountas:2006:HSF

[681] Athanasios P. Kakarountas, Haralambos Michail, Athanasios Milidonis, Costas E. Goutis, and George Theodoridis. Highspeed FPGA implementation of secure hash algorithm for IPSec and VPN applications. The Journal of Supercomputing, 37(2):179–195, August 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=2&spage=179.

Kargahi:2006:MPA

[682] Mehdi Kargahi and Ali Movaghar. A method for performance analysis of earliest-deadline-first scheduling policy. The Journal of Supercomputing, 37 (2):197–222, August 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=2&spage=197.

Chang:2006:CA

[683] Jyh-Biau Chang, Tyng-Yeu Liang, and Ce-Kuen Shieh. Contributing authors. The Journal of Supercomputing, 37(2):223, August 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=2&spage=223.

Reza:2006:MAD

[684] Hassan Reza. A methodology for architectural design of concurrent and distributed software systems. The Journal of Supercomputing, 37(3):227– 248, September 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=37&issue=3&spage=227.

Parsa:2006:NGA

[685] Saeed Parsa and Shahriar Lotfi. A new genetic algorithm for loop tiling. The Journal of Supercomputing, 37(3): 249–269, September 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=3&spage=249.

Gajin:2006:EPD

[686] Slavko Gajin and Zoran Jovanović.

Explanation of performance degradation in turn model. The Journal of Supercomputing, 37(3):271– 295, September 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=3&spage=271.

Huang:2006:ECS

[687] Jih-Woei Huang and Chih-Ping Chu.

An efficient communication scheduling method for the processor mapping technique applied data redistribution. The Journal of Supercomputing, 37(3):297–318, September 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=3&spage=297.

Karra:2006:FBS

[688] M. Ch. Karra and M. P. Bekakos.

A FPGA-based systolic array prototype implementing the quadrant interlocking factorization method. The Journal of Supercomputing, 37(3):319– 331, September 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=37&issue=3&spage=319.

Abderazek:2006:HLM

[689] Ben A. Abderazek, Tsutomu Yoshinaga, and Masahiro Sowa. High-level modeling and FPGA prototyping of produced order parallel queue processor core. The Journal of Supercomputing, 38(1):3–15, October 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=1&spage=3.

Galanis:2006:PMH

[690] Michalis D. Galanis, Gregory Dimitroulakos, and Costas E. Goutis. Partitioning methodology for heterogeneous reconfigurable functional units. The Journal of Supercomputing, 38 (1):17–34, October 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=1&spage=17.

Deng:2006:PIK

[691] Junjun Deng, Hengyong Yu, Jun Ni, Tao He, Shiying Zhao, Lihe Wang, and Ge Wang. A parallel implementation of the katsevich algorithm for 3-D CT image reconstruction. The Journal of Supercomputing, 38 (1):35–47, October 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=1&spage=35.

Petit:2006:AWC

[692] Salvador Petit, Julio Sahuquillo, Ana Pont, and David Kaeli. Addressing a workload characterization study to the design of consistency protocols. The Journal of Supercomputing, 38(1):49–72, October 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=1&spage=49.

Ugur:2006:SPA

[693] Ahmet Ugur and Henry Thompson.

The p-sized partitioning algorithm for fast computation of factorials of numbers. The Journal of Supercomputing, 38(1):73–82, October 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=1&spage=73.

Myoupo:2006:WEB

[694] Jean-Frédéric Myoupo and David Semé. Work-efficient BSR-based parallel algorithms for some fundamental problems in graph theory. The Journal of Supercomputing, 38(1):83– 107, October 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=1&spage=83.

Al-Ayyoub:2006:GMA

[695] A. Al-Ayyoub, A. Awwad, K. Day, and M. Ould-Khaoua. Generalized methods for algorithm development on optical systems. The Journal of Supercomputing, 38(2):111–125, November 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=2&spage=111.

Huh:2006:ARM

[696] Eui-Nam Huh and Lonnie R. Welch. Adaptive resource management for dynamic distributed real-time applications. The Journal of Supercomputing, 38(2):127–142, November 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=2&spage=127.

Hritonenko:2006:CDC

[697] Natali Hritonenko and Yuri Yatsenko. Creative destruction of computing systems: analysis and modeling. The Journal of Supercomputing, 38(2):143– 154, November 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=2&spage=143.

Wang:2006:GML

[698] Junfeng Wang, Hongxia Zhou, Mingtian Zhou, and Lei Li. A general model for long-tailed network traffic approximation. The Journal of Supercomputing, 38(2):155–172, November 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=38&issue=2&spage=155. See

comment [797].

Blais:2006:SHT

[699] J. A. R. Blais, D. A. Provins, and M. A. Soofi. Spherical harmonic transforms for discrete multiresolution applications. The Journal of Supercomputing, 38(2):173–187, November 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=2&spage=173.

Xiao:2006:ACN

[700] Li Xiao, Xiaodong Zhang, Zhengqian Kuang, Baiming Feng, and Jichang Kang. Auto-CFD-NOW: a pre-compiler for effectively parallelizing CFD applications on networks of workstations. The Journal of Supercomputing, 38(2): 189–217, November 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=2&spage=189.

Jigang:2006:AAA

[701] Wu Jigang and Thambipillai Srikanthan. Algorithmic aspects of areaefficient hardware/software partitioning. The Journal of Supercomputing, 38(3):223–235, December 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=3&spage=223.

Ro:2006:DEH

[702] Won W. Ro, Stephen P. Crago, Alvin M. Despain, and Jean-Luc Gaudiot. Design and evaluation of a hierarchical decoupled architecture. The Journal of Supercomputing, 38(3):237– 259, December 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=3&spage=237.

Mohamed:2006:HPM

[703] Nader Mohamed, Jameela Al-Jaroodi, Hong Jiang, and David Swanson. Highperformance message striping over reliable transport protocols. The Journal of Supercomputing, 38(3):261– 278, December 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=3&spage=261.

Volckaert:2006:FGS

[704] Bruno Volckaert, Pieter Thysebaert, Marc De Leenheer, Filip De Turck, Bart Dhoedt, and Piet Demeester. Flexible Grid service management through resource partitioning. The Journal of Supercomputing, 38(3):279– 305, December 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=3&spage=279.

Chen:2006:DSD

[705] Ching-Wen Chen. Design schemes of dynamic rerouting networks with destination tag routing for tolerating faults and preventing collisions. The Journal of Supercomputing, 38(3): 307–326, December 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=38&issue=3&spage=307.

Hababeh:2007:HPC

[706] Ismail Omar Hababeh, Muthu Ramachandran, and Nicholas Bowring. A high-performance computing method for data allocation in distributed database systems. The Journal of Supercomputing, 39(1):3–18, January 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=1&spage=3.

Li:2007:PBM

[707] Yanjun Li and Soon M. Chung.

Parallel bisecting k-means with prediction clustering algorithm. The Journal of Supercomputing, 39(1):19– 37, January 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=1&spage=19.

Hsieh:2007:EPS

[708] Sun-Yuan Hsieh. An efficient parallel strategy for the perfect domination problem on distance-hereditary graphs. The Journal of Supercomputing, 39(1):39–57, January 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=1&spage=39.

Pandey:2007:SCM

[709] Nirved Pandey and G. K. Sharma. Startup comparison for message passing libraries with DTM on Linux clusters. The Journal of Supercomputing, 39(1):59–72, January 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=1&spage=59.

Salleh:2007:SRM

[710] Shaharuddin Salleh, Stephan Olariu, Albert Y. Zomaya, Kiew Leh Yieng, and Nur Arina B. Aziz. Single-row mapping and transformation of connected graphs. The Journal of Supercomputing, 39(1):73–89, January 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=1&spage=73.

Frigo:2007:MBC

[711] Matteo Frigo and Volker Strumpen. The memory behavior of cache oblivious stencil computations. The Journal of Supercomputing, 39(2):93– 112, February 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=2&spage=93.

Obimbo:2007:PAD

[712] Charlie Obimbo and Behzad Salami. A parallel algorithm for determining the inverse of a matrix for use in blockcipher encryption/decryption.

The Journal of Supercomputing, 39 (2):113–130, February 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=2&spage=113.

Tsoi:2007:PAS

[713] Kelvin Yiu-Lun Tsoi and Yu-Kwong Kwok. Practical algorithms for scheduling video data in a local area network environment. The Journal of Supercomputing, 39(2):131– 147, February 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=2&spage=131.

Zhu:2007:ECC

[714] Yingwu Zhu and Yiming Hu. Exploiting client caches to build large Web caches. The Journal of Supercomputing, 39(2):149–175, February 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=2&spage=149.

Chandra:2007:ESP

[715] Sumir Chandra, Xiaolin Li, Taher Saif, and Manish Parashar. Enabling scalable parallel implementations of structured adaptive mesh refinement applications. The Journal of Supercomputing, 39(2):177–203, February 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=2&spage=177.

Jang:2007:DIP

[716] Jae-Wan Jang and Jin-Soo Kim. Design issues and performance comparisons in supporting the sockets interface over user-level communication architecture. The Journal of Supercomputing, 39(2):205–226, February 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=2&spage=205.

Tan:2007:COA

[717] Guangming Tan, Shengzhong Feng, and Ninghui Sun. Cache oblivious algorithms for nonserial polyadic programming. The Journal of Supercomputing, 39(2):227–249, February 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=2&spage=227.

Galanis:2007:ESE

[718] Michalis D. Galanis, Gregory Dimitroulakos, and Costas E. Goutis. Exploring the speedups of embedded microprocessor systems utilizing a highperformance coprocessor data-path. The Journal of Supercomputing, 39 (3):251–271, March 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=39&issue=3&spage=251.

Holt:2007:PMA

[719] John D. Holt and Soon M. Chung. Parallel mining of association rules from text databases. The Journal of Supercomputing, 39(3):273–299, March 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=3&spage=273.

Prasad:2007:BDD

[720] P. W. C. Prasad, Ali Assi, and Azam Beg. Binary decision diagrams and neural networks. The Journal of Supercomputing, 39(3):301–320, March 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=3&spage=301.

Lin:2007:TED

[721] Chin-Feng Lin, Yeh-Ching Chung, and Don-Lin Yang. TRLE — an efficient data compression scheme for image composition of volume rendering on distributed memory multicomputers. The Journal of Supercomputing, 39(3):321–345, March 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=3&spage=321.

Reza:2007:MTC

[722] Hassan Reza and Emanuel S. Grant.

A method to test concurrent systems using architectural specification. The Journal of Supercomputing, 39 (3):347–357, March 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=39&issue=3&spage=347.

Yu:2007:RRS

[723] Jung-Lok Yu, Jin-Soo Kim, and Seung-Ryoul Maeng. A runtime resolution scheme for priority boost conflict in implicit coscheduling. The Journal of Supercomputing, 40(1): 1–28, April 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=1&spage=1.

Javadi:2007:AMI

[724] Bahman Javadi, Jemal H. Abawajy, and Mohammad K. Akbari. Analytical modeling of interconnection networks in heterogeneous multi-cluster systems. The Journal of Supercomputing, 40(1):29–47, April 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=1&spage=29.

Milovanovic:2007:CAP

[725] E. I. Milovanović, I. Z. Milovanović, M. P. Bekakos, and I. N. Tselepis. Computing all-pairs shortest paths on a linear systolic array and hardware realization on a reprogrammable FPGA platform. The Journal of Supercomputing, 40(1):49–66, April 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=40&issue=1&spage=49.

Acebron:2007:SAN

[726] Juan A. Acebrón and Renato Spigler.

Supercomputing applications to the numerical modeling of industrial and applied mathematics problems. The Journal of Supercomputing, 40(1): 67–80, April 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=1&spage=67.

Plaza:2007:IPH

[727] Antonio Plaza, Javier Plaza, and David Valencia. Impact of platform heterogeneity on the design of parallel algorithms for morphological processing of high-dimensional image data. The Journal of Supercomputing, 40 (1):81–107, April 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=1&spage=81.

Han:2007:SPA

[728] Guowen Han and Yuanyuan Yang.

Scheduling and performance analysis of multicast interconnects. The Journal of Supercomputing, 40(2): 109–125, May 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=2&spage=109.

Dimitroulakos:2007:DSE

[729] Grigoris Dimitroulakos, Michalis D. Galanis, and Costas E. Goutis. Design space exploration of an optimized compiler approach for a generic reconfigurable array architecture. The Journal of Supercomputing, 40(2): 127–157, May 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=2&spage=127.

Choi:2007:CPE

[730] Gyu Sang Choi, Jin-Ha Kim, Deniz Ersoz, Andy B. Yoo, and Chita R. Das. A comprehensive performance and energy consumption analysis of scheduling alternatives in clusters. The Journal of Supercomputing, 40 (2):159–184, May 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=2&spage=159.

Jiang:2007:SIB

[731] Xianyang Jiang, Peiheng Zhang, Xinchun Liu, and Stephen S.-T. Yau. Survey on index based homology search algorithms. The Journal of Supercomputing, 40(2):185–212, May 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=2&spage=185.

Wang:2007:BTS

[732] Chia-Wei Wang, Hung-Chang Hsiao, Wen-Hung Sun, Chung-Ta King, and Ming-Tsung Sun. Building a tuple space on structured peer-to-peer networks. The Journal of Supercomputing, 40(3):213–227, June 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=3&spage=213.

Hsu:2007:SCF

[733] Ching-Hsien Hsu, Shih-Chang Chen, and Chao-Yang Lan. Scheduling contention-free irregular redistributions in parallelizing compilers. The Journal of Supercomputing, 40(3): 229–247, June 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=3&spage=229.

Yang:2007:RBE

[734] Chao-Tung Yang, Po-Chi Shih, Cheng-Fang Lin, and Sung-Yi Chen. A resource broker with an efficient network information model on Grid environments. The Journal of Supercomputing, 40(3):249–267, June 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=3&spage=249.

Yang:2007:IDA

[735] Chao-Tung Yang, I-Hsien Yang, Kuan-Ching Li, and Shih-Yu Wang. Improvements on dynamic adjustment mechanism in co-allocation data Grid environments. The Journal of Supercomputing, 40(3):269–280, June 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=3&spage=269.

Chen:2007:SSS

[736] Chung-Kai Chen, Yu-Hao Chang, Yu-Tin Chen, Chih-Chieh Yang, and Jenq-Kuen Lee. Switching supports for stateful object remoting on network processors. The Journal of Supercomputing, 40(3):281–298, June 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=3&spage=281.

Li:2007:DIV

[737] Kuan-Ching Li and Hsun-Chang Chang. The design and implementation of visual performance monitoring and analysis toolkit for cluster and Grid environments. The Journal of Supercomputing, 40(3):299–317, June 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=3&spage=299.

Wang:2007:CSR

[738] Shao-Yang Wang and Rong-Guey Chang. Code size reduction by compressing repeated instruction sequences. The Journal of Supercomputing, 40 (3):319–331, June 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=40&issue=3&spage=319.

Zhou:2007:HSM

[739] Jiazheng Zhou, Xuan-Yi Lin, and Yeh-Ching Chung. Hardware supported multicast in fat-tree-based InfiniBand networks. The Journal of Supercomputing, 40(3):333–352, June 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=40&issue=3&spage=333.

Zhong:2007:PPS

[740] Wei Zhong, Gulsah Altun, Xinmin Tian, Robert Harrison, Phang C. Tai, and Yi Pan. Parallel protein secondary structure prediction schemes using Pthread and OpenMP over hyper-threading technology. The Journal of Supercomputing, 41(1):1–16, July 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=1&spage=1.

Jin:2007:IPO

[741] Hyun-Wook Jin and Chuck Yoo. Impact of protocol overheads on network throughput over high-speed interconnects: measurement, analysis, and improvement. The Journal of Supercomputing, 41(1):17–40, July 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=1&spage=17.

Hsieh:2007:PTS

[742] Sun-Yuan Hsieh and Huang-Ming Gao.

On the partial terminal Steiner tree problem. The Journal of Supercomputing, 41(1):41–52, July 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=1&spage=41.

Ajwa:2007:CSG

[743] Iyad A. Ajwa. A case study of Grid computing and computer algebra: parallel Gröbner bases and characteristic sets. The Journal of Supercomputing, 41(1):53–62, July 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=1&spage=53.

Lin:2007:DDS

[744] Chun-Yuan Lin and Yeh-Ching Chung. Data distribution schemes of sparse arrays on distributed memory multicomputers. The Journal of Supercomputing, 41(1):63–87, July 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=1&spage=63.

Park:2007:EPA

[745] Sung Hoon Park. The election problem in asynchronous distributed systems with bounded faulty processes. The Journal of Supercomputing, 41 (1):89–104, July 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=1&spage=89.

Chiu:2007:HPA

[746] Steve C. Chiu. High performance I/O architectures and systems. The Journal of Supercomputing, 41(2):105– 108, August 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=2&spage=105.

Nomura:2007:PHM

[747] Ken-Ichi Nomura, Rajiv K. Kalia, Aiichiro Nakano, Priya Vashishta, and Jorge L. Landa. Parallel history matching and associated forecast at the center for interactive smart oilfield technologies. The Journal of Supercomputing, 41(2):109–117, August 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=2&spage=109.

Li:2007:TRC

[748] Yaohang Li, Daniel Chen, and Xiaohong Yuan. Trustworthy remote compiling services for Grid-based scientific applications. The Journal of Supercomputing, 41(2):119–131, August 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=2&spage=119.

Suarez:2007:AGT

[749] Tiki L. Suarez. Access Grid technology in classroom and research environments. The Journal of Supercomputing, 41(2):133–145, August 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=2&spage=133.

Jin:2007:FFD

[750] Hong Jin, Danli Wang, Hongan Wang, and Hui Wang. Feedback fuzzy-DVS scheduling of control tasks. The Journal of Supercomputing, 41(2):147– 162, August 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=2&spage=147.

Chiu:2007:NDL

[751] Steve C. Chiu, Alok N. Choudhary, and Danli Wang. Network and devicelevel impacts: performance and reliability of active I/O storage systems. The Journal of Supercomputing, 41 (2):163–178, August 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=2&spage=163.

Son:2007:REC

[752] Seung Woo Son, Konrad Malkowski, Guilin Chen, Mahmut Kandemir, and Padma Raghavan. Reducing energy consumption of parallel sparse matrix applications through integrated link/CPU voltage scaling. The Journal of Supercomputing, 41(3):179– 213, September 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=41&issue=3&spage=179.

Safaei:2007:PAF

[753] F. Safaei, A. Khonsari, M. Fathy, and M. Ould-Khaoua. Performance analysis of fault-tolerant routing algorithm in wormhole-switched interconnections. The Journal of Supercomputing, 41(3): 215–245, September 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=3&spage=215.

Shih:2007:PBP

[754] Wen-Chung Shih, Chao-Tung Yang, and Shian-Shyong Tseng. A performancebased parallel loop scheduling on Grid environments. The Journal of Supercomputing, 41(3):247– 267, September 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=3&spage=247.

Imani:2007:PLB

[755] N. Imani, H. Sarbazi-Azad, and S. G.

Akl. Perfect load balancing on the star interconnection network. The Journal of Supercomputing, 41(3):269– 286, September 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=41&issue=3&spage=269.

Gravvanis:2007:SIG

[756] George A. Gravvanis, John P. Morrison, and Geoffrey C. Fox. Special issue on “Grid Technologies”. The Journal of Supercomputing, 42 (1):1–2, October 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=1&spage=1.

Stockinger:2007:DGS

[757] Heinz Stockinger. Defining the Grid: a snapshot on the current view. The Journal of Supercomputing, 42(1):3– 17, October 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=1&spage=3.

Scherson:2007:SDG

[758] Isaac D. Scherson, Enrique Cauich, and Daniel S. Valencia. Service discovery for GRID computing using LCAN-mapped hierarchical directories. The Journal of Supercomputing, 42 (1):19–32, October 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=1&spage=19.

Mehta:2007:DRA

[759] Ashish M. Mehta, Jay Smith, H. J. Siegel, Anthony A. Maciejewski, Arun Jayaseelan, and Bin Ye. Dynamic resource allocation heuristics that manage tradeoff between makespan and robustness. The Journal of Supercomputing, 42(1):33–58, October 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=1&spage=33.

Thysebaert:2007:DLS

[760] Pieter Thysebaert, Bruno Volckaert, Marc De Leenheer, Filip De Turck, Bart Dhoedt, and Piet Demeester. Dimensioning and on-line scheduling in Lambda Grids using divisible load concepts. The Journal of Supercomputing, 42(1):59–82, October 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=1&spage=59.

Baker:2007:TWA

[761] Mark A. Baker and Matthew Grove.

Tycho: a wide-area messaging framework with an integrated virtual registry. The Journal of Supercomputing, 42(1):83–106, October 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=1&spage=83.

Castain:2007:CTD

[762] Ralph H. Castain and Jeffrey M.

Squyres. Creating a transparent, distributed, and resilient computing environment: the OpenRTE project. The Journal of Supercomputing, 42 (1):107–123, October 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=1&spage=107.

Kurzyniec:2007:UCA

[763] Dawid Kurzyniec, Magdalena Slawińska, Jarosław Slawiński, and Vaidy Sunderam. Unibus: a contrarian approach to Grid computing. The Journal of Supercomputing, 42(1):125– 144, October 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=1&spage=125.

Giannoutakis:2007:MHP

[764] K. M. Giannoutakis, G. A. Gravvanis, B. Clayton, A. Patil, T. Enright, and J. P. Morrison. Matching high performance approximate inverse preconditioning to architectural platforms. The Journal of Supercomputing, 42(2): 145–163, November 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=2&spage=145.

Konstantopoulos:2007:OHS

[765] Charalampos Konstantopoulos and Christos Kaklamanis. Optimal hypercube simulation on the partitioned optical passive stars network. The Journal of Supercomputing, 42(2):165– 180, November 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=42&issue=2&spage=165.

Weng:2007:GRM

[766] Chuliang Weng, Minglu Li, and Xinda Lu. Grid resource management based on economic mechanisms. The Journal of Supercomputing, 42(2): 181–199, November 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=2&spage=181.

Lin:2007:EAS

[767] Yung-Chia Lin, Yi-Ping You, Chung-Wen Huang, Jenq Kuen Lee, Wei-Kuan Shih, and Ting-Ting Hwang. Energy-aware scheduling and simulation methodologies for parallel security processors with multiple voltage domains. The Journal of Supercomputing, 42(2):201–223, November 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=2&spage=201.

Hsieh:2007:PEP

[768] Sun-Yuan Hsieh, Tsong-Jie Lin, and Hui-Ling Huang. Panconnectivity and edge-pancyclicity of 3-ary N-cubes. The Journal of Supercomputing, 42(2): 225–233, November 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=2&spage=225.

ElFarag:2007:IUR

[769] Ahmed A. El Farag, Hatem M. El-Boghdadi, and Samir I. Shaheen. Improving utilization of reconfigurable resources using two-dimensional compaction. The Journal of Supercomputing, 42(2):235–250, November 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=2&spage=235.

Li:2007:FGE

[770] Kuan-Ching Li and Yong-Kee Jun. Foreword from Guest Editors. The Journal of Supercomputing, 42(3):251– 252, December 2007. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=3&spage=251.

Luna:2007:UOC

[771] Jesus Luna, Manel Medina, and Oscar Manso. Using OGRO and CertiVeR to improve OCSP validation for Grids. The Journal of Supercomputing, 42(3): 253–266, December 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=3&spage=253.

Wang:2007:OSP

[772] Chien-Min Wang, Chun-Chen Hsu, Pangfeng Liu, Hsi-Min Chen, and Jan-Jan Wu. Optimizing server placement in hierarchical Grid environments. The Journal of Supercomputing, 42(3): 267–282, December 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=42&issue=3&spage=267.

Liu:2007:OSA

[773] Pangfeng Liu, Yi-Fang Lin, Jan-Jan Wu, and Zhe-Hao Kang. An optimal scheduling algorithm for an agentbased multicast strategy on irregular networks. The Journal of Supercomputing, 42(3):283–302, December 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=3&spage=283.

Qiu:2007:ESA

[774] Shibin Qiu, Terran Lane, and Cundong Yang. Efficient search algorithms for RNAi target detection. The Journal of Supercomputing, 42(3): 303–319, December 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=3&spage=303.

Hong:2007:GBN

[775] Dong-Suk Hong, Hong-Koo Kang, Dong-Oh Kim, Jae-Kwan Yun, and Ki-Joon Han. A Grid-based node split algorithm for managing current location data of moving objects. The Journal of Supercomputing, 42(3): 321–337, December 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=42&issue=3&spage=321.

Vengerov:2008:RLF

[776] David Vengerov. A reinforcement learning framework for online data migration in hierarchical storage systems. The Journal of Supercomputing, 43(1):1–19, January 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=1&spage=1.

Genaud:2008:EPC

[777] Stéphane Genaud, Pierre Gançarski, Guillaume Latu, Alexandre Blansché, Choopan Rattanapoka, and Damien Vouriot. Exploitation of a parallel clustering algorithm on commodity hardware with P2P-MPI. The Journal of Supercomputing, 43(1):21– 41, January 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=1&spage=21.

Park:2008:RHS

[778] Jin Hwan Park and H. K. Dai. Reconfigurable hardware solution to parallel prefix computation. The Journal of Supercomputing, 43(1):43– 58, January 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=1&spage=43.

Liu:2008:HPP

[779] Yang Liu, Yu-Kwong Kwok, and J. Wang. High performance power control and opportunistic fair scheduling in TH-PPM UWB ad-hoc multimedia networks. The Journal of Supercomputing, 43(1):59–76, January 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=1&spage=59.

Jin:2008:PSM

[780] Shiyuan Jin, Guy Schiavone, and Damla Turgut. A performance study of multiprocessor task scheduling algorithms. The Journal of Supercomputing, 43 (1):77–97, January 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=1&spage=77.

Bahig:2008:PMR

[781] Hazem M. Bahig. Parallel merging with restriction. The Journal of Supercomputing, 43(1):99–104, January 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=1&spage=99.

Guo:2008:AHP

[782] Minyi Guo and Jingling Xue. Advances in high performance computing. The Journal of Supercomputing, 43(2):105–106, February 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=2&spage=105.

Liu:2008:EEP

[783] Ming Liu, Jiannong Cao, Yuan Zheng, Haigang Gong, and Xiaomin Wang. An energy-efficient protocol for data gathering and aggregation in wireless sensor networks. The Journal of Supercomputing, 43(2):107–125, February 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=2&spage=107.

Shao:2008:TET

[784] Zhiyuan Shao, Hai Jin, Bin Cheng, and Wenbin Jiang. ER-TCP: an efficient TCP fault-tolerance scheme for cluster computing. The Journal of Supercomputing, 43(2):127– 145, February 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=2&spage=127.

Xue:2008:IPI

[785] Jingling Xue, Minyi Guo, and Daming Wei. Improving the parallelism of iterative methods by aggressive loop fusion. The Journal of Supercomputing, 43(2):147–164, February 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=2&spage=147.

Lim:2008:RSS

[786] Sang Boem Lim, Hanku Lee, Bryan Carpenter, and Geoffrey Fox. Runtime support for scalable programming in Java. The Journal of Supercomputing, 43(2):165–182, February 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=2&spage=165.

Chen:2008:IRS

[787] Guihai Chen, Tongqing Qiu, and Fan Wu. Insight into redundancy schemes in DHTs. The Journal of Supercomputing, 43(2):183–198, February 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=2&spage=183.

Jin:2008:RSA

[788] Dejiang Jin and Sotirios G. Ziavras. Robust scalability analysis and SPM case studies. The Journal of Supercomputing, 43(3):199–223, March 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=3&spage=199.

Xu-Dong:2008:EED

[789] Chen Xu-Dong, Zhu Qing-Xin, Liao Yong, and Xiong Guang Ze. Endto-end deadline control for aperiodic tasks in distributed real-time systems. The Journal of Supercomputing, 43 (3):225–240, March 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=43&issue=3&spage=225.

Barros:2008:MSP

[790] Fernando J. Barros. Modeling and simulation of parallel adaptive divideand-conquer algorithms. The Journal of Supercomputing, 43(3):241– 255, March 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=3&spage=241.

Shahriar:2008:MBH

[791] Abu Zafar M. Shahriar, M. Mostofa Akbar, M. Sohel Rahman, and Muhammad Abdul Hakim Newton. A multiprocessor based heuristic for multidimensional multiple-choice knapsack problem. The Journal of Supercomputing, 43(3):257–280, March 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=3&spage=257.

Numrich:2008:MSC

[792] Robert W. Numrich. A metric space for computer programs and the principle of computational least action. The Journal of Supercomputing, 43 (3):281–298, March 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=43&issue=3&spage=281.

Yang:2008:DPL

[793] Chao-Tung Yang, Wen-Chung Shih, and Shian-Shyong Tseng. Dynamic partitioning of loop iterations on heterogeneous PC clusters. The Journal of Supercomputing, 44(1): 1–23, April 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=1&spage=1.

Sharifi:2008:YYA

[794] Mohsen Sharifi and Behrouz Zolfaghari.

YAARC: yet another approach to further reducing the rate of conflict misses. The Journal of Supercomputing, 44(1):24–40, April 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=1&spage=24.

Carino:2008:DLB

[795] Ricolindo L. Cariño and Ioana Banicescu. Dynamic load balancing with adaptive factoring methods in scientific applications. The Journal of Supercomputing, 44(1):41–63, April 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=1&spage=41.

Athanasaki:2008:EPL

[796] Evangelia Athanasaki, Nikos Anastopoulos, Kornilios Kourtis, and Nectarios Koziris. Exploring the performance limits of simultaneous multithreading for memory intensive applications. The Journal of Supercomputing, 44(1):64–97, April 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=44&issue=1&spage=64.

Nadarajah:2008:CGM

[797] Saralees Nadarajah. Comment on “A general model for long-tailed network traffic approximation”. The Journal of Supercomputing, 44(1): 98–101, April 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=44&issue=1&spage=98.  See

[698].

Akanda:2008:DEM

[798] Md. Musfiquzzaman Akanda, Ben A. Abderazek, and Masahiro Sowa. Dualexecution mode processor architecture. The Journal of Supercomputing, 44(2):103–125, May 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=44&issue=2&spage=103.

Lai:2008:DPD

[799] Kuan-Chou Lai and Chao-Tung Yang. A dominant predecessor duplication scheduling algorithm for heterogeneous systems. The Journal of Supercomputing, 44(2):126–145, May 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=44&issue=2&spage=126.

Cathey:2008:URD

[800] Rebecca J. Cathey, Steven M. Beitzel, Eric C. Jensen, David Grossman, and Ophir Frieder. Using a relational database for scalable XML search. The Journal of Supercomputing, 44 (2):146–178, May 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=2&spage=146.

Sweeney:2008:HSR

[801] James P. Sweeney and Sanjay P. Ahuja.

Heuristic solutions to resource allocation in Grid computing: a natural approach. The Journal of Supercomputing, 44(2):179–198, May 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=2&spage=179.

Liu:2008:DTP

[802] Yang Liu, Yu-Kwong Kwok, and Jiangzhou Wang. Downlink TCP performance under cross layer rate and power allocation in infrastructure TH-PPM UWB networks. The Journal of Supercomputing, 44(3): 199–216, June 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=3&spage=199.

Liu:2008:MPP

[803] Lu Liu, Nick Antonopoulos, and Stephen Mackin. Managing peer-to-peer networks with human tactics in social interactions. The Journal of Supercomputing, 44(3):217–236, June 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=3&spage=217.

Bernabeu:2008:MPA

[804] Miguel O. Bernabeu, Pedro Alonso, and Antonio M. Vidal. A multilevel parallel algorithm to solve symmetric Toeplitz linear systems. The Journal of Supercomputing, 44(3): 237–256, June 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=3&spage=237.

Sabzehparvar:2008:MMM

[805] Majid Sabzehparvar and S. Mohammad Seyed-Hosseini. A mathematical model for the multi-mode resourceconstrained project scheduling problem with mode dependent time lags. The Journal of Supercomputing, 44 (3):257–273, June 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=3&spage=257.

Santos:2008:EAP

[806] Eunice E. Santos, Jeffrey M. Rickman, Gayathri Muthukrishnan, and Shuangtong Feng. Efficient algorithms for parallelizing Monte Carlo simulations for 2D Ising spin models. The Journal of Supercomputing, 44(3):274–290, June 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=44&issue=3&spage=274.

Li:2008:NCM

[807] Chung-Sheng Li, Yung-Chih Tseng, Han-Chieh Chao, and Yueh-Min Huang. A neighbor caching mechanism for handoff in IEEE 802.11 wireless networks. The Journal of Supercomputing, 45(1):1–14, July 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=1&spage=1.

Bellavista:2008:DCA

[808] Paolo Bellavista, Antonio Corradi, and Luca Foschini. Dynamic and context-aware streaming adaptation to smooth quality degradation due to IEEE 802.11 performance anomaly. The Journal of Supercomputing, 45 (1):15–28, July 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=1&spage=15.

Seo:2008:RST

[809] Jeong-Wook Seo, Sang-Soo Yeo, Kyoo Seok Park, Hang Bae Chang, Sung-Eon Cho, and Dong-Ku Kim. A robust and secure time-domain ICI canceller for OFDM based ubiquitous systems in time-varying multipath channels. The Journal of Supercomputing, 45(1):29–43, July 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=45&issue=1&spage=29.

Kausar:2008:SEK

[810] Firdous Kausar, Sajid Hussain, Laurence T. Yang, and Ashraf Masood. Scalable and efficient key management for heterogeneous sensor networks. The Journal of Supercomputing, 45(1):44–65, July 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=45&issue=1&spage=44.

Li:2008:TCP

[811] Haiyan Li, Chunyuan Zhang, Li Li, and Ju Ren. Transform coding on programmable stream processors. The Journal of Supercomputing, 45 (1):66–87, July 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=45&issue=1&spage=66.

Lee:2008:UMF

[812] Deok Gyu Lee, Jong Hyuk Park, Tai-Hoon Kim, and Laurence T. Yang. Umultimedia framework: a secure and intelligent multimedia service framework based on context information in U-home. The Journal of Supercomputing, 45(1):88–104, July 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=45&issue=1&spage=88.

Cheng:2008:DSA

[813] Hongju Cheng, Naixue Xiong, Larence T.

Yang, and Young-Sik Jeong. Distributed scheduling algorithms for channel access in TDMA wireless mesh networks. The Journal of Supercomputing, 45(1):105–128, July 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=1&spage=105.

Hsu:2008:IRU

[814] Ching-Hsien Hsu, Tai-Lung Chen, and Jong-Hyuk Park. On improving resource utilization and system throughput of master slave job scheduling in heterogeneous systems. The Journal of Supercomputing, 45(1): 129–150, July 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=1&spage=129.

Huang:2008:FPM

[815] Jih-Woei Huang and Chih-Ping Chu.

A flexible processor mapping technique toward data localization for block-cyclic data redistribution. The Journal of Supercomputing, 45(2):151– 172, August 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=2&spage=151.

Zhu:2008:CDR

[816] Qiang Zhu. On conditional diagnosability and reliability of the BC networks. The Journal of Supercomputing, 45(2):173–184, August 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=45&issue=2&spage=173.

Rao:2008:PAS

[817] Imran Rao and Eui-Nam Huh. A probabilistic and adaptive scheduling algorithm using system-generated predictions for inter-Grid resource sharing. The Journal of Supercomputing, 45(2):185–204, August 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=45&issue=2&spage=185.

Verma:2008:CBU

[818] Akshat Verma and Sandeep Sen. Combating I-O bottleneck using prefetching: model, algorithms, and ramifications. The Journal of Supercomputing, 45(2):205–235, August 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=45&issue=2&spage=205.

Zhou:2008:GDDa

[819] Jing Zhou and Guosun Zeng. A general data dependence analysis for parallelizing compilers. The Journal of Supercomputing, 45(2):236– 252, August 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=2&spage=236.

Zhou:2008:GDDb

[820] Jing Zhou and Guosun Zeng. A general data dependence analysis for parallelizing compilers. The Journal of Supercomputing, 45(2):253, August 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=2&spage=253.

Kolcz:2008:LRN

[821] A. Kolcz and A. Chowdhury. Lexicon randomization for near-duplicate detection with I-Match. The Journal of Supercomputing, 45(3):255– 276, September 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=3&spage=255.

Chang:2008:DDR

[822] Ruay-Shiung Chang and Hui-Ping Chang. A dynamic data replication strategy using access-weights in data grids. The Journal of Supercomputing, 45(3):277–295, September 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=3&spage=277.

Wang:2008:SBA

[823] Junfeng Wang, Jin Liu, and Chundong She. Segment-based adaptive hyper-Erlang model for longtailed network traffic approximation. The Journal of Supercomputing, 45(3): 296–312, September 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=3&spage=296.

Noori:2008:AFA

[824] Hamid Noori, Farhad Mehdipour, Kazuaki Murakami, Koji Inoue, and Morteza Saheb Zamani. An architecture framework for an adaptive extensible processor. The Journal of Supercomputing, 45(3):313– 340, September 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=3&spage=313.

Flores:2008:ECC

[825] Antonio Flores, Juan L. Aragón, and Manuel E. Acacio. An energy consumption characterization of on-chip interconnection networks for tiled CMP architectures. The Journal of Supercomputing, 45(3):341– 364, September 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=45&issue=3&spage=341.

Mahabadi:2008:PLI

[826] Aminollah Mahabadi, Hamid Sarbazi-Azad, Ebrahim Khodaie, and Keivan Navi. Parallel Lagrange interpolation on k-ary n-cubes with maximum channel utilization. The Journal of Supercomputing, 46(1):1–14, October 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=1&spage=1.

Kalantari:2008:FAG

[827] Mohammad Kalantari and Mohammad Kazem Akbari. Fault-aware grid scheduling using performance prediction by workload modeling. The Journal of Supercomputing, 46(1):15– 39, October 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=1&spage=15.

Souravlas:2008:MPS

[828] Stavros Souravlas and Manos Roumeliotis. A message passing strategy for array redistributions in a torus network. The Journal of Supercomputing, 46(1):40–57, October 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=1&spage=40.

Liu:2008:SCH

[829] Yang Liu, Yu-Kwong Kwok, and Jiangzhou Wang. On scheduling and clustering in hierarchical TH-PPM UWB wireless ad hoc networks. The Journal of Supercomputing, 46 (1):58–83, October 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=1&spage=58.

Herrero:2008:HOS

[830] José R. Herrero and Juan J. Navarro. Hypermatrix oriented supernode amalgamation. The Journal of Supercomputing, 46(1):84–104, October 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=1&spage=84.

Chiu:2008:CIH

[831] Steve C. Chiu. Current issues in high performance computing I/O architectures and systems. The Journal of Supercomputing, 46(2):105– 107, November 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=2&spage=105.

Chandy:2008:RDI

[832] John A. Chandy. RAID0.5: design and implementation of a low cost disk array data protection method. The Journal of Supercomputing, 46(2): 108–123, November 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=2&spage=108.

Li:2008:ATS

[833] Yaohang Li and Yong-Duan Song. An adaptive and trustworthy software testing framework on the grid. The Journal of Supercomputing, 46(2):

124–138, November 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=2&spage=124.

Wang:2008:SFD

[834] Danli Wang, Guozhong Dai, Hongan Wang, and Steve C. Chiu. Scenariofocused development method for a penbased user interface: model and applications. The Journal of Supercomputing, 46(2):139–149, November 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=2&spage=139.

Subramaniyan:2008:OCR

[835] Rajagopal Subramaniyan, Eric Grobelny, Scott Studham, and Alan D. George. Optimization of checkpointingrelated I/O for high-performance parallel and distributed computing. The Journal of Supercomputing, 46(2):150– 180, November 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=2&spage=150.

Salem:2008:PMA

[836] Fatima K. Abu Salem and Laurence T. Yang. Parallel methods for absolute irreducibility testing. The Journal of Supercomputing, 46(3):181– 212, December 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=3&spage=181.

Isaila:2008:MFD

[837] Florin Isaila and Walter F. Tichy. Mapping functions and data redistribution for parallel files. The Journal of Supercomputing, 46(3):213– 236, December 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=3&spage=213.

Xiao:2008:DSU

[838] Bin Xiao, Jiannong Cao, and Qin Lu. Dynamic SPT update for multiple link state decrements in network routing. The Journal of Supercomputing, 46(3): 237–256, December 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=3&spage=237.

Song:2008:HRM

[839] Eun-Ha Song, Laurence T. Yang, and Young-Sik Jeong. 3-hierarchical resource management model on web grid service architecture. The Journal of Supercomputing, 46(3):257– 275, December 2008. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=46&issue=3&spage=257.

Mahafzah:2008:LBP

[840] Basel A. Mahafzah and Bashira A. Jaradat. The load balancing problem in OTIS-hypercube interconnection networks. The Journal of Supercomputing, 46(3):276–297, December 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=46&issue=3&spage=276.

Li:2009:WDC

[841] Hui Li. Workload dynamics on clusters and grids. The Journal of Supercomputing, 47(1):1–20, January 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=1&spage=1.

Rice:2009:SAN

[842] Kenneth L. Rice, Tarek M. Taha, and Christopher N. Vutsinas. Scaling analysis of a neocortex inspired cognitive model on the Cray XD1. The Journal of Supercomputing, 47 (1):21–43, January 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=1&spage=21.

Gravvanis:2009:OBP

[843] George A. Gravvanis. OpenMP based parallel normalized direct methods for sparse finite element linear systems. The Journal of Supercomputing, 47 (1):44–52, January 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=1&spage=44.

Singh:2009:CIB

[844] David E. Singh, Florin Isaila, Juan C. Pichel, and Jesús Carretero. A collective I/O implementation based on inspector-executor paradigm. The Journal of Supercomputing, 47(1):53– 75, January 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=47&issue=1&spage=53. See er-

ratum [855].

Yang:2009:DIW

[845] Chao-Tung Yang, Kuan-Chou Lai, and Po-Chi Shih. Design and implementation of a workflow-based resource broker with information system on computational grids. The Journal of Supercomputing, 47(1):76–109, January 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=47&issue=1&spage=76.

Wang:2009:TIN

[846] Hsiao-Hsi Wang, Kuan-Ching Li, Ssu-Hsuan Lu, and Chun-Chieh Yang. Towards implementation of a novel scheme for data prefetching on distributed shared memory systems. The Journal of Supercomputing, 47(2):111– 126, February 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=47&issue=2&spage=111.

Wang:2009:ALU

[847] Guojun Wang, Tian Wang, Weijia Jia, Minyi Guo, and Jie Li. Adaptive location updates for mobile sinks in wireless sensor networks. The Journal of Supercomputing, 47(2):127– 145, February 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=2&spage=127.

Kola:2009:SAF

[848] Kiran Kola, Ruppa K. Thulasiram, and Parimala Thulasiraman. A software architecture framework for on-line option pricing. The Journal of Supercomputing, 47(2):146–170, February 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=2&spage=146.

Yang:2009:MBS

[849] Xuejun Yang, Jing Du, Xiaobo Yan, and Yu Deng. Matrix-based streamization approach for improving locality and parallelism on FT64 stream processor. The Journal of Supercomputing, 47(2):171–197, February 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=2&spage=171.

Roy:2009:FDE

[850] Banani Roy, Michael Einhaus, and Chanchal Kumar Roy. A framework for development and evaluation of a dynamic subchannel allocation scheme in an OFDMA system. The Journal of Supercomputing, 47 (2):198–227, February 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=2&spage=198.

Amamiya:2009:CBN

[851] Satoshi Amamiya, Makoto Amamiya, Ryuzo Hasegawa, and Hiroshi Fujita. A continuation-based noninterruptible multithreading processor architecture. The Journal of Supercomputing, 47 (2):228–252, February 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=2&spage=228.

Carretero:2009:SDM

[852] Jesus Carretero and Jose Daniel Garcia. Scalability in data management. The Journal of Supercomputing, 47 (3):253–254, March 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=3&spage=253.

Fukuda:2009:IPF

[853] Munehiro Fukuda and Jumpei Miyauchi. An implementation of parallel file distribution in an agent hierarchy. The Journal of Supercomputing, 47 (3):255–285, March 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=47&issue=3&spage=255.

Secretan:2009:EAC

[854] Jimmy Secretan, Malachi Lawson, and Ladislau Bölöni. Efficient allocation and composition of distributed storage. The Journal of Supercomputing, 47(3):286–310, March 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=3&spage=286.

Singh:2009:ECI

[855] David E. Singh, Florin Isaila, Juan C.

Pichel, and Jesús Carretero. Erratum: “A collective I/O implementation based on inspector-executor paradigm”. The Journal of Supercomputing, 47(3):311, March 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=3&spage=311. See

[844].

Calderon:2009:FTF

[856] A. Calderón, F. García-Carballeira, L. M. Sánchez, J. D. García, and J. Fernandez. Fault tolerant file models for parallel file systems: introducing distribution patterns for every file. The Journal of Supercomputing, 47(3):312–334, March 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=47&issue=3&spage=312.

Deng:2009:PIC

[857] Junjun Deng, Hengyong Yu, Jun Ni, Lihe Wang, and Ge Wang. Parallelism of iterative CT reconstruction based on local reconstruction algorithm. The Journal of Supercomputing, 48(1):1–14, April 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=1&spage=1.

Qiu:2009:PES

[858] Tongqing Qiu, Edward Chan, Mao Ye, Guihai Chen, and Ben Y. Zhao. Peer-exchange schemes to handle mismatch in peer-to-peer systems. The Journal of Supercomputing, 48(1): 15–42, April 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=1&spage=15.

Li:2009:PBP

[859] Kuan-Ching Li and Tien-Hsiung Weng. Performance-based parallel application toolkit for high-performance clusters. The Journal of Supercomputing, 48(1):43–65, April 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=1&spage=43.

Chang:2009:SSC

[860] Chung-Hao Chang, Cheng-Kuan Lin, Jimmy J. M. Tan, Hua-Min Huang, and Lih-Hsing Hsu. The super spanning connectivity and super spanning laceability of the enhanced hypercubes. The Journal of Supercomputing, 48(1):66–87, April 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=1&spage=66.

Aliaga:2009:TPG

[861] Jose Ignacio Aliaga, Francisco Almeida, Jose Manuel Badía, Sergio Barrachina, Vicente Blanco, et al. Toward the parallelization of GSL. The Journal of Supercomputing, 48(1): 88–114, April 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=1&spage=88.

Dimitroulakos:2009:CAA

[862] Grigorios Dimitroulakos, Nikos Kostaras, Michalis D. Galanis, and Costas E. Goutis. Compiler assisted architectural exploration framework for coarse grained reconfigurable arrays. The Journal of Supercomputing, 48 (2):115–151, May 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=2&spage=115.

Chen:2009:MC

[863] Zhao Chen. Mapping computations.

The Journal of Supercomputing, 48 (2):152–162, May 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=2&spage=152.

Caire:2009:DIL

[864] Mario E. Caire, Francisco J. Lopez, and David H. Williams. Distributed identification of the lineality space of a cone. The Journal of Supercomputing, 48(2):163–182, May 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=2&spage=163.

dAuriol:2009:SEP

[865] Brian J. dÁuriol. The systems edge of the Parameterized Linear Array with a Reconfigurable Pipelined Bus System (LARPBS(p)) optical bus parallel computing model. The Journal of Supercomputing, 48(2):183–209, May 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=2&spage=183.

Cheng:2009:PBD

[866] Kuan-Wei Cheng, Tzong-Yen Lin, and Rong-Guey Chang. Profilebased dynamic pipeline scaling. The Journal of Supercomputing, 48(2): 210–226, May 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=48&issue=2&spage=210.

Xiang:2009:MEO

[867] Yang Xiang. Managing email overload with an automatic nonparametric clustering system. The Journal of Supercomputing, 48(3):227–242, June 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=3&spage=227.

Lin:2009:OAQ

[868] Chuang Lin, Xin Jiang, Hao Yin, Yuanzhuo Wang, Yada Hu, et al. Optimizing availability and QoS of heterogeneous distributed system based on residual lifetime in uncertain environment. The Journal of Supercomputing, 48(3):243–263, June 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=3&spage=243.

Xu:2009:SPN

[869] Jie Xu and Hai Jin. A structured P2P network based on the small world phenomenon. The Journal of Supercomputing, 48(3):264–285, June 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=3&spage=264.

Konstantopoulos:2009:EPT

[870] Charalampos Konstantopoulos, Basilis Mamalis, Grammati Pantziou, and Damianos Gavalas. Efficient parallel text retrieval techniques on Bulk Synchronous Parallel (BSP)/Coarse Grained Multicomputers (CGM). The Journal of Supercomputing, 48(3): 286–318, June 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=3&spage=286.

Meng:2009:SIS

[871] Sha Meng, Xin Kai Li, and Ronaldo Mercado. Smooth interfaces for spectral element method for the solution of incompressible Newtonian fluid flow. The Journal of Supercomputing, 48(3):319–331, June 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=48&issue=3&spage=319.

Cerin:2009:SIS

[872] Christophe Cérin, Jean-Luc Gaudiot, and Kuan-Ching Li. Special issue of Supercomputing Journal on secure, manageable and controllable grid services. The Journal of Supercomputing, 49(1):1–3, July 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=1.

Deng:2009:ACO

[873] Yuhui Deng, Frank Wang, and Adrian Ciura. Ant colony optimization inspired resource discovery in P2P Grid systems. The Journal of Supercomputing, 49(1):4–21, July 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=4.

deMello:2009:PDN

[874] Rodrigo F. de Mello and Laurence T.

Yang. Prediction of dynamical, nonlinear, and unstable process behavior. The Journal of Supercomputing, 49(1):22–41, July 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=22.

Cheng:2009:QAA

[875] Chieh-Wen Cheng, Jan-Jan Wu, and Pangfeng Liu. QoS-aware, accessefficient, and storage-efficient replica placement in grid environments. The Journal of Supercomputing, 49(1): 42–63, July 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=42.

Hurault:2009:AST

[876] Aurélie Hurault, Michel Daydé, and Marc Pantel. Advanced service trading for scientific computing over the grid. The Journal of Supercomputing, 49(1):64–83, July 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=64.

Sopena:2009:BEM

[877] Julien Sopena, Luciana Arantes, Fabrice Legond-Aubry, and Pierre Sens. Building effective mutual exclusion services for grids. The Journal of Supercomputing, 49(1):84–107, July 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=84.

Mazzoleni:2009:EIF

[878] P. Mazzoleni, B. Crispo, S. Sivasubramanian, and E. Bertino. Efficient integration of fine-grained access control and resource brokering in grid. The Journal of Supercomputing, 49(1):108–126, July 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=108.

Liu:2009:LDL

[879] Dan Liu, Yuan-Da Cao, and Chun-Qing Li. Liana: a decentralized loaddependent scheduler for performancecost optimization of grid service. The Journal of Supercomputing, 49 (1):127–156, July 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=127.

Anonymous:2009:R

[880] Anonymous.  Reviewers.  The

Journal of Supercomputing, 49(1): 157, July 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=1&spage=157.

Li:2009:IHP

[881] Deng Li, Zhi-Gang Chen, Hui Liu, Athanasios V. Vasilakos, and Yi Pan. IPBGA: a hybrid P2P based grid architecture by using information pool protocol. The Journal of Supercomputing, 49(2):159–189, August 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=2&spage=159.

Li:2009:HCP

[882] Chunlin Li and Layuan Li. Hierarchical control policy for dynamic resource management in grid virtual organization. The Journal of Supercomputing, 49(2):190–218, August 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=2&spage=190.

Fu:2009:EFF

[883] Jung-Sheng Fu, Hao-Shun Hung, and Gen-Huey Chen. Embedding fault-free cycles in crossed cubes with conditional link faults. The Journal of Supercomputing, 49(2):219–233, August 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=2&spage=219.

Cho:2009:HSP

[884] Hsun-Jung Cho and Chien-Lun Lan. Hybrid shortest path algorithm for vehicle navigation. The Journal of Supercomputing, 49(2):234–247, August 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=49&issue=2&spage=
                                        234.

Eshaghian-Wilner:2009:EPP

[885] Mary M. Eshaghian-Wilner and Shiva Navab. Efficient parallel processing with spin-wave nanoarchitectures. The Journal of Supercomputing, 49 (2):248–267, August 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=49&issue=2&spage=248.

Milovanovic:2009:SSO

[886] E. I. Milovanović, M. P. Bekakos, and I. Z. Milovanović. Synthesis of space optimal systolic arrays for band matrix-vector multiplication. The Journal of Supercomputing, 49(3): 269–290, September 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=49&issue=3&spage=269.

Sider:2009:FLB

[887] Abderrahmane Sider and Raphaël Couturier. Fast load balancing with the most to least loaded policy in dynamic networks. The Journal of Supercomputing, 49(3):291–317, September 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=3&spage=291.

Zhang:2009:ICH

[888] Yan-Qing Zhang and Yi Pan. Incomplete crossed hypercubes. The Journal of Supercomputing, 49(3):318– 333, September 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=3&spage=318.

Isazadeh:2009:NFM

[889] Ayaz Isazadeh and Jaber Karimpour. A new formalism for mathematical description and verification of component-based systems. The Journal of Supercomputing, 49(3):334– 353, September 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=3&spage=334.

Hsieh:2009:OFT

[890] Sun-Yuan Hsieh and Chang-De Wu.

Optimal fault-tolerant Hamiltonicity of star graphs with conditional edge faults. The Journal of Supercomputing, 49(3): 354–372, September 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=49&issue=3&spage=354.

Yarmish:2009:DSS

[891] Gavriel Yarmish and Richard Van Slyke. A distributed, scaleable simplex method. The Journal of Supercomputing, 49(3):373–381, September 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=49&issue=3&spage=
                                        373.

Nazir:2009:ACS

[892] Babar Nazir, Kalim Qureshi, and Paul Manuel. Adaptive checkpointing strategy to tolerate faults in economy based grid. The Journal of Supercomputing, 50(1):1–18, October 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=50&issue=1&spage=1.

Khanli:2009:AGI

[893] L. Mohammad Khanli and M. Analoui. Active Grid Information Server for grid computing. The Journal of Supercomputing, 50(1):19–35, October 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=50&issue=1&spage=19.

Goumas:2009:PES

[894] Georgios Goumas, Kornilios Kourtis, Nikos Anastopoulos, Vasileios Karakasis, and Nectarios Koziris. Performance evaluation of the sparse matrixvector multiplication on modern architectures. The Journal of Supercomputing, 50(1):36–77, October 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=50&issue=1&spage=36.

Lin:2009:CSG

[895] Yu-Te Lin and Peng-Sheng Chen.

Compiler support for general-purpose computation on GPUs. The Journal of Supercomputing, 50(1):78– 97, October 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=50&issue=1&spage=78.

Lee:2009:FPA

[896] Yong-Ju Lee, Yoo-Hyun Park, Song-Woo Sok, Hag-Young Kim, and Cheol-Hoon Lee. Fast-path I/O architecture for high performance streaming server. The Journal of Supercomputing, 50(2):99–120, November 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=50&issue=2&spage=99.

deMendivil:2009:FAD

[897] José Ramón González de Mendívil, José Enrique Armendáriz-Iñigo, José Ramón Garitagoitia, and Francesc D. Muñoz-Escoí. A formal analysis of database replication protocols with SI replicas and crash failures. The Journal of Supercomputing, 50(2):121– 161, November 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=50&issue=2&spage=121.

Huang:2009:PPC

[898] Chien-Hung Huang and Jywe-Fei Fang. The panconnectivity and the pancycle-connectivity of the generalized base- b hypercube. The Journal of Supercomputing, 50(2):162– 176, November 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=50&issue=2&spage=162.

Yang:2009:DBM

[899] Chao-Tung Yang and Kuan-Chou Lai. A directive-based MPI code generator for Linux PC clusters. The Journal of Supercomputing, 50(2):177– 207, November 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=50&issue=2&spage=177.

Walters:2009:FTS

[900] John Paul Walters and Vipin Chaudhary. A fault-tolerant strategy for virtualized HPC clusters. The Journal of Supercomputing, 50(3):209– 239, December 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=50&issue=3&spage=209.

Chen:2009:BBR

[901] Ching-Wen Chen and Chuan-Chi Weng. Bandwidth-based routing protocols in mobile ad hoc networks. The Journal of Supercomputing, 50(3):240– 268, December 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=50&issue=3&spage=240.

Hsu:2009:SAB

[902] Ching-Hsien Hsu and Bing-Ru Tsai.

Scheduling for atomic broadcast operation in heterogeneous networks with one port model. The Journal of Supercomputing, 50(3):269– 288, December 2009. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=50&issue=3&spage=269.

Lotfi:2009:PLG

[903] Shahriar Lotfi and Saeed Parsa. Parallel loop generation and scheduling. The Journal of Supercomputing, 50(3): 289–306, December 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=50&issue=3&spage=289.

Rau:1993:ILPb

[904] Bantwal Ramakrishna Rau and Joseph A.

Fisher, editors. Instruction-level parallelism: a special issue of The Journal of Supercomputing, volume SECS 235 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-9367-8. 282 pp. LCCN QA76.58 .I49 1993.